A Complete Bibliography of Publications in the Journal of Automated Reasoning

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

08 October 2025 Version 1.30

Title word cross-reference

− [280]. 2 [1222, 610]. 6 [1114]. αx_{1} + ⋯ + αx_{p} ≐_{AC} βy_{1} + ⋯ + βy_{q} [203]. BCDL [816]. ALC [839, 905]. SHOIQ [752, 763]. δ [718, 312]. DL(ALCS)▽_{⇕} [1143]. E [732, 442, 597]. ∃^{∗}∀^{∗} [1235]. ∀∃ [443]. k [678, 713, 1252]. K_{(m)} [559]. Łℵ_{0} [539]. λ [91, 83]. BEL [1034]. EL [867]. SHOIQ [740]. K_{n} [1180]. μ [412]. n [881]. O(n log n) [545]. π [1227, 1091]. π_{4}(S^{3}) [1131]. Q [1126]. Σ [1240]. T [464]. + [1074]. = [1074]. Z [72, 114].

-ary [881]. -Basis [610]. -calculus [412]. -match [280]. -module [72, 114]. -Protocols [1240]. -Resolution [464, 1126]. -Rule [718, 312]. -SAT [678]. -SPACE [545]. -unifiability [91]. -Unification [442, 597, 732].

0.2 [704].

1 [192]. 11 [277, 276]. 13

[394, 397, 403, 385, 387, 402, 386]. 14 [439]. 15 [468]. 16 [500]. 17 [543]. 18 [612].

2 [966]. 2005 [686]. 2008 [943]. 2009 [888]. 2010 [863]. 2015 [1043, 1118]. 2016 [1107, 1078, 1066]. 2017 [1133, 1159, 1182]. 2018 [1194, 1213, 1205]. 2019 [1249]. 2020 [1270]. 22 [854]. 23 [927].

3 [1070].

40 [994].

5.2 [1292].

61131 [1070]. 61131-3 [1070].

92 [485].

= [790, 463].

Abduction [839, 311]. Abductive [1023, 637]. Abelian [60, 241]. ABox [839]. ABoxes [965]. ABR [598]. absence [132]. Absorption [970, 965]. Abstract [728, 617, 1013, 1121, 417, 1230, 894, 1046, 999, 901, 890, 1344, 264, 48]. Abstraction [889, 811, 61]. Abstractions [1029]. Abuse [1362]. AC [1361, 1367, 270, 101, 203, 330]. AC-unification [101, 203]. Acceleration [1230]. Acceptance [618]. Accuracy [1068]. Accurate [1384]. Achievements [984]. Ackermann [189]. ACL2 [760, 547, 727, 793, 851, 616, 859]. acquiring [3]. across [239]. Actions [885]. Active [1164, 1377]. ACUI [643]. Acyclic [905, 1143]. Adding [1017, 59]. addition [82]. Additive [651]. Address [1199, 11, 53]. Addresses [28, 110, 169]. Adequacy [894]. adjunction [231]. Adoption [1366]. Advanced [1222]. Adversarial [418]. against [690]. Agda [1197]. Agents [624]. AI [1162]. Aid [997]. Aided [520, 1310, 195]. AIE [434]. Aircraft [1071]. AKS [1231]. Alan [686]. ALF [595]. Algebra [1203, 1014, 978, 793, 448, 887, 694, 1151, 573, 1379, 610, 1044, 147, 160, 1050]. Algebraic [834, 435, 1177, 1029, 433, 1273]. Algebras [679, 364]. Algorithm [522, 548, 788, 1374, 557, 1231, 1358, 1191, 404, 725, 996, 656, 761, 706, 598, 455, 738, 452, 1143, 519, 1196, 1290, 420, 205, 166, 236, 284, 480]. Algorithmic [943]. Algorithms [1101, 673, 523, 1091, 383, 408, 1109, 491, 502, 1112, 1350, 602, 748, 668, 587, 190, 504, 905, 29, 104, 107]. Allocation [783]. Allowing [925]. Alternative [568, 138, 260, 74, 308, 114]. Always [422]. Amortized [1115, 1116]. Analogy [421, 454, 261]. Analyses [1053]. Analysis [950, 842, 835, 687, 1224, 689, 845, 1363, 1095, 1198, 1338, 654, 547, 1051, 773, 1219, 776, 1054, 848, 758, 1071, 1052, 784, 30, 274]. Analytic [868, 721, 797, 122, 314]. Analytica [447]. analytical [303]. analytically [93, 100]. Analyzing [1032, 637, 929, 1269]. anatomy [344]. Andrew [686]. Android [1315]. Andrzej [989]. Annotation [853]. Announcement [21, 47, 58, 262, 615, 43, 73, 89, 102, 139, 154, 162, 199]. Announcements [10, 54, 66, 81, 111, 444, 473]. Announcing [27, 35]. Answer [747, 701, 366]. Answering [744, 765, 1337, 341]. Answers [777]. Anti [1042, 949, 381, 238, 237]. Anti-Links [381]. Anti-Unification [1042, 949, 238, 237]. Any [368]. Application [1326, 1050, 414, 1167, 505, 1070, 1024, 538, 339, 317, 228, 223]. Applications [434, 1370, 1361, 1367, 1153, 1147, 889, 702, 1109, 916, 1283, 934, 956, 75, 246, 171]. Applied [1350, 418]. Applying [679]. Approach [459, 680, 833, 510, 713, 449, 1163, 895, 620, 446, 700, 848, 1112, 1220, 1313, 213, 127, 370, 308, 254]. Approaches [1124]. Approaching [198]. Approximation [1187, 1036]. Approximations [1068, 1113]. AProVE [1032]. Architecture [1241, 1180]. Area [884, 1303]. argument [239]. Argumentation [579]. Arithmetic [1119, 872, 864, 766, 1300, 694, 1364, 668, 1331, 1033, 1235, 353, 357, 236, 187]. Array [981, 1382, 956]. Arrays [1302]. article [224]. Artificial [434]. ary [881]. Aspects [712]. Assembly [622]. assertion [133]. Assertional [1193]. Assertions [492]. Assistant [734, 1261, 1356, 1331, 1254]. Assistants [586]. Assisted [682, 461, 796, 1363, 964, 189, 975].

Associated [802, 1065, 133]. Associative [554, 711, 230, 368, 61]. associative-commutative [230, 368, 61]. associativity [40]. Assumption [774, 120, 406, 511]. Assumptions [911, 345]. Asymmetric [846]. asymptotic [216]. Asynchronous [1197]. ath [681]. ATP [284, 403, 384, 385, 387, 402, 468, 500, 543, 564, 612, 386, 439, 920]. Attacker [849]. Attacking [692]. Attacks [690]. Attractor [1092]. Auctions [1378]. Augmenting [563]. Autarkic [565]. Autarky [475]. Author [465, 513, 529, 550, 570, 590, 607, 625]. Authoring [735]. Auto2 [1142]. Autoepistemic [304, 248, 305, 323]. Automata [1164, 824, 940, 1163, 1252, 1376]. Automata-Based [824]. Automata-Theoretic [1163]. Automatable [594]. Automated [138, 200, 687, 844, 1277, 584, 1176, 974, 982, 1276, 984, 1087, 1201, 443, 249, 250, 329, 375, 376, 510, 855, 1256, 44, 846, 1238, 1248, 1046, 383, 795, 834, 697, 702, 426, 1003, 1378, 520, 973, 709, 1109, 435, 719, 964, 975, 1080, 30, 986, 1347, 1380, 437, 707, 188, 218, 1190, 432, 1245, 1336, 86, 109, 206, 1258, 745, 16, 642, 1246, 118, 888, 1313, 278, 351, 1060, 133, 353, 4, 52, 22, 57, 358, 2, 137, 247]. Automath [588]. Automatic [810, 1340, 909, 449, 775, 1041, 1051, 1074, 986, 1124, 1086, 471, 1024, 757, 33]. Automatically [1257, 1315, 1032, 216, 1033, 1185, 660]. Automating [1167, 683, 581, 997, 440]. Automation [1099, 953, 1312, 142, 51]. Automations [913]. Automotive [492]. Autonomy [1069]. avoid [221]. avoiding [212]. Award [618]. Aware [1136]. Axiom [824, 1128, 1311, 575, 541, 578]. Axiomatic [644]. axiomatizations [94, 218]. Axiomatizing [703]. Axioms [643, 1103, 1017, 695, 430, 573, 1293, 1307, 42, 334, 233, 241].

Bachmair [1206]. Back

[1208, 801, 442, 597]. Back-end [801]. Backchain [263]. Backdoor [778, 674]. Backend [803]. Background [593]. Backtracking [671, 23, 105]. Backward [754]. Balanced [678]. Balancing [787]. Barcelona [388]. Barriers [953]. Based [1134, 579, 824, 1081, 1237, 1366, 671, 1027, 1340, 657, 935, 1216, 1352, 562, 1173, 1349, 701, 1225, 611, 1202, 954, 947, 1054, 658, 752, 763, 848, 1347, 707, 1180, 980, 1220, 998, 1245, 1336, 1321, 594, 1187, 1021, 899, 518, 958, 1337, 349, 220, 3, 149, 205, 291, 730, 423, 192, 928, 721, 724, 299, 257, 323, 77, 313, 321, 254, 558, 680]. Bases [459, 147, 273]. Basic [43, 759, 314, 816, 37, 36]. Basin [686]. Basis [1196, 610, 1060, 104]. Bayesian [1034, 841, 820]. be [263]. Behaved [450]. behavior [216]. belief [291]. Bell [1238]. Benchmark [903, 497, 72]. Benchmarking [1377]. Bendix [831, 596]. Berlekamp [1191]. Bernays [206]. Bernstein [934]. Best [1298, 1285]. Beth [1275]. Between [1268, 176, 260, 137, 159, 145, 212]. Beyond [819, 163]. Bi [928, 1268]. Bi-Directional [1268]. Bi-intuitionistic [928]. Bidirectional [1079]. Big [729]. Binary [1309, 1306, 1215, 1174, 1198, 1323, 499, 78]. Binders [935, 968, 988, 1170]. Binding [893]. Bindings [1188, 1321]. Bisequent [1335]. Bisimulation [614]. Bit [707, 1253, 1255, 1149]. Bit-Level [707]. Bit-Vector [1149]. Bit-Vectors [1255, 1253]. Black [1193]. Bledsoe [378, 429]. Block [1003, 1070, 67]. Blockchain [1370]. Blocking [1172]. Board [110, 169, 428, 28]. Boogie [803, 1288]. Book [80, 255, 943, 639, 737, 87, 240, 281, 288, 686, 167, 888]. boolean [367, 160, 731, 557, 1148, 677, 1162, 694, 840, 544, 1324, 96, 573, 997, 778, 610, 703, 1337].

Borel [429]. born [200, 1]. Both [1239, 116]. Bottom [1172, 181]. Bottom-Up [1172, 181]. Bound [1051, 840, 1052, 1339]. boundaries [239]. boundary [216]. Bounded [835, 622, 1219, 266, 956, 82]. Bounded-overhead [266]. Bounding [1101, 1108]. Bounds [842, 673, 1333, 1055, 667, 501, 1026, 1024]. Boyer [235, 357, 343, 6]. Branch [840]. Branch-and-Bound [840]. Branching [600, 350]. Bridging [1169]. Bruijn [898, 581, 899]. Buchberger [519]. bugs [150]. Building [1228, 560, 647]. Built [535]. Built-in [535]. Bundy [686]. Bytecode [606, 603, 605, 602, 601, 623, 604].

C [1136, 798, 803, 926, 1098, 1030, 764, 784, 781]. C-like [764]. C11 [1265]. C11-Style [1265]. Cached [1199]. Caching [924, 1166, 266]. CADE [927, 1043, 1249, 277, 394, 397, 276, 403, 854, 385, 387, 402, 468, 500, 543, 612, 386, 439, 1182]. CADE-11 [277, 276]. CADE-13 [394, 397, 403, 385, 387, 402, 386]. CADE-14 [439]. CADE-15 [468]. CADE-16 [500]. CADE-17 [543]. CADE-18 [612]. CADE-22 [854]. CADE-23 [927]. CakeML [1211, 1140]. Calculi [1223, 941, 1002, 1345, 425, 730, 933, 946, 595, 316, 218, 190, 300]. Calculus [858, 969, 534, 892, 864, 29, 1251, 1023, 1137, 1351, 714, 1225, 430, 1335, 482, 720, 847, 483, 685, 1057, 1100, 1073, 1088, 703, 578, 234, 356, 286, 153, 13, 412, 222, 164, 361, 83]. Call [34, 161, 1137]. Call-by-Value [1137]. Cambridge [686, 888]. Canonical [941, 893, 68, 368]. Canonicity [1129]. Capable [629]. car [22]. car-sequencing [22]. Cardinals [655]. Carrying [620]. Cartan [862]. CAS [1381]. Case [909, 894, 501, 491, 848, 1303, 511, 358, 1058, 72, 651, 836, 52, 22, 274]. Categoricity [1138]. Categories [1356]. Category [1167]. Cauchy [1175]. causal [297]. cc [377]. CCS [790]. CD [430]. CDCL [1352]. CEGAR [1202]. CEGAR-Based [1202]. Centered [1258]. Central [1062]. Certificate [1183, 1364, 972]. Certificates [1363, 1104]. Certification [481, 1312, 107]. Certified [1361, 1244, 862, 840, 1367]. Certify [1121]. Certifying [952, 785, 604]. Chain [1215]. Chaining [754]. Chains [1061, 1384]. Challenge [898, 153, 999, 429, 540, 1281, 899, 294, 146]. Chance [1069]. Chance-Constrained [1069]. Change [53]. Changes [11]. changing [185]. characteristic [249, 250]. Characterization [754, 928, 634, 313]. Characterizations [425, 252]. Chase [931, 1289]. Check [1024]. Checked [519, 166]. Checker [621, 1067, 490, 1299, 246]. Checkers [1378, 972]. Checking [835, 606, 1366, 417, 479, 1212, 1218, 1163, 783, 1202, 1154, 808, 566, 1183, 774, 997, 904, 705, 870]. Checks [381]. Chief [28, 1229]. Choice [868, 1345, 799, 174]. Choices [646]. Choosing [70, 56, 93, 100, 137, 159, 197, 186, 212]. Choreographic [1319]. CHSH [1339]. Church [517]. CIC [807]. Cipher [1003]. Circles [961, 347]. Circuit [906]. Circuit-Level [906]. Circuits [1157, 336]. Circumscription [404, 18, 74, 248]. Citizens [753]. claims [133]. Class [1284, 460, 1239, 658, 753, 214, 297, 45]. Classes [644]. ’Classical [424, 558, 866, 360, 198]. Classification [1147, 815, 679, 432]. Classifiers [820]. Classroom [630]. Clausal [907, 647, 1020, 1243, 180]. Clause [1373, 813, 1073, 1088, 676, 176, 31, 306, 48, 92, 266, 273, 145]. Clause-Learning [813]. Clause-Sharing [1373]. Clauses [42, 1349, 749, 431, 1329, 220, 207, 221]. Claw [944]. Claw-Free [944]. Clean [853]. clear [67]. Clifford [435]. Clight [798]. CLIN [389, 390]. CLIN-E [389]. CLIN-S [390]. Close [998]. Closed [842, 120, 406, 511, 83]. Closed-Form [842]. Closed-World [406, 511]. Closure [617, 937, 1325]. CLP [229]. CNF [906, 405]. CNFs [569]. Co [730, 1045]. Coarsest [614]. CoCon [1234]. Code [1013, 987, 620, 1079, 782, 784, 781, 130]. Coded [633]. Codes [1204, 1323, 1310]. coefficients [104]. Coercive [533]. Coherent [750, 1287]. Coinductive [1037]. Collaborative [850]. Collatz [1313]. Collections [1105]. Collector [1140, 1372, 1141]. Colored [507]. Coloring [407]. Column [1282, 43]. Combination [1186, 1263, 1275, 1169, 883, 822, 1273]. combinations [32]. Combinatorial [833, 789, 1044]. combinatory [46, 335, 52, 196, 256]. Combined [858, 650]. Combining [447, 1173, 446, 563, 1318, 648, 1332, 663, 655, 1044, 365]. Commafree [1310]. Commitment [1240, 774, 822]. Common [1108]. commonsense [322, 337]. community [284]. Commutative [554, 711, 292, 230, 368, 61]. Commutativity [1214]. COMP [684, 921, 985]. Comparing [215, 719]. Comparison [677, 1124, 589, 61]. CompCert [1119]. CompCertS [1136]. Compendium [582]. Competing [270]. Competition [684, 394, 1064, 403, 385, 387, 402, 468, 500, 543, 564, 612, 386, 439, 278]. competitions [276]. Compilation [821, 977, 632, 761]. compile [4]. compile-time [4]. Compiler [1370, 1136, 803, 1082, 801, 1096, 604, 95]. Compiling [181, 396]. Complemented [992]. Complete [969, 808, 1057, 1295, 1375, 1021, 294, 195, 263, 117, 132, 124]. Completely [580]. Completeness [1037, 1056, 1186, 1251, 1015, 945, 204, 258, 705, 604, 371, 361, 31]. Completion [596, 562, 469, 923, 369]. completions [141]. Complex [1175, 1336]. Complexities [1049]. Complexity [1000, 1153, 1115, 1217, 1055, 1032, 474, 922, 675, 230, 1259, 1116, 929, 765, 1262, 1052, 452, 652, 237, 331]. Components [1161]. Composite [460]. Compositional [1094, 1161]. Comprehensive [926, 1279]. Compression [770]. Computation [1368, 380, 447, 1091, 1244, 1137, 526, 972, 980]. Computational [712, 1153, 845, 474, 438, 1295, 82]. Computationally [852]. Computations [952, 565, 977, 193, 895, 661]. Computer [844, 461, 460, 796, 1363, 1189, 520, 448, 696, 1310, 148, 1044, 195, 189]. Computer-Aided [520, 1310, 195]. Computer-Assisted [461, 1363, 189]. Computing [1091, 877, 404, 777, 748, 841, 1384, 405, 373]. Concatenation [1302]. concept [8]. conceptual [3]. Conclusions [403]. Concrete [1274, 725, 150]. concurrency [4]. Concurrent [1053, 936, 1141]. Condensed [195]. Condition [882]. Conditional [1085, 1222, 327, 251, 24]. Conditions [1214, 160, 122]. Cone [851]. Conference [434, 1234]. Confidentiality [1093, 850, 1234]. Confidentiality-Verified [1093]. Configuration [1246]. Conflict [1186, 1263, 1165, 1073, 1088]. Conflict-Driven [1186, 1263, 1165, 1073, 1088]. Confluence [1040]. Confluent [425, 45]. Conformance [598]. Congruence [617]. Conjecture [860, 1114, 1313, 1271]. Conjectures [421, 692, 1044]. conjunction [192]. Conjunctive [1337]. Connection [1233, 928, 755, 149, 316, 98]. Connection-based [928]. connections [119]. Conquer [1049]. Conquering [541]. consed [967]. Considered [405]. Consistency [492, 705]. Consistent [459, 1122]. Constant [1123, 1224, 189, 61]. constant-abstraction [61]. Constant-Time [1123, 1224]. Constants [1317, 1010, 880]. Constrained [833, 1069]. Constraint [472, 367, 412, 650, 489, 947, 1287, 458, 904, 792, 645, 136]. Constraints [872, 940, 1174, 520, 1165, 748, 1052, 236]. construct [14]. Constructing [1379, 364]. Construction [1326, 584, 1131, 1382, 520, 629, 757, 83]. Constructions [809, 482, 1138, 1310]. Constructive [1363, 1015, 1102, 816, 947, 1207, 811, 362]. constructs [173]. Contents [467, 515, 531, 552, 572, 592, 609, 627]. contest [277]. Context [633, 1259, 244]. Context-Sensitive [1259]. Contexts [534, 730, 1334, 646, 951, 99]. Continuations [800]. Continued [537]. Continuity [1103]. Continuous [582, 1160]. Contract [689]. Contraction [933, 71]. Contraction-Free [933]. contradictions [283]. contrapositives [317, 92]. Contribution [1282]. Control [973, 407, 78]. Controlled [316]. Controller [1305]. Convergent [877]. Converse [935]. Conversions [800]. Convex [1134, 1114]. Cooperation [593]. Coq [1227, 891, 1267, 1369, 1103, 967, 1340, 1099, 1102, 1195, 481, 789, 1137, 1355, 585, 901, 1311, 1026, 1331, 761, 961]. Corpus [950]. Correct [380, 1168, 962]. Correcting [1204, 416]. Correction [1367, 1237, 1297, 1298, 1266, 1307]. Correctness [29, 1098, 1115, 786, 851, 1080, 706, 166]. CoSMed [1093]. Cost [842, 1053, 1054, 1219]. Countable [1283]. Counter [933, 1236, 63]. Counter-Models [933]. Counting [1048, 1175]. Coverage [835]. Cox [1192]. Craig [1243]. Cramming [599]. Credits [1115]. criteria [29]. Critical [835, 831]. Cross [460]. Cryptanalysis [493]. CryptHOL [1240]. Cryptographic [1153, 845, 847]. Cryptography [1123, 1224]. Crystal [806]. CSP [774]. CSPs [881]. CTL [1015, 106]. Cubical [1128, 1130, 1129, 1356]. Current [532, 984, 1381]. Curve [789, 914]. Custom [913]. Cut [812, 1335, 1283, 316]. Cut-Free [812]. Cutting [931, 1269]. Cyber [1161, 1357]. Cyber-Physical [1161, 1357]. Cycle [732]. Cycles [562]. Cyclic [1325, 1185].

D [1222]. Dafny [1193]. Dally [882]. Damas [481]. Daniel [943]. Data [770, 459, 1119, 967, 472, 771, 1328, 1163, 492, 1124, 765, 1029]. Database [510]. Databases [409, 635, 777, 553, 290, 370, 272, 76, 251, 97, 308, 7, 307]. Datalog [743]. datastructure [25]. Datatypes [867, 1273, 1045]. David [686]. Davis [496]. Deadlock [882]. Deadlock-Free [882]. Dealing [881, 1170]. Debugging [742]. Decades [1276, 990]. Decidability [878, 883, 1015, 691, 462, 870, 1025, 175]. Decidable [1274, 1239, 1359, 647, 1235, 155]. Deciding [1272, 857, 694, 1104, 817]. Decimals [1091]. Decision [456, 981, 729, 1304, 681, 723, 1215, 143, 172, 173, 657, 935, 1359, 690, 1017, 174, 545, 558, 1252, 1061, 559, 740, 563, 957, 752, 763, 876, 724, 1207, 707, 968, 988, 980, 175, 1045, 1073, 1088, 651, 1036, 323, 32, 943]. Decisions [1342, 1162]. Declarative [805, 115]. Decoding [1079]. decomposition [135]. Decreasing [683, 869, 976]. Dedekind [1284]. Dedicated [836]. Deduction [664, 639, 982, 1201, 1046, 383, 1106, 1301, 401, 409, 707, 919, 432, 7, 652, 176, 349, 303, 314, 23, 202, 258, 206, 16, 311, 49, 145]. Deductive [510, 409, 635, 777, 553, 314, 370, 76, 251, 308, 338, 307]. Deep [1353, 1135, 1354]. Default [380, 821, 819, 338, 339, 214, 313, 341, 187]. defaults [326, 339]. Definability [1275, 18]. Defined [566, 1386, 63]. Definite [1113, 266]. definite-clause [266]. Definiteness [1024]. Definition [1111, 213, 71]. Definitional [995]. Definitions [1010, 1153, 460, 1328, 914, 814, 31].

Deletion [414]. demodulating [239]. Demodulation [1278, 223]. demodulator [231]. Dempster [822]. dense [204]. Dependencies [1109, 753]. Dependency [1145, 1352, 886, 856, 710, 1047, 1222, 1324, 929, 1126]. Dependent [1326, 1099, 1138]. Depth [933, 1301]. derivation [62, 4, 221]. Derivational [1259]. Derivations [878, 293, 181]. Derivatives [1192, 1322]. Description [739, 1274, 744, 885, 816, 815, 767, 954, 743, 839, 725, 867, 765, 432, 753, 741]. Design [684, 386, 704, 228, 150, 128]. Designing [76]. detachment [195]. Detect [504]. Detecting [693]. Detection [775, 834]. determining [117]. Deterministic [1305, 297]. Developing [456]. Development [1090, 521, 109]. Device [1094]. Diagnosis [423]. diagnostic [3]. Diagram [745]. Diagrams [1156, 1215, 869, 1070, 976]. Diameters [1101]. Dickson [522]. Dictionary [690]. Dieter [686]. Difference [1052, 792]. different [94]. Differential [435, 768, 1057, 433, 249, 250, 144, 298, 171, 1266]. Dimension [433]. Direct [1371]. Directed [957, 185, 192]. Directional [1268]. directions [70]. Disassembling [1013]. Disc [1226]. Disconnection [720]. DISCOUNT [391]. Discovering [122, 510]. Discovery [683, 1216, 795, 1245, 471, 578, 1060, 218]. Discrete [789, 772]. Discrimination [420, 245, 225, 224]. discrimination-tree [225]. Disequality [940]. disjunction [257]. Disjunctive [414, 425, 743, 526, 777, 553, 272, 257, 97, 308, 307]. disparate [85]. Disproving [421]. Distance [1126, 712]. distance-sat [712]. Distant [1091]. Distilling [1254]. Distributed [391, 973, 1109, 1373, 418, 293]. Distribution [1353]. Distributions [1384, 1305]. Distributivity [643, 539]. Divide [1049]. DL [744, 811]. DLL [674].

DNF [405]. do [1323]. Document [1234]. Domain [484, 1246, 362, 319, 252]. domain-minimal [252]. Domains [1353, 1274, 1284, 412, 1163, 725]. Dots [457]. Double [1378]. Down [987, 635, 1350, 181, 311]. DPLL [673, 817, 738]. DQBF [1145]. Driven [1186, 1263, 1165, 1073, 1088, 553, 158]. Drivers [1094]. Dual [548]. Dualities [260]. Duality [553]. Dubois [1197]. Dummett [545, 946, 724]. duplication [219]. during [336, 223]. DWFS [635]. Dynamic [885, 783, 947, 820, 768, 1057, 1266, 828, 829].

E-SETHEO [397]. E-Unification [422, 880]. Eager [732]. Earliest [1005]. Early [973]. Editor [28, 1229, 36, 39, 50]. Editor-in-Chief [28, 1229]. Editorial [43, 110, 139, 169, 1, 28]. Effect [1139]. Effective [656, 852, 165]. Effectively [817]. effectiveness [358]. efficiency [299]. Efficient [1164, 1306, 744, 238, 713, 1202, 1162, 1124, 1183, 1350, 722, 1298, 1285, 455, 504, 367]. Efficiently [1309, 841]. Eisbach [1012]. elaborating [280]. electrical [229]. Elegant [440]. Elementary [795, 554, 474, 1026, 471, 153, 143, 172, 173, 62, 174, 157, 36, 175, 177, 118, 37]. Elements [956]. Eleventh [434]. Eliciting [911]. Eliminating [383, 219]. Elimination [411, 414, 464, 1236, 1335, 675, 711, 732, 827, 431, 475, 315, 317, 156, 311]. ELK [959]. Elliptic [914]. Embedding [326, 1354]. Embeddings [577, 853]. Empirical [502, 215]. Empirically [697, 702]. employ [56]. Emptiness [940]. Enabling [1331]. Encoding [771, 890, 847, 1079]. Encodings [677]. Encryption [874, 846, 690]. End [1119, 801]. Engel [1316]. Engine [1355]. Engineering [434, 1292, 229]. Engines [1220, 1246]. Enhancements [1172]. enough [263]. Ensure [1072]. Entailment [1371, 222, 252, 821]. Environment [892, 638]. Epistemic [1301, 1143]. EQ [1320]. Equality [721, 1320, 837, 938, 442, 597, 9, 59, 211, 212, 278]. Equation [926, 1225, 1110, 790, 203]. Equational [523, 460, 877, 391, 401, 407, 876, 410, 647, 121]. Equations [729, 506, 1245, 433, 652, 185, 104, 269]. Equivalence [760, 151, 1076, 878, 1239, 887, 876, 790, 94]. Equivalential [430, 578]. Equivalents [477]. Equivariant [830]. Erdos [1114]. Errata [79, 88, 168, 302, 384]. Erratum [508, 561, 975, 988, 384, 1088]. Error [1204, 1338, 1024]. Error-Correcting [1204]. ETPS [631]. Euclidean [915]. Euler [1225, 745, 1058]. European [1192]. Evaluate [557]. Evaluating [559, 1175, 150]. Evaluation [31, 557, 985, 502, 1064, 406, 511, 299]. Evidence [1059, 996]. evident [1336]. Evolution [858]. Evolutionary [820]. Exact [1091, 116]. Example [417]. Examples [522, 63, 369]. Excerpts [1282]. Exchange [875]. Executable [1067, 181]. Execution [979, 1344]. Exercises [737]. existence [124]. Expansion [1117, 164, 71]. Expansions [995]. Expectation [772]. Expected [1384]. Experience [953]. Experiences [1261]. Experiment [447, 1007, 6]. Experimental [1167, 557]. Experimenting [242]. Experiments [179, 12, 225, 1180, 653, 704, 335, 57]. Expert [434, 3, 116, 17]. experts [150]. Explaining [1162, 696, 85]. Explanations [637, 841]. Explicit [892, 595, 63, 896]. Exploitation [202]. Exploring [315, 1203]. Exponent [695, 334]. Exponential [673, 634]. Exporting [1261]. Expression [746, 887]. Expressions [1026, 1322, 958, 24]. Expressive [935, 765, 741, 55, 305]. Expressiveness [1135, 136]. ExpTime [924, 1166]. Extended [1194, 1133, 1159, 1213, 927, 1107, 1078, 1359, 1043, 1249, 1205, 1188, 863, 691, 808, 1209, 968, 1066, 1270, 1321, 854, 1354, 1118, 1182, 172, 173, 23, 174, 125, 95]. Extending [932, 64]. Extensible [771, 1268]. Extension [1326, 727, 611, 632, 1184, 988, 504, 214, 298, 235, 285]. Extensional [1306, 1272, 1247]. Extensions [380, 867, 1289, 822]. exterior [147]. External [997, 586]. extracting [150]. Extraction [522, 1035, 1117, 919, 535].

Fact [1027]. Factorization [1191, 1196]. Factors [1309]. Failure [31, 475, 354]. Fair [875]. Fairness [786]. FALCON [364]. Falsification [1161]. Family [744, 885, 192]. Fast [1368, 633, 1114, 1384, 1289, 381, 245]. Faster [808, 1290]. Fault [1236]. Fault-Tolerant [1236]. FD [377]. Feedback [637]. Fermat [1358, 1058]. Few [1385]. Field [914]. Fields [1284, 1147, 1379, 2]. fifty [146]. file [346]. Filtering [472]. Find [1274, 1115]. Finder [1291]. Finding [1309, 1038, 52, 538, 176, 190, 78, 108, 145, 152, 165, 247]. Fine [1181, 1217, 978]. Fine-Grained [1217]. Fine-Structure [978]. Finitary [1334]. Finite [693, 941, 1147, 855, 412, 1174, 914, 520, 942, 679, 944, 1376, 904, 368, 364]. Finite-valued [941]. Finitely [562]. Finiteness [940]. Firewall [1095]. First [389, 1361, 1367, 684, 548, 779, 963, 1201, 390, 1256, 1174, 1383, 665, 84, 1025, 1351, 567, 635, 1178, 1311, 1320, 837, 1086, 996, 749, 1312, 568, 1100, 1073, 1088, 955, 418, 252, 1187, 753, 653, 1235, 1243, 450, 42, 235, 141, 155, 305]. First-Class [753]. First-Order [1361, 1367, 548, 963, 1201, 390, 1256, 1174, 1383, 665, 1025, 1351, 567, 635, 1311, 837, 1086, 996, 749, 1312, 568, 1100, 1073, 1088, 955, 418, 1235, 1243, 450, 84, 252, 42, 235, 141, 155]. Fitch [664]. Fitch-Style [664]. five [33]. Fixed [862, 1197]. Flat [981]. Flatterms [245]. Flaw [834]. flexibility [273]. Flexibly [563]. Floating [1309, 977, 1331, 661, 1255]. Floating-Point [1309, 977, 1331, 661, 1255]. Flow [1110, 1112, 1283, 784]. Flow-Sensitive [784]. Flowpipes [1160]. Floyd [1098, 133]. Flyspeck [964, 975]. FOC [587]. focusing [152]. Focussing [1208]. FOL [693]. Follow [1342]. ‘Fool [196]. foray [46]. Foreword [873]. Forget [1097]. Form [842, 1297, 1296, 789, 180, 307]. Formal [631, 902, 565, 1135, 1091, 1067, 833, 689, 861, 479, 1319, 773, 1109, 993, 478, 1030, 764, 457, 1114, 641, 859, 736, 1070, 998, 761, 1024, 706, 1071, 1199, 782, 910, 781, 882, 234, 209]. Formalisation [1150, 1050, 1241, 1011, 1356, 958]. Formalising [1240, 1300, 1293, 1307]. formalisms [326]. Formalization [960, 1204, 1134, 1284, 1267, 1297, 1296, 1339, 914, 831, 793, 1378, 1225, 1110, 482, 1151, 1077, 1158, 791, 1312, 934, 918, 1295, 1294, 1299, 1100, 971, 1226, 1365, 1142, 1260]. Formalizations [1318, 602, 920]. Formalized [789, 1188, 483]. Formalizing [894, 1192, 797, 595, 1112, 484, 1206, 1196]. Formally [1101, 1062, 1382, 1063, 1148, 1338, 954, 801, 1113, 972, 980, 1234]. Formally-Verified [1338, 980]. Formula [1181, 1333, 290, 1242, 490, 594]. Formula-Based [594]. Formulae [557, 1162]. Formulas [673, 865, 678, 731, 778, 149, 62, 267]. formulation [55]. Formulations [26]. Forward [754]. Found [542]. Foundation [714, 1122, 123]. Foundational [620]. Foundations [844, 1127, 279]. Founded [425, 526, 1171, 555]. Foundedness [477]. Four [990, 1335]. Four-Valued [1335]. Fourier [236]. Fourth [576]. FPGA [708]. Fragment [657, 935, 1025, 968, 988, 1187]. Fragments [1235, 1337, 193]. Framework [952, 881, 1097, 718, 600, 1059, 986, 840, 437, 1222, 707, 826, 1279, 1036, 641, 251]. Frameworks [456, 579, 532]. Free [1167, 864, 718, 1174, 1359, 933, 907, 812, 1184, 1080, 1207, 944, 882, 312, 94, 166, 848].

Free-Variable [718]. Fresh [890]. Front [1119]. Front-End [1119]. Full [482]. Fully [1041]. Function [814, 1070, 971, 159, 268]. Functional [1216, 396, 776, 1054, 1080, 1295, 925, 753, 164]. Functionality [1178]. Functions [1211, 810, 1267, 909, 991, 731, 1169, 382, 1351, 1189, 516, 616, 1026, 925, 651, 866, 82, 157, 345, 24, 279]. Fundamental [1142, 1050]. Fundamentals [140]. Fusion [892, 819]. Futile [656]. future [232]. Fuzzy [243, 205, 116].

G [581]. Gabriel [696]. Gale [1348].

Gamma [971]. Gandalf [392]. Ganzinger [1206]. Garbage [1140, 1372, 1141]. Gauge [1333]. Gazing [213]. GCWA [425]. General [523, 1349, 1188, 563, 725, 732, 880, 452, 191, 141, 300, 137]. general-purpose [137]. Generalising [945]. Generalization [1353, 1028, 166, 314]. Generalizations [421]. Generalized [120, 269]. Generalizing [1087]. Generate [455]. Generated [1008]. Generating [182]. Generation [1172, 505, 1125, 719, 1347, 375, 376]. Generational [1140]. Generator [1340, 130]. Generators [708]. Generic [1383, 1105, 641, 123]. Genetic [1074]. Gentzen [253]. Gentzen-type [253]. GeoGebra [1381, 984]. Geometric [520, 947, 1151, 1245, 1336, 1303, 1060, 375, 376, 357]. geometrie [36]. geometries [157, 171, 37]. Geometry [1038, 1039, 1340, 510, 795, 449, 947, 435, 809, 506, 1347, 629, 736, 471, 1226, 828, 829, 38, 62, 90, 249, 250, 329, 44, 144, 298, 109]. GHC [939]. Global [1284, 935, 940, 924, 934, 1166, 1375]. GMP [583]. Goal [1006, 1056, 1304, 1106, 957, 852, 553, 192]. Goal-Directed [957]. Goal-Driven [553]. Goal-Oriented [1106]. Goal-Sensitive [1006, 1056, 1304]. Gödel [460, 42, 724, 1327, 983, 1254, 86, 206].

GOEDEL [912]. Good [503]. GPU [1080]. Grained [1181, 1217]. Grammars [969, 1079]. GRAMY [629]. Graph [1195, 724, 306, 98]. Graph-based [724]. Graphical [736]. Graphs [944, 23, 48, 175]. Green [1150]. Gries [1265]. Gröbner [147, 1060]. Ground [1272, 982, 368]. Group [1379, 692, 148, 1142, 369, 218]. Groups [1284, 562, 695, 1189, 334, 94, 233, 1380, 241]. Guarantee [1053, 1141]. guaranteeing [132, 124]. Guarantees [1091]. Guarded [1130, 1230, 1337]. Guidance [1233, 686]. Guide [637]. Guided [1006, 1056, 1304, 390, 837, 494].

hacker [20]. Hammer [1099]. Handbook [888]. Handling [1256, 721]. hard [198, 998]. Hardest [972]. Hardest-to-Round [972]. Hardness [1084]. Hardware [487, 785, 113]. Harmful [405]. HARP [77]. Harrison [888]. Harrop [267]. Hash [967]. Hash-consed [967]. Heap [622]. Heap-Bounded [622]. heat [365]. Heavy [489]. Heavy-Tailed [489]. Hedges [949]. HeerHugo [490]. Heine [429]. Help [631]. Herbrand [618, 252]. Hereditarily [944]. hereditary [267]. HermiT [966]. Hessenberg [750]. Heuristic [1016, 671]. Heuristic-Based [671]. Hierarchical [511, 49]. Hierarchies [1170]. hierarchy [183]. High [401, 700, 290, 202, 208]. High-Performance [401, 290, 202, 208]. Higher [868, 1042, 1308, 1001, 865, 930, 1216, 894, 853, 1297, 1296, 999, 1236, 507, 1318, 814, 1011, 1111, 1139, 749, 1157, 708, 662, 979, 1247, 1280, 119, 265]. Higher-Order [868, 1042, 1308, 1001, 865, 930, 1216, 999, 1236, 507, 1318, 1011, 1111, 1139, 749, 1157, 708, 662, 1247, 1280, 894, 814, 119, 265]. Higman [859]. Hilbert [226, 576]. Hints [1385, 358]. Historical [1278]. history [178]. HL [968]. Hoare [133]. HOL [1150, 1211, 1368, 1050, 1010, 1167, 898, 737, 1027, 803, 1244, 771, 1265, 1049, 1192, 1339, 1300, 1351, 1225, 446, 915, 776, 1061, 1241, 1177, 1122, 1112, 1120, 1104, 1151, 1175, 1077, 1327, 480, 799, 1293, 1307, 1379, 971, 1346, 955, 1196, 762]. HOL-Boogie [803]. HOL4 [851]. holds [244]. Homogeneous [1189]. Homomorphic [874]. Homotopy [1131, 1127]. Honor [429]. honour [581]. Hopping [1358]. horn [5, 31, 359, 207, 121, 848, 163, 92, 320, 273, 1269, 7]. Horn-like [320]. Hot [451]. Huet [831]. Huffman [788]. Human [1189, 1041, 1258]. Human-Centered [1258]. Human-Style [1041]. Hutter [686]. Hybrid [838, 935, 871, 1264, 957, 988, 1166, 768, 1266, 1160]. Hyper [389, 509, 1020, 518, 574, 219]. Hyper-Linking [389, 509, 219]. Hyper-Linking-Based [518]. Hyper-Resolution [1020]. Hyperbase [499]. Hyperbolic [1226]. Hypergeometric [993]. Hyperidentities [499]. hyperlinking [299]. hyperlinking-based [299]. Hypermaps [789]. hyperparamodulation [295, 301]. Hyperresolution [505, 685, 85, 289]. Hypersequent [1325].

I-SATCHMO [546]. ICE [1216].

ICE-Based [1216]. ID [1257]. Ideal [598]. idempotence [40]. idempotent [41]. Identification [336]. identifying [216]. Identities [1272, 138]. Identity [1347, 1336, 372]. IEA [434]. IEA/AIE [434]. IEC [1070]. II [456, 1224, 1001, 1222, 342, 511]. III [944, 1247]. IJCAR [751, 1205, 863, 1270, 564]. Illustrated [1121]. ILTP [726]. Impact [1048]. Imperative [1052, 1120]. implement [234]. Implementation [533, 1115, 786, 1191, 1177, 1222, 791, 95, 519, 653, 704, 588, 8, 129]. implementations [321]. Implemented [800]. Implementing [967, 265, 698, 196, 496, 202]. Implicants [455, 405]. Implicants/ [405]. Implicate [722]. Implicates [405]. implication [192]. Implicit [911, 1153, 327, 18]. Importance [1246]. imprecise [228]. improve [136]. improved [318]. Improvement [546, 1074]. Improving [710, 299, 998]. IMPS [271]. Incarnations [394]. Incidence [1340, 13, 29, 361]. Incoherent [742]. Incomplete [1220]. Incompletely [1386]. Incompleteness [1145, 417, 1311, 983, 1254]. incompletensess [86]. Inconsistency [579, 227]. Inconsistent [459, 264, 283]. Incorrect [692]. increase [358]. Incredible [959]. Incremental [815, 598, 642, 299]. Incrementality [1097]. incrementally [280]. Indefinite [409]. Index [465, 466, 513, 514, 529, 550, 530, 551, 570, 590, 571, 591, 607, 625, 608, 626]. Indexing [633, 225]. Indices [898, 1175, 899]. Induction [699, 352, 1173, 516, 379, 342, 327, 179, 353, 355, 282, 351]. Inductive [730, 1328, 856, 482, 860, 453, 454, 164, 692, 336, 133, 354]. Industrial [835, 434]. Inequalities [1016, 217]. Inequality [1339, 122]. inevitability [264]. inexact [205, 116]. Inference [523, 865, 1056, 855, 481, 567, 996, 99, 480, 418, 538, 273, 450, 280, 75, 209, 263, 56, 108, 152, 197, 223]. Inferences [857, 69]. Inferring [579, 747, 97]. Infinite [1163, 663, 870]. Infinite-State [870]. Infinitely [1210]. Infiniteness [1332]. Influence [851]. Information [507, 97, 51]. Informational [426]. Infrastructure [802, 1065]. Inheritance [587, 1168, 48]. Inhibitor [528]. Initial [704]. Innermost [929]. Input [945]. Inputs [1230]. Instance [389, 1362, 767, 719]. Instances [738]. Instantiation [879, 567, 1021]. Instantiation-Based [1021]. instantiations [280]. Instantion [393]. Instruction [1241]. Integer [1136, 104]. integral [210]. Integrals [1113]. Integrated [700]. Integrating [735, 1265, 806, 448]. Integration [1326, 1267, 681, 316, 333]. Integrations [594]. Integrity [700]. Intelligence [434]. Intelligent [737, 23, 637, 30]. intelligible [1008]. Intensional [1248, 646, 328]. Interaction [734, 639, 357]. Interactions [852]. Interactive [902, 1089, 1090, 803, 1369, 1357, 1009, 271, 1194]. Interconnection [882]. Interest [158]. Interface [1268, 736]. Interfaces [733]. interference [1123, 1224]. Interleavings [1053]. International [434]. Interpolants [1251, 1275]. Interpolating [864]. Interpolation [1366, 974, 982, 1022, 1335, 1019, 1020, 1021, 1243]. Interpretation [1013, 1046, 362, 311]. Interpretations [1121, 669, 756, 1286]. Interpreting [257]. Interprocedural [784]. Interruptible [1094]. Interrupts [785, 324]. Intertwined [1189]. interviewing [150]. Interweaving [150]. Intractable [105]. Introducing [383, 413]. Introduction [1089, 170, 188, 284, 90]. Intuitionistic [1103, 789, 933, 1106, 1349, 1301, 726, 442, 597, 193, 928, 318]. Invariant [907]. Invariant-Free [907]. Invariants [1220, 1160, 757, 375, 376]. Inverse [754, 667, 103]. investigation [189]. Investigations [1360]. invocation [185]. Involving [651]. iptables [1095]. IR [1082]. Ireland [686]. Isabelle [1150, 1002, 898, 737, 1027, 1244, 1265, 1049, 1192, 1339, 1300, 1351, 1061, 1241, 1177, 1144, 1122, 1112, 1175, 1012, 480, 242, 983, 1299, 1293, 1307, 1379, 1346, 955, 1196, 762, 1104]. Isabelle/ [1104]. Isabelle/HOL [1150, 898, 737, 1027, 1244, 1265, 1049, 1192, 1339, 1300, 1351, 1061, 1241, 1177, 1122, 1112, 1175, 480, 1293, 1307, 1379, 1346, 955, 1196, 762]. Isar [1008, 589]. IsaVODEs [1357]. Isomorphisms [1195]. Isotopy [757]. Issue [733, 739, 844, 1176, 927, 1152, 873, 697, 702, 1249, 1205, 709, 863, 1051, 1019, 1127, 1270, 854, 818, 322, 337, 309, 351]. Issues [691, 258]. Iterated [425, 657]. iteration [263]. ITP [1133, 1107, 1118]. ITPs [1368].

James [1131]. Java [603, 808, 602, 1096, 601, 604]. Java-to-JVM [604]. JML [836]. Job [851, 22]. Job-shop [22]. John [888]. Jordan [789]. journal [27, 35, 1]. Justification [822]. JVM [604].

Kan [1289]. KBO [698, 794]. Keller [1271]. kept [178]. Kernel [950, 256]. Kernels [1094]. KeY [962]. Keys [753]. Kit [131]. Kluwer [346]. Know [422, 698]. Knowledge [459, 877, 883, 632, 326, 3, 228, 150, 155, 273]. knowledge-based [3]. Knuth [596, 831]. Koala [1304]. Kroening [943]. KT [497]. Kutta [1338].

Labelings [976]. Labelled [941, 1020, 314]. Lagrange [1225]. lambda [234, 300, 699, 534, 1137, 646, 483]. Lambdas [1250]. Language [622, 798, 1034, 412, 806, 1102, 603, 808, 809, 457, 1158, 1012, 587, 1295, 265, 129, 377]. Languages [902, 682, 1359, 1054, 715, 1157]. LaPadula [1238]. Large [1138]. Larry [1277, 1276, 1282]. Last [578]. Lattice [499]. Lattices [582, 992]. Layout [708]. Lazy [1252, 1054, 938, 755]. Leaf [652]. lean [349, 1362, 1268]. Learn [1097]. Learning [1164, 1353, 1135, 1027, 963, 391, 1161, 1233, 1377, 1232, 1162, 964, 1220, 648, 820, 813, 503, 1376, 1073, 1088, 975]. Learning-Assisted [964, 975]. Learning-Based [1027, 1220]. Least [1050, 1108, 822]. Lebesgue [1267]. Left [1289, 218]. Legacy [537]. Legibility [998]. Lemma [759, 522, 1300, 859, 919, 213, 187]. Lemmas [411, 1263, 1125]. Lemmata [413]. Length [1189, 792]. Leo [1001, 1247].

Leo-II [1001]. Leo-III [1247]. Lessons [1369]. Level [682, 1123, 1224, 785, 895, 906, 707, 686, 782, 1149, 651, 1013, 75]. lexicographical [135]. Lexing [1322]. LF [533]. LIB [985]. Liberalized [717, 312]. Libraries [1261]. Library [1204, 1362, 1090, 916, 726, 441, 802, 1065]. Lie [1379]. Light [1225, 1327, 915, 1151, 1077]. Lightweight [623]. like [764, 320]. LIM [294]. Limit [1062]. Limited [1178]. limits [1381]. Line [705]. Linear [1156, 1204, 1050, 729, 359, 1042, 872, 1075, 766, 937, 933, 1301, 1343, 634, 1155, 1158, 1312, 827, 1346, 1187, 1235, 1290, 204, 353, 104, 236, 1364, 320, 296]. Linearity [1152]. Linearization [1121]. Link [393, 48]. Linked [538, 209]. Linking [389, 509, 518, 219]. Links [381]. LINUS [393]. lion [15, 125]. List [903, 451]. List-Machine [903]. Lists [825]. Lite [744]. Literal [675, 239]. Literals [1073, 1088, 366]. Living [639]. LK [1028]. LLL [1196]. LLVM [1350]. Load [787]. Löb [1327, 86]. Local [677, 491, 502, 1111, 841, 494, 1285, 1020, 676, 99, 1298]. Locales [1203, 948]. Localization [696]. Locally [900, 569, 893]. Logic [487, 1156, 1085, 424, 1223, 380, 680, 868, 1308, 1167, 750, 548, 577, 1103, 825, 1075, 723, 416, 833, 1098, 935, 1115, 367, 1325, 1296, 1328, 1371, 816, 933, 1106, 1210, 1349, 545, 946, 1301, 426, 1351, 895, 1343, 767, 954, 567, 1221, 1287, 1069, 1144, 1318, 521, 839, 814, 1030, 634, 1011, 1111, 1320, 715, 1139, 1327, 867, 1158, 988, 1298, 1285, 432, 1341, 708, 647, 458, 662, 817, 768, 1057, 1266, 726, 925, 333, 1100, 904, 442, 597, 1354, 888, 866, 325, 119, 363, 339, 214, 260, 194, 304, 42, 82, 55, 193, 181]. logic [150, 373, 265, 310, 46, 335, 227, 343, 141, 257, 52, 285, 196, 318, 195, 267, 20, 164, 155, 32, 92, 320, 115, 328, 313, 136, 308, 296, 106, 243, 146, 137, 256, 248, 305, 1297]. Logical [532, 881, 754, 443, 1086, 493]. Logics [838, 739, 1274, 497, 1345, 744, 885, 558, 815, 539, 843, 743, 1335, 724, 437, 725, 996, 498, 568, 765, 462, 753, 741, 287, 214, 191, 314, 348, 188, 341]. Long [1126]. Long-Distance [1126]. Loop [1223, 1109]. Loop-Type [1223]. Loops [1053]. Lorenz [1092]. LOTOS [594]. Low [682, 1013, 785, 782]. Low-Level [785, 782, 1013]. Lower [673, 1333, 1055]. LTI [1230]. LTL [1385, 1067]. Lua [1292]. Łukasiewicz [1210, 904, 575]. Lyapunov [1363].

M [681]. Machine [903, 1008, 963, 987, 1161, 1233, 470, 1079, 519, 234]. Machine-Checked [519]. Machine-Code [1079]. Machine-Generated [1008]. Machines [707]. Major [1261]. make [160]. Making [1280]. MaLeS [986]. MALL [1208]. Management [492, 1234, 228]. Managing [507]. Manuals [705]. Many [1076, 55]. Many-Sorted [1076]. Map [1110, 173]. Maple [446]. mapping [176, 145]. Maps [1190]. Mark [1005, 1004]. Markov [1061, 1384]. Match [667, 280]. Match-Bounds [667]. Matching [554, 474, 68, 238, 331]. Mathematica [1268]. Mathematical [1362, 948, 1090, 699, 577, 681, 916, 686, 351, 271]. Mathematicians [910]. Mathematics [950, 524, 714, 581, 438, 638, 910]. Mathlib [1316]. MathsTiles [737]. Matijasevich [187]. Matita [734, 902]. Matrix [1083, 756, 634, 205, 190]. Matroid [1340]. Matroid-Based [1340]. Maude [1085]. Max [1283, 494]. McCune [224]. MDP [1219]. mean [357]. Means [594, 23]. Measure [1326]. Measures [683]. Measuring [1303]. Mechanical [506, 107, 758, 598, 340, 171, 83, 62, 90, 157, 36, 201, 19, 37]. Mechanically [669, 1372, 166, 129, 130]. mechanically-checked [166]. mechanics [249, 250, 171]. Mechanisation [1231]. Mechanised [983]. Mechanising [1096, 1327]. mechanism [13].

Mechanization [750, 485, 668]. Mechanized [903, 759, 798, 926, 1103, 524, 521, 1283, 1146, 939]. Mechanizing [788, 1014, 710, 992, 581, 951, 374]. Media [1093]. Meeting [146]. Meets [1145, 1201, 1343]. Membership [657]. Memoriam [989]. Memory [1224, 1004, 1136, 1119, 1276, 1212, 1265, 783, 1241, 1030, 764, 1033, 782, 784]. Memory-Aware [1136]. Meredith [541]. Mergesort [939]. Message [1229]. Meta [75, 686, 311]. meta-interpretation [311]. Meta-Level [686, 75]. MetaCoq [1200]. Metalanguages [532]. Metalogic [1299]. Metalogical [456]. metamathematics [19]. Metatheory [903, 902, 1383, 1158, 1146, 951]. METEOR [315]. Method [1134, 497, 754, 435, 379, 884, 1012, 1303, 496, 1060, 1105, 149, 62, 90, 249, 250, 348, 103, 323, 263, 273]. Methodological [578]. Methodology [1141, 341]. Methods [950, 1172, 1037, 1169, 845, 1338, 650, 567, 719, 1162, 919, 133, 61]. MetiTarski [810]. Metric [1221, 1077]. Microkernel [786]. Microprocessor [128]. Middle [355]. Middle-out [355]. Milawa [987]. Milestone [536]. Milestones [1089]. Milner [481]. Min [1283]. Minds [1358]. Minimal [505, 933, 777, 748, 203, 222, 252]. Minimality [504]. Minimization [1028, 74, 319]. miniscoped [320]. Minkowski [1293, 1307]. Minors [1195]. Mirrored [944]. Missing [542]. mixed [181]. MIZAR [582, 994, 911, 1090, 735, 913, 914, 990, 992, 916, 1144, 997, 920, 589, 917, 995]. ML [481]. Mobile [624]. Modal [497, 558, 559, 560, 498, 568, 1298, 1285, 462, 1262, 287, 134, 191, 84]. Modalities [935]. Modality [324]. Modalized [905]. Model [835, 459, 411, 693, 1274, 1123, 606, 317, 414, 858, 1172, 1119, 1366, 1006, 549, 1067, 505, 1251, 479, 1212, 1218, 370, 464, 1137, 1163, 423, 783, 650, 1202, 1154, 843, 74, 1241, 1138, 1030, 1048, 764, 1096, 1291, 774, 1295, 1226, 1246, 475, 870, 315, 291, 156, 311, 307, 175, 196]. Model-based [423, 291]. Model-Checking [1163]. Model-Theoretic [459, 650]. Modeling [808]. Models [771, 731, 1230, 1265, 933, 773, 560, 1086, 647, 782, 190, 182]. Modern [791, 813]. Modes [1003]. modified [285]. Modifying [594, 1013]. Modular [642]. Modularity [1084]. Module [948, 72, 114]. Modules [1263]. Modulo [1085, 643, 874, 684, 872, 1201, 613, 879, 938, 1253, 1179, 1330, 1255, 1007, 652]. Monadic [1211, 1187]. Monads [1132, 1383]. Monotone [651]. Monotonic [908]. Monotonicity [865]. Moore [235, 357, 343, 6]. Morley [1087]. Most [1114, 841]. Mostly [1109]. Motivated [417]. Motivation [653]. Moufang [138, 372]. Moves [761]. Moving [1071]. MPTP [653, 704]. MSP [821]. Multi [849, 1219, 923]. Multi-Attacker [849]. Multi-Completion [923]. Multi-cost [1219]. Multicast [773]. Multilevel [172, 173, 174]. Multiparty [689]. Multiple [469, 437]. Multiple-Valued [437]. Multiples [1108]. Multiplicative [1156, 634, 651]. multisets [25]. Multithreaded [1096]. Mutation [1164]. Mutually [408]. Myhill [958].

N. [581]. Nagging [418]. Naive [820]. Named [893]. Nameless [900]. Names [890, 896]. Naming [889, 268]. Narrowing [1085]. Natural [664, 1218, 383, 1106, 919, 176, 145]. natural-deduction [176, 145]. Naturals [1221]. Near [163, 160]. near-boolean [160]. Near-Horn [163]. Necessary [882]. Negation [424, 31]. negative [104, 97]. Neighborhood [723]. Nerode [958]. Nested [408, 945, 901, 814, 331]. nested-object [331]. nets [245]. Network [1014, 1112, 299]. Networks [1063, 773, 1283, 841, 882]. Neumann [206]. Next [999]. NFM [1159, 1066]. Nice [732]. Nicolog [377]. No [503]. No-Good [503]. Nominal [983, 970, 762]. Nominals [1170]. Non [1317, 1123, 1224, 908, 937, 677, 104, 1364, 5, 1341, 92, 207, 1380, 297, 198, 7]. non-classical [198]. non-deterministic [297]. Non-horn [5, 92, 207, 7]. Non-interference [1123, 1224]. Non-Linear [937, 1364]. Non-Monotonic [908]. Non-negative [104]. non-orderability [1380]. Non-termination [1341]. Non-trivial [1317]. nonassociative [57]. nonclassical [188]. Nonconstructive [438]. nonlinear [216]. Nonmonotonic [424, 596, 325, 260, 322, 337, 323, 184]. Nonnegative [1267]. Nonstably [663]. Nonstandard [547]. Nonsystematic [503]. Nontotal [596]. Normal [1309, 1297, 1296, 180, 307]. Normalization [937, 595, 24]. Normalized [1085]. notation [159]. Note [112, 224, 68, 203, 237, 345, 39, 50]. notion [113, 300]. Notions [883]. NP [998]. NQTHM [485]. NQTHM-92 [485]. NRF [258]. Nuclear [1070]. nuclei [301]. nullary [40]. Nulls [1294]. Number [461, 797, 843, 1336, 1365, 1058, 203, 177]. Numbers [1025, 1177, 1175, 1255]. Numeral [1155]. Numerical [926, 867, 644].

O [729]. OBDD [284]. Obfuscation [1072]. Obituary [640]. Object [771, 730, 1144, 331]. Object-based [730]. Object-oriented [771]. obstacles [51]. Obvious [69]. OCaml [1372]. Octopus [648]. odd [334]. ODE [1092]. ODEs [1110]. Ofer [943]. Off [1338]. Offline [690]. OMDoc [916]. Omission [911]. On-Line [705]. Once [1269]. one [166]. Only [777]. Ontologies [815]. Ontology [1034]. Open [624, 540, 114]. Opening [101]. OpenJDK [1105]. Operating [783, 780, 782].

Operating-System [782]. Operation [1003]. Operational [1171]. Operations [997, 241]. operator [243]. operators [172, 174, 230, 30]. Optimal [723, 472, 792, 1305]. Optimality [1063]. OptiMathSAT [1179]. Optimization [840, 453, 934, 1179, 1255, 122]. Optimized [559]. Optimizing [180, 741]. Order [838, 1361, 1367, 868, 1042, 1308, 1001, 548, 779, 865, 1342, 1028, 963, 930, 1067, 1201, 1216, 479, 390, 1256, 1174, 853, 1297, 1296, 1383, 999, 665, 1025, 1351, 1236, 567, 507, 635, 1318, 1178, 1138, 1311, 1011, 1111, 1320, 1139, 837, 1086, 996, 749, 1312, 568, 1157, 708, 662, 979, 1100, 1073, 1088, 955, 1247, 418, 1187, 1235, 1280, 1243, 450, 119, 42, 894, 265, 84, 235, 814, 141, 155, 252, 305]. Order-Sorted [1086]. orderability [1380]. Ordered [509, 1206, 307, 574]. ordering [330, 135]. Orderings [596, 908, 469, 135, 259]. Orders [886, 352, 204]. Ordinal [461, 668]. ORE [1064]. Organizing [644, 150]. Orientability [794]. Oriented [1106, 771, 341]. Orthogonal [1040]. Other [1172, 1087, 539, 540, 135, 32]. Otter [699, 394, 335, 372]. Otter- [394]. Out-of-Distribution [1353]. Output [1041]. Overbeek [278]. overhead [266]. overview [2]. Owicki [1265]. OWL [966, 1084, 1064].

PACE [905]. Pair [831, 1222]. Pairs [991, 886, 856, 710, 1047, 929]. Papers [1194, 1133, 1159, 1213, 927, 1107, 1078, 1276, 1043, 1249, 1205, 863, 1066, 1270, 854, 1118, 1182, 34, 161]. Pappu [1039]. Paraconsistent [427]. Parallel [1103, 1350, 648, 761, 399, 207, 238, 202, 237]. Parallelism [463]. Parallelization [303, 658, 518]. parameter [215]. Parameterized [1212]. Parameters [1362]. Parametric [1253]. Paramodulation [596, 908, 1278, 755, 1247, 290, 121, 12, 85, 78].

Pardinus [1291]. Part [1123, 1224, 314, 192, 1222, 511]. Parthenon [207]. Partial [1067, 479, 1218, 382, 516, 567, 814, 616, 624, 925, 866, 82]. Partial-Order [479]. Partially [1210, 292]. Partition [614]. Path [886, 330, 225]. Pattern [1042, 185]. Pattern-directed [185]. Patterns [746]. PC [485]. PDL [957, 1166]. Pearl [788, 1108, 1083, 887, 1139, 859, 944, 939, 882, 962]. Perception [1069]. Performance [776, 401, 290, 202, 208, 85]. Performant [1355]. Permissions [1315]. Permutations [523]. Permutative [652]. Permuting [410]. Persistence [1160]. personal [276]. Perspective [1278]. Perturbation [759]. perturbed [216]. Phenomena [489]. Phylogenetic [747]. Physical [1161, 1357]. pigeonhole [149]. Ping [691]. Pinpointing [824]. Pipelined [707]. Plane [1269, 828, 829]. Planning [973, 850, 448, 1246, 645, 210, 338]. plans [179, 8, 67]. Platform [1093]. plausible [283]. Plus [643]. Poincaré [1110, 1226]. Point [1309, 943, 977, 862, 1197, 1347, 715, 1331, 661, 1255, 118]. point-set [118]. Pointer [1136, 1119, 1033, 1185]. Points [1114]. Pointwise [1221]. Policies [1072]. Polite [1076, 1273]. Politeness [1169, 1332]. Poly [1165]. Poly-Power [1165]. Polyadic [847]. Polygons [1114]. Polyhedra [1134, 1121, 472]. polymodal [348]. Polymorphism [1139]. Polynomial [872, 669, 843, 1104, 980]. Polynomials [436, 934]. Pong [691]. PoplMark [898, 899]. Portfolio [1149]. Portfolios [841]. Positive [414, 505, 1024, 156]. Positiveness [436]. POSIX [1322]. possibilities [1381]. Posthumous [1282]. Postulates [1103]. Power [861, 843, 1165, 1190, 365, 305]. powerful [377]. Powerlist [793]. powerset [172]. Practical [1359, 1209, 888, 263]. Practically [1350]. Practice [1201, 665]. Precise [1214]. Predicate [1264, 20, 318, 159]. Predicates [1235, 212]. Predicting [693]. Preemptive [785]. Preface [419, 476, 823, 733, 769, 1194, 739, 1159, 844, 716, 1213, 1176, 927, 1107, 804, 1152, 1031, 619, 1043, 889, 1249, 751, 1205, 659, 666, 863, 1051, 581, 445, 628, 636, 649, 1019, 322, 337, 1127, 1066, 897, 854, 309, 818, 351, 1118, 1182, 36]. Preference [579, 819]. Preference-Based [579]. Prefix [1342]. Premise [950]. Preprocessing [1035, 1048, 1329]. prerequisites [339]. Presburger [864, 694]. presence [166]. Presentation [920, 828, 829]. Presented [562]. Preserve [1323]. Preserving [1022]. Presheaf [892]. Press [686, 888]. preventing [150]. Pricing [1192]. Primal [969]. Prime [914, 797, 843, 405, 455, 1365]. Primitive [1355]. Primitives [847]. Primitivity [1323]. Principle [998, 822, 209]. principles [36, 37]. prior [254]. Priorities [339]. Priority [1168]. Privacy [1072]. Probabilistic [690, 1210, 379, 1071, 13]. probabilities [251]. Probability [1283]. Probable [841]. Probably [1210]. Problem [1050, 1272, 1148, 947, 409, 1364, 493, 415, 429, 726, 656, 441, 802, 1065, 576, 488, 176, 213, 1381, 68, 133, 226, 285, 22, 26, 56, 71, 64, 51, 85, 78, 93, 100, 117, 108, 132, 124, 137, 159, 145, 152, 197, 186, 165, 212, 231, 239, 223, 217, 247, 282, 268, 261, 274, 254, 295, 301, 289, 372]. Problems [930, 713, 999, 677, 614, 489, 474, 1104, 1347, 494, 384, 540, 43, 134, 153, 122, 294, 229, 230, 5, 33, 177, 216, 198, 114, 105, 278]. Procedural [807, 115]. Procedure [456, 411, 729, 723, 657, 935, 562, 1359, 1328, 1371, 545, 1252, 740, 1165, 957, 752, 763, 968, 988, 1045, 651, 144, 298, 285, 98, 267, 32, 311, 83]. Procedures [981, 943, 1304, 681, 690, 1017, 558, 559, 563, 707, 980, 387, 1036, 143, 172, 173, 174, 175]. Process [31, 1014, 790]. Processes [1061]. Processing [1181, 635, 722, 553, 202].

Processors [1222]. Producing

[1211, 1370, 1373]. Product [492, 705]. production [329, 297]. Productive [354]. Program [912, 522, 926, 1032, 1068, 1082, 764, 453, 758, 535, 1344, 136, 358]. Programming [424, 380, 902, 577, 747, 367, 1319, 701, 1074, 715, 1158, 1341, 458, 184, 488, 325, 363, 260, 194, 304, 265, 310, 92, 137, 248]. Programs [1317, 825, 416, 1098, 1216, 443, 785, 1080, 1288, 662, 936, 979, 925, 1052, 1033, 1185, 1386, 1354, 31, 214, 151, 133, 14, 181, 373, 141, 257, 164, 115, 297, 308]. Progress [875]. Progressions [1300]. Project [1200]. Projection [1218, 1375]. Projective [1340]. PROLOG [125, 106, 151, 44, 163, 95]. Promela [775]. Proof [1211, 1156, 770, 621, 759, 734, 1062, 1370, 1090, 1135, 583, 584, 750, 737, 788, 926, 1028, 1039, 1340, 417, 766, 1108, 1059, 1352, 1022, 862, 805, 1238, 789, 1328, 1106, 1349, 1355, 807, 620, 797, 1154, 585, 1184, 478, 448, 1209, 1074, 1261, 887, 1207, 1347, 1139, 1283, 837, 1324, 453, 1356, 1114, 1331, 859, 1012, 997, 586, 826, 983, 1193, 1254, 1299, 1258, 1262, 638, 939, 296, 1185, 882, 540, 442, 597, 1354, 1360, 645, 962, 210, 179, 113, 271, 354, 166, 285, 195, 98, 267, 6, 201, 83, 944]. Proof-Carrying [620]. Proof-Guided [837]. Proof-Producing [1211, 1370]. Proof-Search [1106, 1207, 597]. Proof-Theoretic [1352]. Proofs [911, 631, 664, 1153, 1181, 565, 832, 1038, 461, 460, 1091, 1228, 1008, 1037, 982, 1263, 886, 846, 1363, 542, 1189, 1003, 667, 516, 709, 993, 413, 595, 1373, 736, 517, 919, 998, 587, 1303, 979, 1024, 1020, 575, 1344, 642, 538, 1269, 440, 828, 829, 138, 149, 329, 375, 376, 14, 86, 347, 148, 278]. Propagation [1329, 1149]. Properties [981, 678, 1239, 1238, 566, 662, 925, 1185, 336]. Property [911, 933, 946, 1301]. Propositional [497, 671, 723, 933, 1106, 1349, 545, 946, 701, 490, 559, 1048, 568, 817, 462, 475, 703, 286, 151, 193, 215, 32, 190]. Protected [141]. Protocol [687, 849, 776, 848, 1168]. Protocols [688, 1014, 1240, 877, 883, 1239, 795, 690, 691, 692]. Prototype [1257, 1315, 1355, 1040]. Provability [1327]. Prove [1232, 144, 298]. Proved [851, 660, 1168, 263]. Prover [810, 902, 1016, 1001, 803, 1340, 909, 390, 987, 391, 1351, 1041, 527, 235, 393, 629, 1170, 1180, 395, 1206, 399, 1187, 518, 210, 207, 82, 191, 204, 357, 299, 208, 319, 77, 123, 6, 57, 95, 1304]. Prover-Backend [803]. Provers [986, 192, 369, 33, 321, 641]. Provide [853]. Proving [733, 1089, 1194, 447, 857, 974, 1369, 984, 963, 930, 38, 510, 669, 1063, 383, 786, 421, 449, 613, 1049, 1173, 756, 856, 569, 875, 611, 772, 435, 1221, 1287, 94, 1009, 492, 506, 1171, 637, 1026, 454, 65, 24, 685, 1336, 745, 1033, 1279, 72, 187, 450, 1060, 287, 360, 315, 133, 90, 44, 294, 157, 330, 353, 4, 1380, 277, 52, 36, 328, 313, 341, 266, 311, 171, 37]. Pruning [418, 475]. PS [905]. Psi [1002]. Psi-Calculi [1002]. PSpace [1143]. PTTP [317]. Publisher [73, 89, 102, 154, 162, 199]. Purchase [688]. Pure [675, 347]. purpose [137]. Putnam [496]. Putting [259]. puzzle [125]. Puzzles [428]. PVS [1070, 1260]. Pythagorean [1148].

QBF [1228, 1352, 1035, 1018]. QBFs [1208]. QCDCL [1333, 1342]. QPCF [1157]. QSMA [1374]. Quadratic [706, 201]. Qualitative [512]. Quantification [1337, 82, 235]. Quantified [731, 557, 1125, 778, 140]. Quantifier [864, 1174, 1359, 1236, 827, 1018]. Quantifier-Free [864, 1174, 1359]. Quantifiers [1374, 1359, 383]. Quantitative [825, 600]. Quantum [1244, 1158, 1157]. Quasi [1025, 1335]. Quasi-decidability [1025]. Quasi-Relevant [1335]. Queries [806, 406, 511]. Query [31, 744, 635, 722, 765, 553, 1337, 341]. Question [578]. Questioning [656]. Questions [540]. Quipper [1158].

R [888, 727, 229]. Race [270, 1080, 101]. Race-Free [1080]. Ramsey [343]. Random [678, 713, 1198, 772]. Rational [1235]. RCD [430]. Reachability [1239, 654, 947]. Reached [536]. Reactive [1385]. Read [1269]. Read-Once [1269]. Readability [1303]. readable [375, 376]. Real [810, 1016, 909, 449, 1025, 1070]. Real-Time [1070]. Real-Valued [810, 909]. Realistic [522]. Realization [1087]. Reasoner [966, 1064]. Reasoners [954, 593]. Reasoning [424, 200, 687, 739, 844, 1277, 427, 1176, 1006, 1056, 1276, 1304, 796, 967, 1201, 744, 885, 600, 730, 1256, 1248, 1023, 999, 697, 702, 426, 895, 525, 743, 407, 719, 964, 975, 819, 1082, 437, 9, 1084, 1190, 458, 1029, 1330, 512, 970, 1199, 852, 593, 741, 1386, 818, 325, 260, 13, 29, 205, 133, 249, 250, 44, 215, 227, 355, 116, 322, 337, 1380, 188, 323, 22, 158, 221, 184, 338, 358, 72, 114, 118, 2, 51, 137, 217, 261, 274, 686, 888]. Reasons [942]. Recalled [537]. received [80]. Reciprocity [201]. Recognizing [220]. Reconfigurable [487]. Reconsidering [991]. Recovering [459]. Recursion [1321, 342]. Recursive [443, 886, 408, 814, 453, 470, 31, 135]. Red [1193]. Reduced [722]. Reduces [475]. Reducible [1060]. Reducing [930, 876, 848]. Reduction [1067, 1215, 479, 404, 746, 851, 435, 743, 469, 646, 662, 300, 1196, 285, 98]. Reductions [1298, 1285, 117, 132, 124]. Redundancy [1207, 475]. Redundancy-Free [1207]. Redundant [1329, 221, 51]. Reference [998]. Refinement [414, 1121, 1216, 1124, 1112, 1120, 1350, 811, 1187, 156]. Refinements [464, 1180, 936]. Reflected [456]. Reflection [766, 585, 175]. Reflective [987]. Reflexive [940].

Refutation [945, 318]. Refutations [1242, 792, 359]. Refuting [692]. Regular [678, 1102, 978, 887, 580, 504, 1322, 958]. Regularity [1300]. reified [287]. Reinterpreting [1145]. Related [1210, 914, 135, 2]. Relation [887, 1170]. Relational [1082, 1291]. Relations [760, 1256, 1174, 1371, 714, 1171]. Relationships [512]. Relative [1068, 869, 1047]. Relativity [955]. Relaxation [671, 1197]. Relaxations [488]. relevance [195]. RElevancy [332]. Relevant [194, 1335, 182]. Rely [1053, 1141]. Rely-Guarantee [1053, 1141]. Remainder [1365]. Renaming [1321]. Renaming-Based [1321]. Rensets [1321]. Reordering [1018]. Replacement [395, 268]. Replacing [136]. Replay [770]. Report [953, 1064]. Representation [1006, 900, 805, 807, 646, 893, 176, 326, 63, 155, 159, 145, 197]. Representations [891, 1102, 999]. Representing [885, 647]. Reproved [1083]. Requirements [1254]. research [43]. Researchers [1369]. Residue [593]. Resolution [287, 838, 584, 926, 549, 1271, 464, 907, 945, 812, 413, 560, 752, 763, 1209, 1324, 1180, 1126, 495, 1020, 1100, 1206, 462, 1262, 1073, 1088, 504, 792, 134, 253, 220, 290, 306, 16, 243, 482, 555, 211, 371]. Resolution-Based [752, 763]. resolution/paramodulation [290]. Resolving [283, 431]. resonance [365]. Resource [1051, 1052]. Resources [1346]. Restart [1097]. Restore [504]. Restricted [1248, 1359]. Restriction [504, 165]. Result [870]. resulting [359]. Results [134, 684, 883, 1015, 1138, 402]. Results- [402]. Retrieval [767, 420, 225]. Reusing [836]. Revealed [536]. Reverse [1000]. Review [943, 686, 888, 255, 87, 240, 281, 288]. reviews [167]. revision [291]. Revisited [1366, 404, 1146]. Rewrite [937, 1222, 1312, 38, 330]. Rewriting [1085, 728, 760, 873, 756, 654, 1055, 667, 1355, 711, 1259, 1312, 586, 800, 929, 1341, 1040, 1286, 660, 1337, 917, 245, 45]. Right [723, 851, 218, 114]. Rigid [422, 442, 597]. Ring [1260, 57, 72]. rings [138, 96, 114]. Rippling [853, 356, 686]. Robbins [415, 160]. Robot [17]. ROGET [3]. Role [1090, 210, 4]. Root [583]. Roots [1175]. Rose [17]. Ross [1192]. Rosser [517]. Roth [1300]. Round [1338, 972]. Round-Off [1338]. Rounding [1024]. Routing [882]. RRTP [395]. Rubinstein [1192]. rue [294, 258]. Rule [718, 611, 632, 732, 395, 822, 280, 205, 312, 316, 98, 56, 108, 197, 223]. Rules [538, 1354, 450, 917, 38, 350, 99, 70, 152]. Runge [1338]. Running [673]. Runs [987, 738]. Runtime [1055, 929].

S [390, 1252]. S4 [497]. Safe [1069, 1096]. Safety [835, 1230, 1217, 1068, 942, 1160, 1033, 784]. Safety-Critical [835]. sages [52]. sat [712, 487, 680, 1366, 1097, 872, 678, 681, 930, 886, 1349, 1377, 1343, 558, 501, 491, 502, 658, 1183, 791, 641, 493, 679, 1373, 494, 997, 495, 738, 676, 1007, 1044, 713]. SAT-Based [558, 680, 1366]. SAT-provers [641]. Satchmo [396, 546, 332]. Satchmo- [396]. SATCHMORE [332]. Satisfaction [489, 904, 136]. Satisfiability [684, 881, 671, 857, 1186, 1263, 879, 1210, 677, 1343, 486, 556, 650, 670, 701, 489, 559, 560, 1364, 544, 1324, 938, 1253, 672, 813, 904, 1330, 1007, 488, 350, 246]. Satisfiable [673]. SATS [1071, 1071]. saturated [320]. Saturation [1173, 1279, 1337]. Saturation-Based [1173, 1337]. Scalability [767]. Scalable [1181, 1355]. Scale [1357]. Scheduler [786]. scheduling [22]. Schemas [970]. Schemata [1242]. Schematic [1242, 1344]. Scheme [879, 658]. Schemes [1145, 1240, 1352, 1126, 215]. Schubert [26]. Schutz [1293, 1307]. SCL [1320]. Search [693, 1106, 1349, 677, 491, 502, 1074, 1207, 1364, 996, 841, 494, 648, 503, 495, 1258, 418, 676, 442, 597, 440, 30]. Search-Pruning [418]. Searching [735, 695, 347]. SECD [470]. Second [1028, 1178, 1138]. Second-Order [1138]. Secret [536, 178]. section [200]. Secure [773]. Security [687, 844, 877, 873, 883, 1239, 1238, 690]. Seitz [882]. Selected [1194, 1133, 1159, 1213, 927, 1107, 1078, 1043, 1249, 1205, 863, 1066, 1270, 854, 1118, 1182]. selecting [254]. Selection [950, 838, 837, 17]. Selector [1027]. Self [1013, 1011, 431, 1336, 65, 93, 100]. self-analytically [93, 100]. Self-evident [1336]. Self-Formalisation [1011]. Self-modifying [1013]. Self-Resolving [431]. self-utterances [65]. Semantic [718, 1059, 421, 509, 574, 312, 12, 59, 152]. semantical [251]. Semantically [1006, 1056, 1304, 390]. Semantically-Guided [1006, 1056, 1304]. Semantics [1156, 941, 1136, 798, 549, 425, 1264, 526, 1144, 555, 1292, 260, 366, 283, 115, 297, 308]. Semi [1008, 528]. Semi-intelligible [1008]. Semi-Thue [528]. Semidefinite [488]. Semigroups [580, 41, 292, 60]. sense [257]. Sensitive [1006, 1056, 1304, 1259, 784]. Sentence [1085]. Sentence-Normalized [1085]. Sentential [539]. Separability [472]. Separated [1312]. Separation [680, 825, 1098, 1115, 853, 1328, 1371, 1030]. Sequence [930]. Sequences [1330]. sequencing [22]. Sequent [1223, 1345, 864, 933, 1351, 222, 156]. sequent-style [156]. sequential [336]. Series [861]. Service [920]. Set [1278, 42, 747, 657, 524, 701, 1225, 1241, 1318, 1356, 997, 944, 279, 342, 452, 651, 1142, 91, 280, 143, 172, 173, 249, 250, 348, 294, 174, 242, 175, 374, 206, 118, 64, 93, 117, 108, 132, 124, 165, 688, 275]. set-theoretic [348].

SET-VAR [275]. SETHEO [208, 397, 463]. Sets [1128, 991, 1174, 1248, 1111, 944, 817, 778, 674, 655]. Setting [835, 382, 563, 1178]. Seventy [33]. Seventy-five [33]. Shallow [1272, 1187, 1354]. Shannon [960]. Shapley [1348]. Sharing [1373, 593, 140]. Sheffer [610]. shell [116]. Shiny [1076]. shop [22]. Short [149, 573, 131]. Shortest [695, 430, 538, 610, 578]. Should [1342]. Signatures [1132]. Signed [427, 1255]. Signing [689]. Simple [566, 367, 377, 263]. Simplex [1134, 273]. Simplification [908, 259, 246]. Simplifications [906]. Simplifying [664]. Simulating [906, 1209]. Simulation [614]. Simulations [1236]. Simultaneous [474, 410, 442, 597]. Simultaneously [1071]. Single [554, 695, 334, 430, 233, 1356, 498, 241, 573, 410, 541, 578]. Single-Set [1356]. singleton [172, 174]. singularly [216]. Six [1276]. Size [1063, 117]. skeletal [8]. Skeptic [446]. Skeptical [549, 586]. Sketches [540]. Skew [703]. Skolem [345]. SLD [482]. SLDENF [371]. SLDENF-resolution [371]. SLDNF [211]. SLDNF-resolution [211]. Sledgehammer [932]. Slicing [639]. SLT [555]. SLT-Resolution [555]. SLUR [945]. SMA [1071]. Smallest [389]. SMELS [938]. Smith [1297, 1296]. Smooth [1379]. SMT [684, 1237, 921, 1081, 932, 1374, 985, 1017, 1045]. SMT-Based [1237, 1081]. SMT-COMP [684]. Social [1093, 799]. Software [835, 1081, 1237, 1366, 700, 953]. solid [329]. Solids [458]. Solution [1050, 898, 1148, 1035, 415, 899, 122]. solutions [26]. solve [125]. Solver [487, 1097, 1377, 1092, 997, 1149, 1302, 367]. Solvers [932, 1017, 658, 791, 1373, 813, 1045, 1036, 1044]. Solving [1381, 872, 1352, 886, 1174, 713, 677, 947, 428, 1165, 1287, 506, 679, 494, 790, 672, 540, 114, 645]. Some [712, 1091, 369, 483, 384, 685, 822, 57, 51, 29, 173, 229, 171]. Sort [1105]. Sorted [1076, 1086, 55]. Sorting [1063, 1350, 962].

Sound [718, 987, 924, 1375, 263]. Soundness [1145, 1037]. Space [1101, 915, 545]. Spaces [520, 875, 1077, 264]. Spacetime [1293, 1307]. SPARC [1241]. Sparse [1162]. SPASS [1187, 398]. SPASS-AR [1187]. Spatial [472, 512]. Special [810, 733, 739, 844, 1176, 927, 909, 1152, 873, 1249, 1205, 709, 863, 1051, 1019, 1127, 1270, 854, 818, 322, 337, 309, 351]. specialized [32]. Specification [682, 443, 1257, 936, 1292, 836]. Specifications [834, 1163, 566, 837, 594, 535]. specificity [339]. Speculative [857]. Speech [618]. Splitting [717, 909, 1314]. Springer [943]. SPS [463]. SPS-Parallelism [463]. SPTHEO [399, 463]. SQL [1294]. Square [583]. Squares [1050, 1358, 1363]. Stålmarck [779]. Stability [1363]. Stable [214, 549, 1332, 962]. stack [131]. State [1101, 875, 1305, 870, 30]. statements [144, 298]. Static [842, 1224, 808, 876]. Station [1257]. Stationary [1384]. status [135]. Steady [1305]. Steady-State [1305]. Stealing [658]. Stealth [1224]. Steamroller [26]. Stencil [1109]. Step [498]. Stickel [1005, 1004]. STMM [524]. Stochastic [677, 544, 1324, 841]. Stop [776]. Stop-and-Wait [776]. Store [979]. Strand [875]. Strategies [134, 1228, 423, 1202, 428, 646, 1180, 758, 495, 303, 277, 340, 296]. Strategy [451, 599, 219, 156, 187, 64, 78, 152, 197, 165, 256, 289]. stratified [272]. Strichman [943]. Strict [925]. String [667, 660]. Stroke [610]. Strong [1184, 595, 1209, 1332, 257]. Strongly [1272, 891, 1076]. Structural [507]. Structure [1203, 978, 140, 3, 150, 341]. structure-oriented [341]. Structured [806, 1383, 521, 285]. Structures [728, 967, 1198, 1124, 1360, 299]. Structuring [413]. Student [737, 637]. Students [631]. Studies [200, 535, 1058, 836, 358, 72]. Study [1103, 894, 1352, 491, 303, 215, 357, 52, 22, 256]. Sturm [980]. Style [664, 1265, 1041, 346, 156]. Subclasses [647]. Subformula [933, 946, 1301]. Subgoal [428]. Subject [466, 514, 530, 551, 571, 591, 608, 626]. subjective [251]. Sublanguage [651]. sublanguages [143, 172, 173, 174, 175]. Subrecursive [1153]. Subroutines [605]. Subset [798]. Subsets [748]. Subsolvers [674]. Substitution [889, 595, 817, 1057, 896]. substructural [314]. Subsumption [414, 1084, 381, 293, 186]. subtype [300]. Subtyping [533]. success [254]. Successful [697, 702]. Sufficient [882]. Suffix [1382]. Sums [1363, 993]. Superposition [858, 969, 1250, 1308, 1251, 1023, 721, 938, 1280]. Superposition-based [721]. Support [1278, 393, 774, 235, 64, 93, 165]. Supporting [1119, 251]. suppositional [158]. Suppressed [1215]. Survey [982, 1369, 845]. syllogistic [142, 172, 173, 174]. Sylow [478]. Symbolic [447, 878, 845, 672, 979]. Symbols [1272]. Symmetries [693, 286]. Symmetry [775]. Synchronisation [847]. synchronous [336]. Syntactic [620, 252]. Syntax [1156, 894, 1383, 999, 1188, 901, 890, 1321]. Synthesis [1211, 1385, 766, 443, 860, 1086, 453, 1376, 1305, 1072, 355]. Synthesising [1317]. synthesize [641]. Synthesizing [1214, 1220, 164]. Synthetic [1039, 1311, 484]. System [631, 948, 1123, 1224, 1056, 1186, 1315, 1325, 603, 992, 780, 1324, 126, 403, 1234, 1040, 433, 385, 387, 402, 468, 500, 543, 564, 612, 386, 439, 782, 325, 360, 3, 291, 294, 271, 335, 116, 318, 368, 297, 819]. System-Level [1123, 1224]. Systems [434, 664, 427, 1176, 982, 845, 1161, 654, 697, 746, 1163, 783, 667, 1236, 1184, 1264, 1357, 711, 1209, 1311, 566, 1222, 1155, 637, 528, 1312, 397, 826, 432, 768, 624, 1040, 1020, 321, 1160, 745, 792, 781, 1269, 870, 1044, 234, 253, 220, 127, 290, 314, 23, 202, 330, 283, 221, 45, 1266]. systolic [107]. Szekeres [1114]. Szemerédi [1300].

Tableau [657, 935, 954, 740, 720, 725, 313, 321, 905, 349, 316, 319, 77]. Tableau-Based [657, 954, 313, 321, 349, 77]. Tableaux [868, 505, 718, 1233, 310, 1343, 721, 924, 498, 1166, 222, 765, 755, 1243, 574, 253, 84, 312, 320, 59, 309, 191]. tables [30]. TACAS [1213]. Tactic [806]. tacticals [265]. Tactics [1232, 265]. TacticToe [1232]. Tail [470]. Tail-Recursive [470]. Tailed [489]. TAP [349]. Tarski [1039, 980, 109]. Tarskian [1038]. TBoxes [725]. Teaching [638]. technical [203]. Technique [416, 479, 418, 30]. Techniques [1013, 709, 1048, 1222, 568, 672, 432, 685, 762]. technology [95]. Temporal [1223, 1075, 600, 665, 907, 1343, 1221, 1069, 1291, 106, 1185, 32, 320, 184]. temporal-nonmonotonic [184]. tenth [226]. Term [891, 853, 756, 654, 1055, 633, 929, 1341, 1040, 1365, 1286, 352, 245, 238, 225, 45]. Terminating [946, 263]. Termination [1053, 886, 669, 756, 856, 667, 408, 709, 1032, 1068, 869, 922, 1047, 1171, 685, 662, 661, 1033, 642, 923, 1286, 660, 356, 306, 330, 24, 1341]. Terminological [741, 326, 339]. Terminologies [742]. Terms [805, 807, 1252, 407, 949, 646, 610, 63, 140, 83]. Test [1371, 837, 1084, 91, 259]. Testing [1164, 833, 436, 560, 1292, 836, 33]. Tests [832]. Text [1203]. Textbook [788]. Their [505, 1048, 1356, 339]. Them [1091, 259]. Theorem [810, 1050, 902, 733, 1089, 1194, 447, 750, 857, 974, 1369, 984, 909, 963, 930, 417, 1358, 510, 390, 383, 987, 449, 613, 789, 1173, 1300, 856, 831, 1041, 611, 797, 772, 527, 435, 1221, 1287, 478, 235, 1009, 986, 506, 1283, 637, 629, 454, 1180, 1316, 395, 1336, 328, 1365, 745, 399, 1187, 1279, 450, 518, 1060, 961, 287, 360, 315, 210, 207, 82, 191, 90, 44, 294, 192, 157, 353, 357, 343, 299, 4, 208, 369, 277, 52, 36, 77, 123, 33, 86, 313, 6, 341, 321, 266, 189, 57, 95, 311, 171, 247, 37, 1150, 1062, 1039]. Theorem [958]. theorem-proving [360, 294, 311]. Theorems [960, 1087, 862, 795, 914, 980, 983, 1245, 1254, 471, 38, 86, 72]. Theoretic [459, 1352, 1163, 650, 348, 370, 362, 369]. Theories [948, 684, 821, 1201, 1076, 1169, 877, 879, 525, 1334, 860, 876, 1086, 918, 410, 644, 1179, 663, 1255, 359, 327, 121, 32]. Theory [1326, 1274, 461, 460, 1167, 584, 1130, 1186, 1263, 1131, 1201, 657, 730, 1319, 1099, 1197, 1015, 1102, 1195, 524, 889, 1025, 727, 1188, 915, 1154, 1129, 1318, 521, 1138, 848, 1283, 1127, 483, 1312, 799, 1295, 484, 535, 1330, 955, 593, 1058, 148, 651, 1142, 956, 1260, 360, 41, 91, 214, 68, 42, 143, 172, 173, 174, 362, 67, 368, 242, 175, 279, 342, 374, 177, 206, 57, 16, 72, 108]. Things [698]. Third [709]. Threads [785]. Three [1058, 961, 285]. three-valued [285]. Thue [566, 528]. Tight [681, 1026]. Tilting [761]. Time [673, 1123, 1224, 1042, 1230, 1115, 600, 1070, 4]. Timed [1376]. Times [1384]. TMPR [285]. Tokeneer [1257]. Tolerant [1236]. Tool [682, 1098, 481, 426, 973, 1179]. Tool-Assisted [682]. Tools [851, 923, 1381]. Top [635, 181]. Top-Down [635, 181]. Topics [1210]. topology [333, 118]. Total [382]. Toupie [412]. TPS [360]. TPTP [441, 802, 1065]. Tractability [286]. Tractable [1274, 744, 867]. Tradeoff [1219]. Tradition [537]. traditional [329]. Transcendental [157, 1364]. Transformation [548, 453, 269, 180]. Transformations [711, 764, 205]. Transformer [1264]. Transition [1186, 566, 232]. Transitive [1256, 1325, 944]. Transitivity [1170]. Translating [749]. Translation [916, 568, 1070, 1199, 348]. Translations [1221]. treating [339]. treatment [209]. Tree [1022, 940, 1198, 715, 225, 285]. tree-structured [285]. Trees [747, 633, 409, 1117, 1193, 420, 224, 307]. Treewidth [1195]. Trends [532]. Triangles [1108]. Tribute [378]. tricks [20]. Tries [1306, 722]. Triggers [1017]. Triple [457]. Triples [1148]. trivial [1317]. Triviality [433]. True [1210]. Trust [624]. Trustworthy [621]. Truth [843]. Trybulec [989]. Tsirelson [1339]. TSO [1212, 1241]. Tsun [36]. Tuning [986]. Tuple [715, 1286]. Turing [1295]. tutorial [357]. Tutoring [637]. Twenty [576]. Twenty-Fourth [576]. Two [522, 75, 1358, 883, 895, 1124, 495, 925, 651, 86]. Two-Level [895, 651]. Two-Valued [925]. Type [1326, 1223, 584, 1130, 1131, 1216, 730, 1099, 1015, 1102, 481, 603, 992, 1334, 1129, 1054, 1138, 1111, 1096, 1127, 483, 480, 644, 784, 360, 253, 41, 226, 362, 40, 186]. Type-Based [1054]. Type-Safe [1096]. type-theoretic [362]. Typed [682, 891, 234, 83]. Types [1010, 1111, 580, 1029, 1375]. Typing [951, 624].

Unbounded [1230]. Unbounded-Time [1230]. Uncertain [818, 29]. Uncertainty [228, 1069]. Uncurrying [922]. Undecidability [1311, 258]. undecidable [68, 183]. unicorn [15, 125]. unifiability [91, 105]. Unification [643, 874, 1361, 1367, 1042, 292, 25, 830, 383, 422, 60, 507, 96, 580, 410, 706, 40, 452, 442, 597, 1290, 41, 68, 238, 23, 203, 226, 230, 237, 949, 732, 183, 880, 300, 136, 269, 61, 45, 270, 101]. Unifier [1028]. Unifiers [420, 203]. Uniform [1251, 1275, 277, 1057]. Unifying [1081, 1237, 1314]. UniMath [1132]. Uninitialised [1119]. Uninterpreted [1235]. Union [1115]. Unit [505, 945, 393, 359]. Unit-Refutation [945]. unit-resulting [359]. Unity [485]. Univalence [1128]. Univalent [1127].

Univariate [1104, 1026, 980]. Universes [533]. University [686, 888]. unnecessary [220]. Unnormalized [822]. Unpublished [1282]. Unranked [949]. Unsatisfiability [855, 569, 1373]. Unsatisfiable [748]. Unsolved [177]. Untrusted [1104]. Untyped [1142]. Update [409]. updates [370, 272]. Upper [1101, 842, 1339, 501]. Upper-bound [1339]. Upper-Bounding [1101]. Upside [311]. Upside-down [311]. Üresin [1197]. Use [411, 1362, 853, 213, 354, 186, 256]. Useful [1214]. User [734, 733, 736]. Users [1369]. Uses [764, 353]. Using [487, 835, 234, 1274, 898, 1136, 584, 737, 1103, 747, 1240, 669, 1017, 914, 851, 924, 632, 772, 560, 1069, 1162, 1082, 1104, 1171, 1289, 644, 983, 1193, 458, 817, 381, 125, 1052, 1160, 955, 358, 540, 1149, 1142, 488, 352, 29, 133, 249, 250, 357, 116, 99, 22, 212]. Utility [1072]. utterances [65].

Validation [849, 1035]. Validity [928].

Value [1137, 216]. Valued

[810, 909, 1210, 1335, 437, 925, 941, 285]. Values [843]. vampire [344]. Vanquishing [578]. VAR [275]. Variable [717, 718, 1312, 732, 410, 312, 61]. variable-abstraction [61]. Variable-Permuting [410]. Variable-Separated [1312]. Variables [772, 166]. Variance [772]. Variants [718, 396]. Variational [1225, 1110]. various [241]. Vector [1245, 1379, 1149]. Vectorial [506]. Vectors [1330, 1255, 1253]. Verification [952, 682, 606, 1081, 1366, 1067, 1230, 1094, 1217, 479, 1083, 1095, 665, 520, 973, 1109, 776, 1357, 700, 780, 1080, 602, 764, 953, 1220, 1070, 601, 708, 758, 936, 761, 1040, 623, 598, 1376, 757, 782, 781, 630, 836, 127, 112, 336, 113, 128, 641, 126, 279, 342, 340, 1237]. Verified [1101, 1368, 1062, 1224, 1093, 1136, 1119, 1097, 1013, 977, 1382, 1257, 1315, 1095, 1191, 1198, 1140, 1338, 954, 1092, 1177, 605, 1183, 801, 1096, 1113, 972, 480, 980, 1116, 1348, 1234, 1372, 129, 130]. Verifier [603, 604]. Verify [1098, 772, 955, 1044]. Verifying [1353, 688, 803, 416, 1115, 1148, 1351, 1068, 764, 707, 1288, 662, 1160, 1185, 1354, 1141, 1105, 164]. Verlag [943]. Versa [836]. Version [1188, 1321, 398]. Versus [410, 495]. Via [1376, 1164, 1333, 1275, 314, 1208, 1363, 667, 945, 1221, 1047, 1207, 765, 341, 83]. viability [258]. Vice [836]. Vice-Versa [836]. View [1081, 1237, 272, 409, 715, 370, 276, 943]. Violet [400]. Virtual [487]. Visibly [1075]. Visions [1277]. Visiting [1384]. Visualizing [738]. Visually [828, 829]. VLSI [246]. Volume [465, 467, 466, 513, 515, 514, 529, 550, 531, 552, 530, 551, 570, 590, 572, 592, 571, 591, 607, 625, 609, 627, 608, 626]. vs [880, 1354]. VST [1098]. VST-Floyd [1098]. VSTTE [1078].

W [480]. w.r.t [1272]. Wait [776]. WALDMEISTER [401]. Wanted [422]. Warshall [522]. Watson [527]. Wave [926]. Weak [1212, 937, 120]. Weighted [494]. Weighting [676]. weights [100]. Well [425, 526, 1171, 477, 555, 450, 178]. Well-Behaved [450]. Well-Founded [425, 526, 1171, 555]. Well-Foundedness [477]. well-kept [178]. Wen [36]. Wen-Tsun [36]. while [150]. Whiley [1288]. Wilson [6, 1058]. Winding [1175]. Windmills [1358, 761]. Wired [773]. Wireless [773]. Within [992, 481]. Without [1334, 890, 951, 317, 92]. Wombit [1149]. Woody [378, 429]. Word [1272, 562, 269, 1149]. Word-Level [1149]. Work [658, 1005, 1280]. World [120, 406, 511]. Worst [501, 491]. Worst-Case [501, 491]. Wos [1277, 1276, 1282, 1281]. Write [631].

Writing [14]. WS [1252]. Wu [36, 90].

X [172]. XCB [578]. XI [173]. XII [174]. XIII [175]. XOR [848].

Year [486, 556, 670]. Years [921, 146].

Z [819]. Zassenhaus [1191]. Zero [1215, 41, 226]. Zero-Suppressed [1215]. ZF [1138, 242]. Zisman [696].

References

Wos:1985:EJB

[1] Larry Wos. Editorial: A journal is born. Journal of Automated Reasoning, 1(1):1–3, March 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF00244287.

Wos:1985:OAR

[2] L. Wos, Fernando Pereira, Robert Hong, Robert S. Boyer, J. Strother Moore, W. W. Bledsoe, L. J. Henschen, Bruce G. Buchanan, Graham Wrightson, and Cordell Green. An overview of automated reasoning and related fields. Journal of Automated Reasoning, 1(1):5–48, March 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                              tronic). URL http://link.springer.
                                              com/article/10.1007/BF00244288.

Bennett:1985:RKB

[3] James S. Bennett. ROGET: A knowledge-based system for acquiring the conceptual structure of a diagnostic expert system. Journal of Automated Reasoning, 1(1):49–74, March 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244289.

Lengauer:1985:RAT

[4] Christian Lengauer. On the role of automated theorem proving in the compile-time derivation of concurrency. Journal of Automated Reasoning, 1 (1):75–101, March 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244290.

Lusk:1985:NHP

[5] Ewing Lusk and Ross Overbeek. Nonhorn problems. Journal of Automated Reasoning, 1(1):103–114, March 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244291.

Russinoff:1985:EBM

[6] David M. Russinoff. An experiment with the Boyer–Moore theorem prover: A proof of Wilson’s theorem. Journal of Automated Reasoning, 1(2):121–139, June 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244993.

Yahya:1985:DNH

[7] Adnan Yahya and Lawrence J. Henschen. Deduction in non-Horn databases. Journal of Automated Reasoning, 1(2):141–160, June 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00244994.

Friedland:1985:CIS

[8] Peter E. Friedland and Yumi Iwasaki. The concept and implementation of skeletal plans. Journal of Automated Reasoning, 1(2):161–208, June 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244995.

Lusk:1985:RAE

[9] E. L. Lusk and R. A. Overbeek. Reasoning about equality. Journal of Automated Reasoning, 1(2):209–228, June 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244996.

Anonymous:1985:Aa

[10] Anonymous. Announcements. Journal of Automated Reasoning, 1(2):229, June 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF00244997.

Anonymous:1985:CA

[11] Anonymous. Changes of address. Journal of Automated Reasoning, 1 (2):230, June 1985. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00244998.

McCune:1985:ESP

[12] William McCune and Lawrence Henschen. Experiments with semantic paramodulation. Journal of Automated Reasoning, 1(3):231–261, September 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244271.

Bundy:1985:ICM

[13] Alan Bundy. Incidence calculus: A mechanism for probabilistic reasoning. Journal of Automated Reasoning, 1(3): 263–283, September 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244272.

Constable:1985:WPC

[14] R. L. Constable, T. B. Knoblock, and J. L. Bates. Writing programs that construct proofs. Journal of Automated Reasoning, 1(3):285–326, September 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244273.

Ohlbach:1985:LU

[15] H. J. Ohlbach and M. Schmidt-Schauss. The lion and the unicorn. Journal of Automated Reasoning, 1(3): 327–332, September 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244274.

Stickel:1985:ADT

[16] Mark E. Stickel. Automated deduction by theory resolution. Journal of Automated Reasoning, 1(4):333–355, December 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244275.

Mauceri:1985:RSE

[17] John G. Mauceri. Robot selection expert ‘Rose’. Journal of Automated Reasoning, 1(4):357–390, December 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244276.

Doyle:1985:CID

[18] Jon Doyle. Circumscription and implicit definability. Journal of Automated Reasoning, 1(4):391–405, December 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244277.

Shankar:1985:TMM

[19] N. Shankar. Towards mechanical metamathematics. Journal of Automated Reasoning, 1(4):407–434, December 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244278.

Ohlbach:1985:PLH

[20] Hans Jürgen Ohlbach. Predicate logic hacker tricks. Journal of Automated Reasoning, 1(4):435–440, December 1985. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244279.

Anonymous:1985:Ab

[21] Anonymous. Announcement. Journal of Automated Reasoning, 1(4): 441, December 1985. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00244280.

Parrello:1986:JSS

[22] Bruce D. Parrello, Waldo C. Kabat, and L. Wos. Job-shop scheduling using automated reasoning: A case study of the car-sequencing problem. Journal of Automated Reasoning, 2(1):1–42, March 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00246021.

Dilger:1986:IBD

[23] Werner Dilger and Agnes Janson. Intelligent backtracking in deduction systems by means of extended unification graphs. Journal of Automated Reasoning, 2(1):43–62, March 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00246022.

Paulson:1986:PTN

[24] Lawrence C. Paulson. Proving termination of normalization functions for conditional expressions. Journal of Automated Reasoning, 2(1):63–74, March 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00246023.

Buttner:1986:UDM

[25] Wolfram Büttner. Unification in datastructure multisets. Journal of Automated Reasoning, 2(1):75–88, March 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00246024. Dis-

cusses generalization of extendible hashing to handle partial match retrieval efficiently.

Stickel:1986:SSP

[26] Mark E. Stickel. Schubert’s steamroller problem: Formulations and solutions. Journal of Automated Reasoning, 2(1):89–101, March 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00246025.

Anonymous:1986:ANJa

[27] Anonymous. Announcing a new journal. Journal of Automated Reasoning, 2(1):103, March 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00246026.

Anonymous:1986:AEC

[28] Anonymous. Addresses of Editor-in-Chief and Editorial Board. Journal of Automated Reasoning, 2(1):105–107, March 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00246027.

Bundy:1986:CCS

[29] Alan Bundy. Correctness criteria of some algorithms for uncertain reasoning using incidence calculus. Journal of Automated Reasoning, 2(2):109–126, June 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02432147.

Kramer:1986:AAO

[30] Thomas Kramer. Automated analysis of operators on state tables: A technique for intelligent search. Journal of Automated Reasoning, 2(2):127–153, June 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02432148.

Aquilano:1986:NFC

[31] C. Aquilano, R. Barbuti, P. Bocchetti, and M. Martelli. Negation as failure. Completeness of the query evaluation process for Horn clause programs with recursive definitions. Journal of Automated Reasoning, 2(2):155–170, June 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02432149.

Plaisted:1986:DPC

[32] David A. Plaisted. A decision procedure for combinations of propositional temporal logic and other specialized theories. Journal of Automated Reasoning, 2(2):171–190, June 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02432150.

Pelletier:1986:SFP

[33] Francis Jeffry Pelletier. Seventyfive problems for testing automatic theorem provers. Journal of Automated Reasoning, 2(2):191–216, June 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02432151.

Anonymous:1986:CP

[34] Anonymous. Call for papers. Journal of Automated Reasoning, 2(2):217, June 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02432152.

Anonymous:1986:ANJb

[35] Anonymous. Announcing a new journal. Journal of Automated Reasoning, 2(2):218, June 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02432153.

Moore:1986:EPL

[36] J. Strother Moore. Editor’s preface to ‘Basic principles of mechanical theorem proving in elementary geometries’ by Wu Wen-Tsun. Journal of Automated Reasoning, 2(3):219– 220, September 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02328446. See [37].

Wu:1986:BPM

[37] Wen-Tsun Wu. Basic principles of mechanical theorem proving in elementary geometries. Journal of Automated Reasoning, 2(3):221–252, September 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02328447. See

preface [36].

Chou:1986:PGT

[38] Shang-Ching Chou and William F. Schelter. Proving geometry theorems with rewrite rules. Journal of Automated Reasoning, 2(3):253– 273, September 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02328448.

Siekmann:1986:EN

[39] Joerg H. Siekmann. Editor’s note. Journal of Automated Reasoning, 2(3): 275, September 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02328449.

Schmidt-Schauss:1986:UUA

[40] Manfred Schmidt-Schauss. Unification under associativity and idempotence is of type nullary. Journal of Automated Reasoning, 2(3):277–281, September 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02328450.

Baader:1986:TIS

[41] Franz Baader. The theory of idempotent semigroups is of unification type zero. Journal of Automated Reasoning, 2(3):283–286, September 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02328451.

Boyer:1986:STF

[42] Robert Boyer, Ewing Lusk, William McCune, Ross Overbeek, Mark Stickel, and Lawrence Wos. Set theory in firstorder logic: Clauses for Gödel’s axioms. Journal of Automated Reasoning, 2(3): 287–327, September 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02328452.

Anonymous:1986:EAN

[43] Anonymous. Editorial announcement a new column: Basic research problems. Journal of Automated Reasoning, 2(4):iii, December 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00248248.

Coelho:1986:ARG

[44] Helder Coelho and Luis Moniz Pereira. Automated reasoning in geometry theorem proving with Prolog. Journal of Automated Reasoning, 2(4): 329–390, December 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00248249.

You:1986:CCT

[45] Jia-Huai You and P. A. Subrahmanyam. A class of confluent term rewriting systems and unification. Journal of Automated Reasoning, 2(4): 391–418, December 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00248250.

Glickfeld:1986:FCL

[46] Barney Glickfeld and Ross Overbeek. A foray into combinatory logic. Journal of Automated Reasoning, 2(4): 419–431, December 1986. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00248251.

Anonymous:1986:A

[47] Anonymous. Announcement. Journal of Automated Reasoning, 2(4): 433, December 1986. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00248252.

Ohlbach:1987:LIA

[48] Hans Jürgen Ohlbach. Link inheritance in abstract clause graphs. Journal of Automated Reasoning, 3(1):1–34, March 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00381143.

Wang:1987:HD

[49] Tie-Cheng Wang and W. W. Bledsoe. Hierarchical deduction. Journal of Automated Reasoning, 3(1):35–77, March 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00381144.

Wos:1987:EN

[50] Larry Wos. Editor’s note. Journal of Automated Reasoning, 3(1): 79–80, March 1987. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00381145.

Wos:1987:SOA

[51] Larry Wos. Some obstacles to the automation of reasoning, and the problem of redundant information. Journal of Automated Reasoning, 3(1):81–90, March 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00381146.

McCune:1987:CSA

[52] William McCune and Larry Wos. A case study in automated theorem proving: Finding sages in combinatory logic. Journal of Automated Reasoning, 3(1):91–107, March 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00381147.

Anonymous:1987:CA

[53] Anonymous. Change of address. Journal of Automated Reasoning, 3(1):109, March 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00381148.

Anonymous:1987:Aa

[54] Anonymous. Announcements. Journal of Automated Reasoning, 3(1):109–111, March 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00381149.

Cohn:1987:MEF

[55] Anthony G. Cohn. A more expressive formulation of many sorted logic. Journal of Automated Reasoning, 3 (2):113–200, June 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243207.

Wos:1987:PCI

[56] Larry Wos. The problem of choosing the inference rule to employ. Journal of Automated Reasoning, 3(2):201–209, June 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243208.

Stevens:1987:SEN

[57] Rick L. Stevens. Some experiments in nonassociative ring theory with an automated theorem prover. Journal of Automated Reasoning, 3(2):211–221, June 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243209.

Anonymous:1987:Ab

[58] Anonymous. Announcement. Journal of Automated Reasoning, 3(2):223, June 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00243210.

Reeves:1987:AES

[59] S. V. Reeves. Adding equality to semantic tableaux. Journal of Automated Reasoning, 3(3):225–246, September 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243790.

Herold:1987:UAS

[60] Alexander Herold and Jörg H. Siekmann. Unification in Abelian semigroups. Journal of Automated Reasoning, 3(3):247–283, September 1987.

CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00243791.

Stickel:1987:CVA

[61] Mark E. Stickel. A comparison of the variable-abstraction and constantabstraction methods for associativecommutative unification. Journal of Automated Reasoning, 3(3):285– 289, September 1987. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243792.

Chou:1987:MMD

[62] Shang-Ching Chou. A method for the mechanical derivation of formulas in elementary geometry. Journal of Automated Reasoning, 3(3): 291–299, September 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243793.

Lassez:1987:ERT

[63] J.-L. Lassez and K. Marriott. Explicit representation of terms defined by counter examples. Journal of Automated Reasoning, 3(3): 301–317, September 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243794.

Wos:1987:PES

[64] Larry Wos. The problem of extending the set of support strategy. Journal of Automated Reasoning, 3(3):

319–328, September 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243795.

Miller:1987:PSU

[65] Michael Miller and Donald Perlis. Proving self-utterances. Journal of Automated Reasoning, 3(3):329– 338, September 1987. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243796.

Anonymous:1987:Ac

[66] Anonymous. Announcements. Journal of Automated Reasoning, 3(3): 339–342, September 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243797.

Manna:1987:HCB

[67] Zohar Manna and Richard Waldinger. How to clear a block: A theory of plans. Journal of Automated Reasoning, 3(4):343–377, December 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247434.

Bockmayr:1987:NCT

[68] Alexander Bockmayr. A note on a canonical theory with undecidable unification and matching problem. Journal of Automated Reasoning, 3(4): 379–381, December 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00247435.

Rudnicki:1987:OI

[69] Piotr Rudnicki. Obvious inferences. Journal of Automated Reasoning, 3(4): 383–393, December 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00247436.

Treitel:1987:CDR

[70] Richard Treitel and Michael R. Genesereth. Choosing directions for rules. Journal of Automated Reasoning, 3(4): 395–431, December 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00247437.

Wos:1987:PDE

[71] Larry Wos. The problem of definition expansion and contraction. Journal of Automated Reasoning, 3(4): 433–435, December 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00247438.

Wang:1987:CSM

[72] Tie-Cheng Wang. Case studies of Z-module reasoning: Proving benchmark theorems from ring theory. Journal of Automated Reasoning, 3(4): 437–451, December 1987. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00247439.

Anonymous:1988:PAa

[73] Anonymous. Publisher’s announcement. Journal of Automated Reasoning, 4(1):v, March 1988. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00244509.

Hintikka:1988:MMA

[74] Jaakko Hintikka. Model minimization— an alternative to circumscription. Journal of Automated Reasoning, 4(1):1–13, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244510.

Bundy:1988:MLI

[75] Alan Bundy and Leon Sterling. Metalevel inference: Two applications. Journal of Automated Reasoning, 4 (1):15–27, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244511.

Helman:1988:DDD

[76] Paul Helman and Robert Veroff. Designing deductive databases. Journal of Automated Reasoning, 4(1):29–68, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244512.

Oppacher:1988:HTB

[77] F. Oppacher and E. Suen. HARP: A tableau-based theorem prover. Journal of Automated Reasoning, 4(1):69–100, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244513.

Wos:1988:PFS

[78] Larry Wos. The problem of finding a strategy to control binary paramodulation. Journal of Automated Reasoning, 4(1):101–107, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244514.

Anonymous:1988:Ea

[79] Anonymous. Errata. Journal of Automated Reasoning, 4(1):109–111, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/BF00244515.
pdf.

Anonymous:1988:BR

[80] Anonymous. Book received. Journal of Automated Reasoning, 4(1):113, March 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00244516.

Anonymous:1988:A

[81] Anonymous. Announcements. Journal of Automated Reasoning, 4(1): 115–116, March 1988. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00244517.

Boyer:1988:ABQ

[82] Robert S. Boyer and J. Strother Moore. The addition of bounded quantification and partial functions to a computational logic and its theorem prover. Journal of Automated Reasoning, 4(2):117–172, June 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244392.

Zaionc:1988:MPP

[83] Marek Zaionc. Mechanical procedure for proof construction via closed terms in typed λ calculus. Journal of Automated Reasoning, 4(2):173–190, June 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244393.

Fitting:1988:FOM

[84] Melvin Fitting. First-order modal tableaux. Journal of Automated Reasoning, 4(2):191–213, June 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00244394.

Wos:1988:PED

[85] Larry Wos. The problem of explaining the disparate performance of hyperresolution and paramodulation. Journal of Automated Reasoning, 4(2):215–217, June 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244395.

Quaife:1988:APL

[86] Art Quaife. Automated proofs of Löb’s theorem and Gödel’s two incompletensess theorems. Journal of Automated Reasoning, 4(2):219–231, June 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244396.

Overbeek:1988:BR

[87] Ross Overbeek. Book review. Journal of Automated Reasoning, 4(2): 233–234, June 1988. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00244397.

Anonymous:1988:Eb

[88] Anonymous. Errata. Journal of Automated Reasoning, 4(2):235–236, June 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        content/pdf/10.1007/BF00244398.
                                        pdf.

Anonymous:1988:PAb

[89] Anonymous. Publisher’s announcement. Journal of Automated Reasoning, 4(3):iii, September 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00244941.

Chou:1988:IWM

[90] Shang-Ching Chou. An introduction to Wu’s method for mechanical theorem proving in geometry. Journal of Automated Reasoning, 4(3): 237–267, September 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00244942.

Bailin:1988:UTS

[91] Sidney C. Bailin. A λ-unifiability test for set theory. Journal of Automated Reasoning, 4(3):269–286, September 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244943.

Plaisted:1988:NHC

[92] David A. Plaisted. Non-Horn clause logic programming without contrapositives. Journal of Automated Reasoning, 4(3):287–325, September 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244944.

Wos:1988:PSAa

[93] Larry Wos. The problem of selfanalytically choosing the set of support. Journal of Automated Reasoning, 4(3):327–329, September 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00244945.

Kapur:1988:PED

[94] Deepak Kapur and Hantao Zhang. Proving equivalence of different axiomatizations of free groups. Journal of Automated Reasoning, 4(3): 331–352, September 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244946.

Stickel:1988:PTT

[95] Mark E. Stickel. A Prolog technology theorem prover: Implementation by an extended Prolog compiler. Journal of Automated Reasoning, 4(4): 353–380, December 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00297245.

Martin:1988:UBR

[96] Ursula Martin and Tobias Nipkow. Unification in Boolean rings. Journal of Automated Reasoning, 4(4): 381–396, December 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00297246.

Ross:1988:INI

[97] Kenneth A. Ross and Rodney W. Topor. Inferring negative information from disjunctive databases. Journal of Automated Reasoning, 4(4): 397–424, December 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00297247.

Munch:1988:NRR

[98] Klaus Heje Munch. A new reduction rule for the connection graph proof procedure. Journal of Automated Reasoning, 4(4):425–444, December 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00297248.

Monk:1988:IRU

[99] Leonard G. Monk. Inference rules using local contexts. Journal of Automated Reasoning, 4(4):445–462, December 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00297249.

Wos:1988:PSAb

[100] Larry Wos. The problem of selfanalytically choosing the weights. Journal of Automated Reasoning, 4(4): 463–464, December 1988. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00297250.

Burckert:1988:OAU

[101] Hans-Jürgen Bürckert, Alexander Herold, and Deepak Kapur. Opening the AC-unification race. Journal of Automated Reasoning, 4(4):465–474, December 1988. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00297251.

Anonymous:1989:PA

[102] Anonymous. Publisher’s announcement. Journal of Automated Reasoning, 5(1):v, March 1989. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00245017.

Lifschitz:1989:WIM

[103] Vladimir Lifschitz. What is the inverse method? Journal of Automated Reasoning, 5(1):1–23, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245018.

Lankford:1989:NNI

[104] Dallas Lankford. Non-negative integer basis algorithms for linear equations with integer coefficients. Journal of Automated Reasoning, 5(1):25–35, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00245019.

Wolfram:1989:IUP

[105] D. A. Wolfram. Intractable unifiability problems and backtracking. Journal of Automated Reasoning, 5(1):37–47, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00245020.

Tang:1989:TLC

[106] Tong Gao Tang. Temporal logic CTL + PROLOG. Journal of Automated Reasoning, 5(1):49–65, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00245021.

Purushothaman:1989:MCS

[107] S. Purushothaman and P. A. Subrahmanyam. Mechanical certification of systolic algorithms. Journal of Automated Reasoning, 5(1):67–91, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00245022.

Wos:1989:PFI

[108] Larry Wos. The problem of finding an inference rule for set theory. Journal of Automated Reasoning, 5(1):93–95, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245023.

Quaife:1989:ADT

[109] Art Quaife. Automated development of Tarski’s geometry. Journal of Automated Reasoning, 5(1):97–118, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245024.

Anonymous:1989:AEB

[110] Anonymous. Addresses of editorial board. Journal of Automated Reasoning, 5(1):119–121, March 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00245025.

Anonymous:1989:A

[111] Anonymous. Announcements. Journal of Automated Reasoning, 5(1): 123–124, March 1989. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00245026.

Boyer:1989:NV

[112] Robert Boyer. Note on verification. Journal of Automated Reasoning, 5 (2):125, June 1989. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00242999.

Cohn:1989:NPH

[113] Avra Cohn. The notion of proof in hardware verification. Journal of Automated Reasoning, 5(2):127–139, June 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243000.

Wang:1989:SOP

[114] Tie-Cheng Wang and Rick Stevens. Solving open problems in right alternative rings with Z-module reasoning. Journal of Automated Reasoning, 5(2):141–165, June 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243001.

Przymusinski:1989:DPS

[115] Teodor C. Przymusinski. On the declarative and procedural semantics of logic programs. Journal of Automated Reasoning, 5(2):167–205, June 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243002.

Leung:1989:FES

[116] K. S. Leung and W. Lam. A fuzzy expert system shell using both exact and inexact reasoning. Journal of Automated Reasoning, 5(2):207–233, June 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243003.

Wos:1989:PDS

[117] Larry Wos. The problem of determining the size of a complete set of reductions. Journal of Automated Reasoning, 5(2):235–237, June 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243004.

Wick:1989:ARA

[118] Cynthia A. Wick and William W. Mc-Cune. Automated reasoning about elementary point-set topology. Journal of Automated Reasoning, 5(2):239–255, June 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243005.

Andrews:1989:CHO

[119] Peter B. Andrews. On connections and higher-order logic. Journal of Automated Reasoning, 5(3): 257–291, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00248320.

Rajasekar:1989:WGC

[120] Arcot Rajasekar, Jorge Lobo, and Jack Minker. Weak generalized closed world assumption. Journal of Automated Reasoning, 5(3):293–307, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00248321.

Furbach:1989:HET

[121] Ulrich Furbach, Steffen Hölldobler, and Joachim Schreiber. Horn equational theories and paramodulation. Journal of Automated Reasoning, 5(3): 309–337, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00248322.

Char:1989:DIC

[122] Bruce W. Char, Alan R. Macnaughton, and Paul A. Strooper. Discovering inequality conditions in the analytic solution of optimization problems. Journal of Automated Reasoning, 5(3): 339–362, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00248323.

Paulson:1989:FGT

[123] Lawrence C. Paulson. The foundation of a generic theorem prover. Journal of Automated Reasoning, 5(3): 363–397, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00248324.

Wos:1989:PGE

[124] Larry Wos. The problem of guaranteeing the existence of a complete set of reductions. Journal of Automated Reasoning, 5(3):399–401, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00248325.

Schmid:1989:UEP

[125] Richard Schmid, Hans-Albert Schneider, and Thomas Filkorn. Using an extended PROLOG to solve the lion and unicorn puzzle. Journal of Automated Reasoning, 5(3):403–408, September 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00248326.

Moore:1989:SV

[126] J. Strother Moore. System verification. Journal of Automated Reasoning, 5(4): 409–410, December 1989. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00243130.

Bevier:1989:ASV

[127] William R. Bevier, Warren A. Hunt, Jr., J. Strother Moore, and William D. Young. An approach to systems verification. Journal of Automated Reasoning, 5(4):411–428, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243131.

Hunt:1989:MDV

[128] Warren A. Hunt, Jr. Microprocessor design verification. Journal of Automated Reasoning, 5(4):429–460, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243132.

Moore:1989:MVL

[129] J. Strother Moore. A mechanically verified language implementation. Journal of Automated Reasoning, 5(4): 461–492, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243133.

Young:1989:MVC

[130] William D. Young. A mechanically verified code generator. Journal of Automated Reasoning, 5(4):493–518, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243134.

Bevier:1989:KSS

[131] William R. Bevier. Kit and the short stack. Journal of Automated Reasoning, 5(4):519–530, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243135.

Wos:1989:PGA

[132] Larry Wos. The problem of guaranteeing the absence of a complete set of reductions. Journal of Automated Reasoning, 5(4):531–532, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00243136.

Chisholm:1989:ARP

[133] G. H. Chisholm, B. T. Smith, and A. S. Wojcik. An automated reasoning problem associated with proving claims about programs using Floyd– Hoare inductive assertion methods. Journal of Automated Reasoning, 5(4): 533–540, December 1989. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243137.

Auffray:1990:SMR

[134] Yves Auffray and Jean-Jacques Hebrard. Strategies for modal resolution: Results and problems. Journal of Automated Reasoning, 6(1):1–38, March 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00302639.

Lescanne:1990:RDO

[135] Pierre Lescanne. On the recursive decomposition ordering with lexicographical status and other related orderings. Journal of Automated Reasoning, 6(1):39–49, March 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00302640.

Roach:1990:RUC

[136] J. W. Roach, R. Sundararajan, and L. T. Watson. Replacing unification by constraint satisfaction to improve logic program expressiveness. Journal of Automated Reasoning, 6(1):51–75, March 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00302641.

Wos:1990:PCBa

[137] Larry Wos. The problem of choosing between logic programming and general-purpose automated reasoning. Journal of Automated Reasoning, 6 (1):77–78, March 1990. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00302642.

Anantharaman:1990:APM

[138] Siva Anantharaman and Jieh Hsiang. Automated proofs of the Moufang identities in alternative rings. Journal of Automated Reasoning, 6(1):79–109, March 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00302643.

Anonymous:1990:EA

[139] Anonymous. Editorial announcement. Journal of Automated Reasoning, 6 (2):iii, June 1990. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00245814.

Staples:1990:SSQ

[140] John Staples and Peter J. Robinson. Structure sharing for quantified terms: Fundamentals. Journal of Automated Reasoning, 6(2):115–145, June 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00245815.

Lu:1990:PCF

[141] James J. Lu and V. S. Subrahmanian. Protected completions of firstorder general logic programs. Journal of Automated Reasoning, 6(2):147–172, June 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00245816.

Cantone:1990:AS

[142] D. Cantone, E. G. Omodeo, and A. Policriti. The automation of syllogistic. Journal of Automated Reasoning, 6(2):173–187, June 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00245817.

Cantone:1990:DPE

[143] D. Cantone and V. Cutello. Decision procedures for elementary sublanguages of set theory. Journal of Automated Reasoning, 6(2):189–201, June 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245818.

Ferro:1990:PPS

[144] Giuseppa Carrà Ferro and Giovanni Gallo. A procedure to prove statements in differential geometry. Journal of Automated Reasoning, 6(2):203–209, June 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245819.

Wos:1990:PFM

[145] Larry Wos. The problem of finding a mapping between clause representation and natural-deduction representation. Journal of Automated Reasoning, 6(2):211–212, June 1990. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00245820.

Wos:1990:MCF

[146] Larry Wos. Meeting the challenge of fifty years of logic. Journal of Automated Reasoning, 6(2):213–232, June 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245821.

Stokes:1990:GBE

[147] Timothy Stokes. Gröbner bases in exterior algebra. Journal of Automated Reasoning, 6(3):233–250, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244487.

Yu:1990:CPG

[148] Yuan Yu. Computer proofs in group theory. Journal of Automated Reasoning, 6(3):251–286, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244488.

Bibel:1990:SPP

[149] W. Bibel. Short proofs of the pigeonhole formulas based on the connection method. Journal of Automated Reasoning, 6(3):287–297, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244489.

DiPiazza:1990:IKE

[150] Joseph S. Di Piazza. Interweaving knowledge extracting, organizing and evaluating: A concrete design for preventing logic and structure bugs while interviewing experts. Journal of Automated Reasoning, 6(3): 299–317, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244490.

Buning:1990:EPP

[151] Hans Kleine Büning, Ulrich Löwen, and Stefan Schmitgen. Equivalence of propositional Prolog programs. Journal of Automated Reasoning, 6(3):

319–335, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244491.

Wos:1990:PFS

[152] Larry Wos. The problem of finding a semantic strategy for focusing inference rules. Journal of Automated Reasoning, 6(3):337–339, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244492.

Bledsoe:1990:CPE

[153] W. W. Bledsoe. Challenge problems in elementary calculus. Journal of Automated Reasoning, 6(3): 341–359, September 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244493.

Anonymous:1990:PA

[154] Anonymous. Publisher’s announcement. Journal of Automated Reasoning, 6(4):iii, December 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00244353.

Patel-Schneider:1990:DFO

[155] Peter F. Patel-Schneider. A decidable first-order logic for knowledge representation. Journal of Automated Reasoning, 6(4):361–388, December 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244354.

Plaisted:1990:SSM

[156] David A. Plaisted. A sequent-style model elimination strategy and a positive refinement. Journal of Automated Reasoning, 6(4):389–402, December 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244355.

Gao:1990:TFM

[157] Xiaoshan Gao. Transcendental functions and mechanical theorem proving in elementary geometries. Journal of Automated Reasoning, 6(4): 403–417, December 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00244356.

Pollock:1990:IDS

[158] John L. Pollock. Interest driven suppositional reasoning. Journal of Automated Reasoning, 6(4):419–461, December 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244357.

Wos:1990:PCBb

[159] Larry Wos. The problem of choosing between predicate and function notation for problem representation. Journal of Automated Reasoning, 6(4):463– 464, December 1990. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00244358.

Winker:1990:RAC

[160] S. Winker. Robbins algebra: Conditions that make a near-boolean algebra boolean. Journal of Automated Reasoning, 6(4):465–489, December 1990. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244359.

Anonymous:1990:CP

[161] Anonymous. Call for papers. Journal of Automated Reasoning, 6(4):491– 492, December 1990. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00244360.

Anonymous:1991:PA

[162] Anonymous. Publisher’s announcement. Journal of Automated Reasoning, 7(1):v, March 1991. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00249352.

Loveland:1991:NHP

[163] Donald W. Loveland. Near-Horn Prolog and beyond. Journal of Automated Reasoning, 7(1):1–26, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00249353.

Padawitz:1991:IEC

[164] Peter Padawitz. Inductive expansion: A calculus for verifying and synthesizing functional and logic programs. Journal of Automated Reasoning, 7 (1):27–103, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00249354.

Wos:1991:PFR

[165] Larry Wos. The problem of finding a restriction strategy more effective than the set of support strategy. Journal of Automated Reasoning, 7(1):105–107, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00249355.

Kaufmann:1991:GPF

[166] Matt Kaufmann. Generalization in the presence of free variables: A mechanically-checked correctness proof for one algorithm. Journal of Automated Reasoning, 7(1):109–158, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00249356.

Stevens:1991:BR

[167] Rick Stevens. Book reviews. Journal of Automated Reasoning, 7(1):159–161, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00249357.

Anonymous:1991:E

[168] Anonymous. Errata. Journal of Automated Reasoning, 7(1):163, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/BF00249358.
pdf.

Anonymous:1991:AEB

[169] Anonymous. Addresses of editorial board. Journal of Automated Reasoning, 7(1):165–167, March 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00249359.

Ferro:1991:I

[170] Alfredo Ferro. Introduction. Journal of Automated Reasoning, 7(2): 169–170, June 1991. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00243805.

Wen-Tsun:1991:MTP

[171] Wu Wen-Tsun. Mechanical theorem proving of differential geometries and some of its applications in mechanics. Journal of Automated Reasoning, 7(2):171–191, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243806.

Cantone:1991:DPEa

[172] D. Cantone. Decision procedures for elementary sublanguages of set theory: X. Multilevel syllogistic extended by the singleton and powerset operators. Journal of Automated Reasoning, 7(2):193–230, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243807.

Cantone:1991:DPEb

[173] D. Cantone and J. T. Schwartz. Decision procedures for elementary sublanguages of set theory: XI. Multilevel syllogistic extended by some elementary map constructs. Journal of Automated Reasoning, 7(2):231–256, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243808.

Ferro:1991:DPE

[174] A. Ferro. Decision procedures for elementary sublanguages of set theory: XII. Multilevel syllogistic extended with singleton and choice operators. Journal of Automated Reasoning, 7(2):257–270, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00243809.

Parlamento:1991:DPE

[175] Franco Parlamento and Alberto Policriti. Decision procedures for elementary sublanguages of set theory: XIII. Model graphs, reflection and decidability. Journal of Automated Reasoning, 7(2):271–284, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243810.

Andrews:1991:MPF

[176] Peter B. Andrews. More on the problem of finding a mapping between clause representation and naturaldeduction representation. Journal of Automated Reasoning, 7(2):285–286, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00243811.

Quaife:1991:UPE

[177] Art Quaife. Unsolved problems in elementary number theory. Journal of Automated Reasoning, 7(2):287–300, June 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00243812.

Anonymous:1991:WKS

[178] Anonymous. A well-kept secret of history. Journal of Automated Reasoning, 7(3):301–302, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00249015.

Bundy:1991:EPP

[179] Alan Bundy, Frank van Harmelen, Jane Hesketh, and Alan Smaill. Experiments with proof plans for induction. Journal of Automated Reasoning, 7(3): 303–324, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00249016.

Socher:1991:OCN

[180] Rolf Socher. Optimizing the clausal normal form transformation. Journal of Automated Reasoning, 7(3): 325–336, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00249017.

DeSchreye:1991:CBM

[181] Danny De Schreye, Bern Martens, Gunther Sablon, and Maurice Bruynooghe. Compiling bottom-up and mixed derivations into top-down executable logic programs. Journal of Automated Reasoning, 7(3):337–358, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00249018.

Ramsay:1991:GRM

[182] Allan Ramsay. Generating relevant models. Journal of Automated Reasoning, 7(3):359–368, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00249019.

Nutt:1991:UHU

[183] Werner Nutt. The unification hierarchy is undecidable. Journal of Automated Reasoning, 7(3):369–381, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00249020.

Tang:1991:PTN

[184] Tong Gao Tang. Programming in temporal-nonmonotonic reasoning. Journal of Automated Reasoning, 7(3): 383–401, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00249021.

Feldman:1991:PDI

[185] Yishai A. Feldman and Charles Rich. Pattern-directed invocation with changing equations. Journal of Automated Reasoning, 7(3):403– 433, September 1991. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00249022.

Wos:1991:PCT

[186] Larry Wos. The problem of choosing the type of subsumption to use. Journal of Automated Reasoning, 7(3): 435–438, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00249023.

Wilding:1991:PML

[187] Matthew Wilding. Proving Matijasevich’s lemma with a default arithmetic strategy. Journal of Automated Reasoning, 7(3):439–446, September 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00249024.

McRobbie:1991:ARN

[188] Michael A. McRobbie. Automated reasoning and nonclassical logics: Introduction. Journal of Automated Reasoning, 7(4):447–451, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01880323.

Slaney:1991:ACT

[189] John K. Slaney. The Ackermann constant theorem: A computer-assisted investigation. Journal of Automated Reasoning, 7(4):453–474, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01880324.

Pritchard:1991:AFM

[190] Paul Pritchard. Algorithms for finding matrix models of propositional calculi. Journal of Automated Reasoning, 7(4):475–487, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01880325.

Catach:1991:TGT

[191] Laurent Catach. TABLEAUX: A general theorem prover for modal logics. Journal of Automated Reasoning, 7(4): 489–510, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01880326.

Gabbay:1991:FGD

[192] Dov M. Gabbay and Frank Kriwaczek. A family of goal directed theorem provers based on conjunction and implication: Part 1. Journal of Automated Reasoning, 7(4):511–536, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01880327.

DeJongh:1991:CFI

[193] Dick De Jongh, Lex Hendriks, and Gerard R. Renardel De Lavalette. Computations in fragments of intuitionistic propositional logic. Journal of Automated Reasoning, 7(4):537–561, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01880328.

Bollen:1991:RLP

[194] A. W. Bollen. Relevant logic programming. Journal of Automated Reasoning, 7(4):563–585, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01880329.

Mints:1991:CDC

[195] Grigori Mints and Tanel Tammet. Condensed detachment is complete for relevance logic: A computeraided proof. Journal of Automated Reasoning, 7(4):587–596, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01880330.

Meyer:1991:ILF

[196] Robert K. Meyer, Martin W. Bunder, and Lawrence Powers. Implementing the ‘Fool’s model’ of combinatory logic. Journal of Automated Reasoning, 7(4):597–630, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01880331.

Wos:1991:PCR

[197] Larry Wos. The problem of choosing the representation, inference rule, and strategy. Journal of Automated Reasoning, 7(4):631–634, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01880332.

Thistlewaite:1991:AHN

[198] Paul B. Thistlewaite and Michael A. McRobbie. Approaching hard nonclassical problems. Journal of Automated Reasoning, 7(4):635–637, December 1991. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01880333.

Anonymous:1992:PA

[199] Anonymous. Publisher’s announcement. Journal of Automated Reasoning, 8(1):v, February 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00263444.

Anonymous:1992:SBS

[200] Anonymous. A section is born: Studies in automated reasoning. Journal of Automated Reasoning, 8(1): 1–2, February 1992. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00263445.

Russinoff:1992:MPQ

[201] David M. Russinoff. A mechanical proof of quadratic reciprocity. Journal of Automated Reasoning, 8(1):3– 21, February 1992. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00263446.

Jindal:1992:EPP

[202] Anita Jindal, Ross Overbeek, and Waldo C. Kabat. Exploitation of parallel processing for implementing high-performance deduction systems. Journal of Automated Reasoning, 8 (1):23–38, February 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00263447.

Domenjoud:1992:TNA

[203] Eric Domenjoud. A technical note on AC-unification. The number of minimal unifiers of the equation αx_{1} + ⋯ + αx_{p} ≐_{AC} βy_{1}+⋯+βy_{q}. Journal of Automated Reasoning, 8(1):39–44, February 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00263448.

Hines:1992:CPD

[204] Larry M. Hines. Completeness of a prover for dense linear orders. Journal of Automated Reasoning, 8(1): 45–75, February 1992. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00263449.

Chen:1992:IRA

[205] Shyi-Ming Chen, Jyh-Sheng Ke, and Jin-Fu Chang. An inexact reasoning algorithm based on fuzzy rule matrix transformations. Journal of Automated Reasoning, 8(1):77–90, February 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00263450.

Quaife:1992:ADN

[206] Art Quaife. Automated deduction in von Neumann–Bernays–Gödel set theory. Journal of Automated Reasoning, 8(1):91–147, February 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00263451.

Bose:1992:PPT

[207] Soumitra Bose, Edmund M. Clarke, David E. Long, and Spiro Michaylov. Parthenon: A parallel theorem prover for non-Horn clauses. Journal of Automated Reasoning, 8(2):153–181, April 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244281.

Letz:1992:SHP

[208] R. Letz, J. Schumann, S. Bayerl, and W. Bibel. SETHEO: A highperformance theorem prover. Journal of Automated Reasoning, 8(2):183–212, April 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244282.

Veroff:1992:LIP

[209] Robert Veroff and Larry Wos. The linked inference principle, i: The formal treatment. Journal of Automated Reasoning, 8(2):213–274, April 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244283.

Baker:1992:ITP

[210] James D. Baker and Shahriar Zand-Biglari. An integral theorem prover and the role of proof planning. Journal of Automated Reasoning, 8(2):275–295, April 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244284.

Shepherdson:1992:SRE

[211] J. C. Shepherdson. SLDNF-resolution with equality. Journal of Automated Reasoning, 8(2):297–306, April 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244285.

Wos:1992:PCB

[212] Larry Wos. The problem of choosing between using and avoiding equality predicates. Journal of Automated Reasoning, 8(2):307–309, April 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244286.

Barker-Plummer:1992:GAP

[213] Dave Barker-Plummer. Gazing: An approach to the problem of definition and lemma use. Journal of Automated Reasoning, 8(3):311–344, June 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02341853.

Baral:1992:SEC

[214] Chitta R. Baral and V. S. Subrahmanian. Stable and extension class theory for logic programs and default logics. Journal of Automated Reasoning, 8(3):345–366, June 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02341854.

Hall:1992:CPS

[215] Robert J. Hall. Comparing parameter schemes for propositional reasoning: An empirical study. Journal of Automated Reasoning, 8(3):367–394, June 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02341855.

Russo:1992:AIA

[216] Mark F. Russo and Richard L. Peskin. Automatically identifying the asymptotic behavior of nonlinear singularly perturbed boundary value problems. Journal of Automated Reasoning, 8(3):395–419, June 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02341856.

Wos:1992:PRI

[217] Larry Wos. The problem of reasoning from inequalities. Journal of Automated Reasoning, 8(3):421–426, June 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02341857.

Mccune:1992:ADN

[218] William W. Mccune. Automated discovery of new axiomatizations of the left group and right group calculi. Journal of Automated Reasoning, 9(1):1–24, August 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00247824.

Lee:1992:EDH

[219] Shie-Jue Lee and David A. Plaisted. Eliminating duplication with the hyper-linking strategy. Journal of Automated Reasoning, 9(1):25–42, August 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247825.

Benanav:1992:RUC

[220] Dan Benanav. Recognizing unnecessary clauses in resolution based systems. Journal of Automated Reasoning, 9(1):43–76, August 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00247826.

Socher-Ambrosius:1992:HAD

[221] Rolf Socher-Ambrosius. How to avoid the derivation of redundant clauses in reasoning systems. Journal of Automated Reasoning, 9(1):77–97, August 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247827.

Olivetti:1992:TSC

[222] Nicola Olivetti. Tableaux and sequent calculus for minimal entailment. Journal of Automated Reasoning, 9 (1):99–139, August 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00247828.

Wos:1992:PDD

[223] Larry Wos. The problem of demodulation during inference rule application. Journal of Automated Reasoning, 9(1):141–143, August 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00247829.

Wos:1992:NMA

[224] Larry Wos. Note on McCune’s article on discrimination trees. Journal of Automated Reasoning, 9(2):145– 146, October 1992. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00245457.

McCune:1992:EDT

[225] William McCune. Experiments with discrimination-tree indexing and path indexing for term retrieval. Journal of Automated Reasoning, 9(2): 147–167, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00245458.

Franzen:1992:HTP

[226] Mark Franzen. Hilbert’s tenth problem is of unification type zero. Journal of Automated Reasoning, 9(2): 169–178, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00245459.

Kifer:1992:LRI

[227] Michael Kifer and Eliezer L. Lozinskii. A logic for reasoning with inconsistency. Journal of Automated Reasoning, 9(2):179–215, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245460.

Brown:1992:UMI

[228] Donald E. Brown and Wendy J. Markert. Uncertainty management with imprecise knowledge with application to design. Journal of Automated Reasoning, 9(2):217–230, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245461.

Heintze:1992:CMR

[229] Nevin Heintze, Spiro Michaylov, and Peter Stuckey. CLP(R) and some electrical engineering problems. Journal of Automated Reasoning, 9(2): 231–260, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00245462.

Kapur:1992:CUP

[230] Deepak Kapur and Paliath Narendran. Complexity of unification problems with associative-commutative operators. Journal of Automated Reasoning, 9(2):261–288, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00245463.

Wos:1992:PDAa

[231] Larry Wos. The problem of demodulator adjunction. Journal of Automated Reasoning, 9(2):289–290, October 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00245464.

Wos:1992:TF

[232] Larry Wos. Transition to the future. Journal of Automated Reasoning, 9 (3):iii, December 1992. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00245292.

Kunen:1992:SAG

[233] Kenneth Kunen. Single axioms for groups. Journal of Automated Reasoning, 9(3):291–308, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245293.

Avron:1992:UTL

[234] Arnon Avron, Furio Honsell, Ian A. Mason, and Robert Pollack. Using typed lambda calculus to implement formal systems on a machine. Journal of Automated Reasoning, 9(3): 309–354, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00245294.

Kaufmann:1992:EBM

[235] Matt Kaufmann. An extension of the Boyer–Moore theorem prover to support first-order quantification. Journal of Automated Reasoning, 9(3): 355–372, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00245295.

Lassez:1992:FAL

[236] Jean-Louis Lassez and Michael J. Maher. On Fourier’s algorithm for linear arithmetic constraints. Journal of Automated Reasoning, 9(3):373–379, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245296.

Kuper:1992:NPC

[237] G. M. Kuper, K. W. McAloon, K. V. Palem, and K. J. Perry. A note on the parallel complexity of antiunification. Journal of Automated Reasoning, 9(3):381–389, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245297.

Delcher:1992:EPT

[238] Arthur L. Delcher and Simon Kasif. Efficient parallel term matching and anti-unification. Journal of Automated Reasoning, 9(3):391–406, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00245298.

Wos:1992:PDAb

[239] Larry Wos. The problem of demodulating across argument and literal boundaries. Journal of Automated Reasoning, 9(3):407–408, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00245299.

Parsons:1992:BR

[240] Simon Parsons. Book review. Journal of Automated Reasoning, 9(3): 409–411, December 1992. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00245300.

McCune:1993:SAG

[241] William W. McCune. Single axioms for groups and Abelian groups with various operations. Journal of Automated Reasoning, 10(1):1–13, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881862.

Noel:1993:EIZ

[242] P. A. J. Noel. Experimenting with Isabelle in ZF set theory. Journal of Automated Reasoning, 10(1): 15–58, February 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881863.

Weigert:1993:FOL

[243] Thomas J. Weigert, Jing-Pha Tsai, and Xuhua Liu. Fuzzy operator logic and fuzzy resolution. Journal of Automated Reasoning, 10(1):59–78, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881864.

Corella:1993:WHC

[244] F. Corella. What holds in a context? Journal of Automated Reasoning, 10 (1):79–93, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881865.

Christian:1993:FDN

[245] Jim Christian. Flatterms, discrimination nets, and fast term rewriting. Journal of Automated Reasoning, 10 (1):95–113, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881866.

Vlach:1993:SSC

[246] Frank Vlach. Simplification in a satisfiability checker for VLSI applications. Journal of Automated Reasoning, 10(1):115–136, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881867.

Wos:1993:PAT

[247] Larry Wos. The problem of automated theorem finding. Journal of Automated Reasoning, 10(1):137– 138, February 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881868.

Yuan:1993:ACL

[248] Li Yan Yuan and Jia-Huai You. Autoepistemic circumscription and logic programming. Journal of Automated Reasoning, 10(2):143–160, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881833.

Chou:1993:ARDa

[249] Shang-Ching Chou and Xiao-Shan Gao. Automated reasoning in differential geometry and mechanics using the characteristic set method. Journal of Automated Reasoning, 10(2):161–172, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881834.

Chou:1993:ARDb

[250] Shang-Ching Chou and Xiao-Shan Gao. Automated reasoning in differential geometry and mechanics using the characteristic set method. Journal of Automated Reasoning, 10(2):173–189, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881835.

Ng:1993:SFS

[251] Raymond Ng and V. S. Subrahmanian. A semantical framework for supporting subjective and conditional probabilities in deductive databases. Journal of Automated Reasoning, 10(2):191–235, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881836.

Suchenek:1993:FOS

[252] Marek A. Suchenek. First-order syntactic characterizations of minimal entailment, domain-minimal entailment, and Herbrand entailment. Journal of Automated Reasoning, 10(2):237–263, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881837.

Avron:1993:GTS

[253] Arnon Avron. Gentzen-type systems, resolution and tableaux. Journal of Automated Reasoning, 10(2):265–281, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881838.

Wos:1993:PSA

[254] Larry Wos. The problem of selecting an approach based on prior success. Journal of Automated Reasoning, 10 (2):283–284, June 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881839.

Ashbacher:1993:BR

[255] Charles Ashbacher. Book review. Journal of Automated Reasoning, 10(2): 285–286, June 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881840.

Wos:1993:KSU

[256] Larry Wos. The kernel strategy and its use for the study of combinatory logic. Journal of Automated Reasoning, 10(3):287–343, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881795.

Lu:1993:IDL

[257] James J. Lu, Monica D. Barback, and Lawrence J. Henschen. Interpreting disjunctive logic programs based on a strong sense of disjunction. Journal of Automated Reasoning, 10(3): 345–370, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881796.

Lu:1993:CIR

[258] James J. Lu and V. S. Subrahmanian. Completeness issues in RUE– NRF deduction: The undecidability of viability. Journal of Automated Reasoning, 10(3):371–388, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881797.

Steinbach:1993:SOP

[259] Joachim Steinbach. Simplification orderings: Putting them to the test. Journal of Automated Reasoning, 10 (3):389–397, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881798.

Baral:1993:DBA

[260] Chitta R. Baral and V. S. Subrahmanian. Dualities between alternative semantics for logic programming and nonmonotonic reasoning. Journal of Automated Reasoning, 10(3): 399–420, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881799.

Wos:1993:PRA

[261] Larry Wos. The problem of reasoning by analogy. Journal of Automated Reasoning, 10(3):421–422, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00881800.

Anonymous:1993:A

[262] Anonymous. Announcement. Journal of Automated Reasoning, 10(3): 423, October 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881801.

Walker:1993:BIT

[263] Adrian Walker. Backchain iteration: Towards a practical inference method that is simple enough to be proved terminating, sound, and complete. Journal of Automated Reasoning, 11(1):1–22, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881898.

Giunchiglia:1993:IIA

[264] Fausto Giunchiglia and Toby Walsh. The inevitability of inconsistent abstract spaces. Journal of Automated Reasoning, 11(1):23–41, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881899.

Felty:1993:ITT

[265] Amy Felty. Implementing tactics and tacticals in a higher-order logic programming language. Journal of Automated Reasoning, 11(1):43–81, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881900.

Segre:1993:BOC

[266] Alberto Segre and Daniel Scharstein. Bounded-overhead caching for definiteclause theorem proving. Journal of Automated Reasoning, 11(1):83– 113, February 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881901.

Nadathur:1993:PPL

[267] Gopalan Nadathur. A proof procedure for the logic of hereditary Harrop formulas. Journal of Automated Reasoning, 11(1):115–145, February 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881902.

Wos:1993:PNF

[268] Larry Wos. The problem of naming and function replacement. Journal of Automated Reasoning, 11(1): 147–148, February 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00881903.

Schulz:1993:WUT

[269] Klaus U. Schulz. Word unification and transformation of generalized equations. Journal of Automated Reasoning, 11(2):149–184, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881904.

Boudet:1993:CAU

[270] Alexandre Boudet. Competing for the AC–Unification race. Journal of Automated Reasoning, 11(2):185–212, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881905.

Farmer:1993:IIM

[271] William M. Farmer, Joshua D. Guttman, and F. Javier Thayer. IMPS: An interactive mathematical proof system. Journal of Automated Reasoning, 11(2):213–248, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881906.

Grant:1993:VUS

[272] John Grant, John Horty, Jorge Lobo, and Jack Minker. View updates in stratified disjunctive databases. Journal of Automated Reasoning, 11(2): 249–267, June 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881907.

Wang:1993:IFH

[273] Jinchang Wang. Inference flexibility in Horn clause knowledge bases and the simplex method. Journal of Automated Reasoning, 11(2):269–288, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881908.

Wos:1993:PRC

[274] Larry Wos. The problem of reasoning by case analysis. Journal of Automated Reasoning, 11(2):289–291, June 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881909.

Bledsoe:1993:SV

[275] W. W. Bledsoe and Guohui Feng. SET-VAR. Journal of Automated Reasoning, 11(3):293–314, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881869.

Overbeek:1993:CCP

[276] Ross Overbeek. The CADE-11 competitions: A personal view. Journal of Automated Reasoning, 11(3): 315–316, October 1993. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF00881870.

Lusk:1993:USC

[277] Ewing L. Lusk and William W. Mc-Cune. Uniform strategies: The CADE-11 theorem proving contest. Journal of Automated Reasoning, 11(3): 317–331, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881871.

Zhang:1993:APE

[278] Hantao Zhang. Automated proofs of equality problems in Overbeek’s competition. Journal of Automated Reasoning, 11(3):333–351, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881872.

Paulson:1993:STV

[279] Lawrence C. Paulson. Set theory for verification: I. From foundations to functions. Journal of Automated Reasoning, 11(3):353–389, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881873.

Bailin:1993:MIR

[280] Sidney C. Bailin and Dave Barker-Plummer. −-match: An inference rule for incrementally elaborating set instantiations. Journal of Automated Reasoning, 11(3):391–428, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881874.

Parsons:1993:BR

[281] Simon Parsons. Book review. Journal of Automated Reasoning, 11(3): 429–431, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881875.

Wos:1993:PI

[282] Larry Wos. The problem of induction. Journal of Automated Reasoning, 11 (3):433–434, October 1993. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF00881876.

Lozinskii:1994:RCP

[283] Eliezer L. Lozinskii. Resolving contradictions: A plausible semantics for inconsistent systems. Journal of Automated Reasoning, 12(1):1–31, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881841.

Moore:1994:IOA

[284] J. Strother Moore. Introduction to the OBDD algorithm for the ATP community. Journal of Automated Reasoning, 12(1):33–45, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881842.

Mellouli:1994:TTS

[285] Taäieb Mellouli. TMPR: A treestructured modified problem reduction proof procedure and its extension to three-valued logic. Journal of Automated Reasoning, 12(1):47–87, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881843.

Benhamou:1994:TTS

[286] Belaid Benhamou and Lakhdar Sais. Tractability through symmetries in propositional calculus. Journal of Automated Reasoning, 12(1):89– 102, February 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881844.

Aitken:1994:RTP

[287] J. Stuart Aitken, Han Reichgelt, and Nigel Shadbolt. Resolution theorem proving in reified modal logics. Journal of Automated Reasoning, 12(1): 103–129, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881845.

Parsons:1994:BR

[288] Simon Parsons. Book review. Journal of Automated Reasoning, 12(1): 131–132, February 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881846.

Wos:1994:PSH

[289] Larry Wos. The problem of strategy and hyperresolution. Journal of Automated Reasoning, 12(1):133– 134, February 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881847.

Butler:1994:FDH

[290] Ralph M. Butler and Ross A. Overbeek. Formula databases for highperformance resolution/paramodulation systems. Journal of Automated Reasoning, 12(2):139–156, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00881885.

Chou:1994:MBB

[291] Timothy S. C. Chou and Marianne Winslett. A model-based belief revision system. Journal of Automated Reasoning, 12(2):157–208, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881886.

Burke:1994:UPC

[292] E. K. Burke. Unification in partially commutative semigroups. Journal of Automated Reasoning, 12(2):209–223, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881887.

Bonacina:1994:SDD

[293] Maria Paola Bonacina and Jieh Hsiang. On subsumption in distributed derivations. Journal of Automated Reasoning, 12(2):225–240, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881888.

Digricoli:1994:RTP

[294] Vincent J. Digricoli. The rue theoremproving system: The complete set of LIM+ challenge problems. Journal of Automated Reasoning, 12(2):241–264, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881889.

Wos:1994:PH

[295] Larry Wos. The problem of hyperparamodulation. Journal of Automated Reasoning, 12(2):265–271, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881890.

Tammet:1994:PSL

[296] Tanel Tammet. Proof strategies in linear logic. Journal of Automated Reasoning, 12(3):273–304, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00885763.

Raschid:1994:SCN

[297] Louiqa Raschid and Jorge Lobo. A semantics for a class of non-deterministic and causal production system programs. Journal of Automated Reasoning, 12(3):305–349, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00885764.

Ferro:1994:EPP

[298] Giuseppa Carra’ Ferro. An extension of a procedure to prove statements in differential geometry. Journal of Automated Reasoning, 12(3):351–358, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00885765.

Lee:1994:IEH

[299] Shie-Jue Lee and Chih-Hung Wu. Improving the efficiency of a hyperlinkingbased theorem prover by incremental evaluation with network structures. Journal of Automated Reasoning, 12 (3):359–388, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00885766.

Qian:1994:RUL

[300] Zhenyu Qian and Tobias Nipkow. Reduction and unification in lambda calculi with a general notion of subtype. Journal of Automated Reasoning, 12 (3):389–406, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00885767.

Wos:1994:PHN

[301] Larry Wos. The problem of hyperparamodulation and nuclei. Journal of Automated Reasoning, 12(3): 407–409, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00885768.

Anonymous:1994:E

[302] Anonymous.  Errata.  Journal

of Automated Reasoning, 12(3):411– 412, October 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/BF00885769.pdf.

Bonacina:1994:PDS

[303] Maria Paola Bonacina and Jieh Hsiang. Parallelization of deduction strategies: An analytical study. Journal of Automated Reasoning, 13(1):1–33, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881910.

Bonatti:1994:ALP

[304] Piero A. Bonatti. Autoepistemic logic programming. Journal of Automated Reasoning, 13(1):35–67, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881911.

Yuan:1994:ALF

[305] Li-Yan Yuan. Autoepistemic logic of first order and its expressive power. Journal of Automated Reasoning, 13 (1):69–82, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881912.

Johnson:1994:TCG

[306] C. A. Johnson. On the termination of clause graph resolution. Journal of Automated Reasoning, 13(1): 83–115, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881913.

Yahya:1994:OMT

[307] Adnan Yahya, José Alberto Fernández, and Jack Minker. Ordered model trees: A normal form for disjunctive deductive databases. Journal of Automated Reasoning, 13(1):117–143, February 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881914.

Sakama:1994:AAS

[308] Chiaki Sakama and Katsumi Inoue. An alternative approach to the semantics of disjunctive logic programs and deductive databases. Journal of Automated Reasoning, 13(1):145– 172, February 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881915.

Wrightson:1994:PSI

[309] Graham Wrightson. Preface to the special issue on tableaux. Journal of Automated Reasoning, 13(2):173, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00881953.

Fitting:1994:TLP

[310] Melvin Fitting. Tableaux for logic programming. Journal of Automated Reasoning, 13(2):175–188, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881954.

Stickel:1994:UMI

[311] Mark E. Stickel. Upside-down metainterpretation of the model elimination theorem-proving procedure for deduction and abduction. Journal of Automated Reasoning, 13(2):189–210, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881955.

Hahnle:1994:LRF

[312] Reiner Hähnle and Peter H. Schmitt. The liberalized δ-rule in free variable semantic tableaux. Journal of Automated Reasoning, 13(2):211–221, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881956.

Risch:1994:TBC

[313] Vincent Risch and Camilla B. Schwind. Tableau-based characterization and theorem proving for default logic. Journal of Automated Reasoning, 13(2): 223–242, June 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881957.

DAgostino:1994:GAD

[314] Marcello D’Agostino and Dov M. Gabbay. A generalization of analytic deduction via labelled deductive systems. Part I: Basic substructural logics. Journal of Automated Reasoning, 13(2):243–281, June 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881958.

Astrachan:1994:MEM

[315] Owen Astrachan. METEOR: Exploring model elimination theorem proving. Journal of Automated Reasoning, 13(3):283–296, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881946.

Letz:1994:CIC

[316] R. Letz, K. Mayr, and C. Goller. Controlled integration of the cut rule into connection tableau calculi. Journal of Automated Reasoning, 13(3):

297–337, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881947.

Baumgartner:1994:MEC

[317] Peter Baumgartner and Ulrich Furbach. Model elimination without contrapositives and its application to PTTP. Journal of Automated Reasoning, 13(3):339–359, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881948.

Miglioli:1994:IRS

[318] Pierangelo Miglioli, Ugo Moscato, and Mario Ornaghi. An improved refutation system for intuitionistic predicate logic. Journal of Automated Reasoning, 13(3):361–373, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881949.

Lorenz:1994:TPD

[319] Sven Lorenz. A tableau prover for domain minimization. Journal of Automated Reasoning, 13(3):375–390, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881950.

Pliuskevicius:1994:STL

[320] Regimantas Pliuškevičius. The saturated tableaux for linear miniscoped Horn-like temporal logic. Journal of Automated Reasoning, 13(3):391– 407, October 1994. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881951.

Schumann:1994:TBT

[321] J. Schumann. Tableau-based theorem provers: Systems and implementations. Journal of Automated Reasoning, 13 (3):409–421, October 1994. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881952.

Lifschitz:1995:PSIa

[322] Vladimir Lifschitz. Preface to the special issue on commonsense and nonmonotonic reasoning. Journal of Automated Reasoning, 14(1):1, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00883928.

Niemela:1995:DMN

[323] Ilkka Niemelä. A decision method for nonmonotonic reasoning based on autoepistemic reasoning. Journal of Automated Reasoning, 14(1):3–42, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00883929.

Ginsberg:1995:MI

[324] Matthew L. Ginsberg. Modality and interrupts. Journal of Automated Reasoning, 14(1):43–91, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00883930.

Alferes:1995:LPS

[325] José Júlio Alferes, Carlos Viegas Damásio, and Luís Moniz Pereira. A logic programming system for nonmonotonic reasoning. Journal of Automated Reasoning, 14(1):93–147, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00883931.

Baader:1995:EDT

[326] Franz Baader and Bernhard Hollunder. Embedding defaults into terminological knowledge representation formalisms. Journal of Automated Reasoning, 14(1):149–180, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00883932.

Bouhoula:1995:IIC

[327] Adel Bouhoula and Michaël Rusinowitch. Implicit induction in conditional theories. Journal of Automated Reasoning, 14(2):189–235, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881856.

Ramsay:1995:TPI

[328] Allan Ramsay. Theorem proving for intensional logic. Journal of Automated Reasoning, 14(2):237–255, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881857.

Chou:1995:APT

[329] Shang-Ching Chou, Xiao-Shan Gao, and Jing-Zhong Zhang. Automated production of traditional proofs in solid geometry. Journal of Automated Reasoning, 14(2):257–291, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881858.

Kapur:1995:POP

[330] Deepak Kapur, G. Sivakumar, and Hantao Zhang. A path ordering for proving termination of AC rewrite systems. Journal of Automated Reasoning, 14(2):293–316, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881859.

Vadaparty:1995:CNO

[331] Kumar V. Vadaparty. On the complexity of nested-object matching. Journal of Automated Reasoning, 14(2): 317–323, June 1995. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881860.

Loveland:1995:SSR

[332] Donald W. Loveland, David W. Reed, and Debra S. Wilson. SATCHMORE: SATCHMO with RElevancy. Journal of Automated Reasoning, 14(2): 325–351, June 1995. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881861.

Schlechta:1995:LTI

[333] Karl Schlechta. Logic, topology, and integration. Journal of Automated Reasoning, 14(3):353–381, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881713.

Hart:1995:SAO

[334] Joan Hart and Kenneth Kunen. Single axioms for odd exponent groups. Journal of Automated Reasoning, 14 (3):383–412, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881714.

Jech:1995:OES

[335] Thomas Jech. OTTER experiments in a system of combinatory logic. Journal of Automated Reasoning, 14 (3):413–426, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881715.

Chakrabarti:1995:IIP

[336] I. Chakrabarti, D. Sarkar, and A. K. Majumdar. Identification of inductive properties during verification of synchronous sequential circuits. Journal of Automated Reasoning, 14(3): 427–462, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881716.

Lifschitz:1995:PSIb

[337] Vladimir Lifschitz. Preface to the special issue on commonsense and nonmonotonic reasoning. Journal of Automated Reasoning, 15(1):v, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00881828.

Thielscher:1995:DRD

[338] Michael Thielscher and Torsten Schaub. Default reasoning by deductive planning. Journal of Automated Reasoning, 15(1):1–40, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881829.

Baader:1995:PDP

[339] Franz Baader and Bernhard Hollunder. Priorities on defaults with prerequisites, and their application in treating specificity in terminological default logic. Journal of Automated Reasoning, 15(1):41–68, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881830.

Subramanian:1995:MVS

[340] Sakthi Subramanian. Mechanical verification of strategies. Journal of Automated Reasoning, 15(1):69–93, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881831.

Schaub:1995:NMQ

[341] T. Schaub. A new methodology for query answering in default logics via structure-oriented theorem proving. Journal of Automated Reasoning, 15 (1):95–165, February 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881832.

Paulson:1995:STV

[342] Lawrence C. Paulson. Set theory for verification. II: Induction and recursion. Journal of Automated Reasoning, 15(2):167–215, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00881916.

Kunen:1995:RTB

[343] Kenneth Kunen. A Ramsey theorem in Boyer–Moore logic. Journal of Automated Reasoning, 15(2):217–235, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881917.

Voronkov:1995:AV

[344] Andrei Voronkov. The anatomy of vampire. Journal of Automated Reasoning, 15(2):237–265, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF00881918.

Ohlbach:1995:NAA

[345] Hans Jürgen Ohlbach and Christoph Weidenbach. A note on assumptions about Skolem functions. Journal of Automated Reasoning, 15(2):267–275, June 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881919.

Anonymous:1995:KLS

[346] Anonymous. The Kluwer L^{A}TEX style file. Journal of Automated Reasoning, 15(2):277–278, June 1995. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF00881920.

Wos:1995:SCP

[347] Larry Wos. Searching for circles of pure proofs. Journal of Automated Reasoning, 15(3):279–315, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00881802.

DAgostino:1995:STT

[348] G. D’Agostino, A. Montanari, and A. Policriti. A set-theoretic translation method for polymodal logics. Journal of Automated Reasoning, 15 (3):317–337, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881803.

Beckert:1995:LTL

[349] Bernhard Beckert and Joachim Posegga. lean TAP: Lean tableau-based deduction. Journal of Automated Reasoning, 15(3):339–358, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00881804.

Hooker:1995:BRS

[350] J. N. Hooker and V. Vinay. Branching rules for satisfiability. Journal of Automated Reasoning, 15(3):359–383, October 1995. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00881805.

Zhang:1996:PSI

[351] Hantao Zhang. Preface to the special issue on automated mathematical induction. Journal of Automated Reasoning, 16(1–2):1, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF00244457.

Bronsard:1996:IUT

[352] François Bronsard, Uday S. Reddy, and Robert W. Hasker. Induction using term orders. Journal of Automated Reasoning, 16(1–2):3–37, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00244458.

Kapur:1996:NUL

[353] Deepak Kapur and M. Subramaniam. New uses of linear arithmetic in automated theorem proving by induction. Journal of Automated Reasoning, 16(1–2):39–78, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00244459.

Ireland:1996:PUF

[354] Andrew Ireland and Alan Bundy. Productive use of failure in inductive proof.

Journal of Automated Reasoning, 16 (1–2):79–111, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244460.

Kraan:1996:MRS

[355] Ina Kraan, David Basin, and Alan Bundy. Middle-out reasoning for synthesis and induction. Journal of Automated Reasoning, 16(1–2):113–145, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244461.

Basin:1996:CTR

[356] David A. Basin and Toby Walsh. A calculus for and termination of rippling. Journal of Automated Reasoning, 16 (1–2):147–180, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244462.

Kaufmann:1996:IBM

[357] Matt Kaufmann and Paolo Pecchiari. Interaction with the Boyer–Moore theorem prover: A tutorial study using the arithmetic–geometric mean theorem. Journal of Automated Reasoning, 16(1–2):181–222, March 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244463.

Veroff:1996:UHI

[358] Robert Veroff. Using hints to increase the effectiveness of an automated reasoning program: Case studies. Journal of Automated Reasoning, 16(3):223–239, June 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00252178.

Baumgartner:1996:LUR

[359] Peter Baumgartner. Linear and unitresulting refutations for Horn theories. Journal of Automated Reasoning, 16(3):241–319, June 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF00252179.

Andrews:1996:TTP

[360] Peter B. Andrews, Matthew Bishop, Sunil Issar, Dan Nesmith, Frank Pfenning, and Hongwei Xi. TPS: A theorem-proving system for classical type theory. Journal of Automated Reasoning, 16(3):321–353, June 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00252180.

Wang:1996:CIC

[361] Lusheng Wang, S. K. M. Wong, and Y. Y. Yao. On the completeness of incidence calculus. Journal of Automated Reasoning, 16(3):355–368, June 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF00252181.

Hedberg:1996:TTI

[362] Michael Hedberg. A type-theoretic interpretation of constructive domain theory. Journal of Automated Reasoning, 16(3):369–425, June 1996.

CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/BF00252182.

Ashbacher:1996:LLP

[363] Charles Ashbacher. From logic to logic programming. Journal of Automated Reasoning, 16(3):427, June 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF00252183.

Zhang:1996:CFA

[364] Jian Zhang. Constructing finite algebras with FALCON. Journal of Automated Reasoning, 17(1):1–22, August 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247667.

Wos:1996:PCR

[365] Larry Wos. The power of combining resonance with heat. Journal of Automated Reasoning, 17(1):23–81, August 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247668.

Kunen:1996:SAL

[366] Kenneth Kunen. The semantics of answer literals. Journal of Automated Reasoning, 17(1):83–95, August 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247669.

Codognet:1996:SEB

[367] Philippe Codognet and Daniel Diaz. A simple and efficient boolean solver for constraint logic programming. Journal of Automated Reasoning, 17(1): 97–128, August 1996. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00247670.

Narendran:1996:GAC

[368] Paliath Narendran and Michaël Rusinowitch. Any ground associativecommutative theory has a finite canonical system. Journal of Automated Reasoning, 17(1):131–143, August 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00247671.

Linton:1996:SGT

[369] S. Linton and D. Shand. Some group theoretic examples with completion theorem provers. Journal of Automated Reasoning, 17(2):145–169, October 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244494.

Fernandez:1996:MTA

[370] José Alberto Fernández, John Grant, and Jack Minker. Model theoretic approach to view updates in deductive databases. Journal of Automated Reasoning, 17(2):171–197, October 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244495.

Thielscher:1996:CSR

[371] Michael Thielscher. On the completeness of SLDENF-resolution. Journal of Automated Reasoning, 17(2): 199–214, October 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00244496.

Wos:1996:OMI

[372] Larry Wos. OTTER and the Moufang identity problem. Journal of Automated Reasoning, 17(2):215–257, October 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00244497.

Dimopoulos:1996:CLP

[373] Yannis Dimopoulos. On computing logic programs. Journal of Automated Reasoning, 17(3):259–289, December 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00283131.

Paulson:1996:MST

[374] Lawrence C. Paulson and Krzysztof Grabczewski. Mechanizing set theory. Journal of Automated Reasoning, 17 (3):291–323, December 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF00283132.

Chou:1996:AGRa

[375] Shang-Ching Chou, Xiao-Shan Gao, and Jing-Zhong Zhang. Automated generation of readable proofs with geometric invariants. Journal of Automated Reasoning, 17(3):325–347, December 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00283133.

Chou:1996:AGRb

[376] Shang-Ching Chou, Xiao-Shan Gao, and Jing-Zhong Zhang. Automated generation of readable proofs with geometric invariants. Journal of Automated Reasoning, 17(3):349–370, December 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00283134.

Sidebottom:1996:NSY

[377] Gregory Sidebottom and William S. Havens. Nicolog: A simple yet powerful cc(FD) language. Journal of Automated Reasoning, 17(3):371–403, December 1996. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF00283135.

Hines:1997:TWB

[378] Larry Hines. A tribute to Woody Bledsoe. Journal of Automated Reasoning, 18(1):1–4, February 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005712718688.

Hou:1997:NPI

[379] Rong-Huei Hou, Tzung-Pei Hong, Shian-Shyong Tseng, and Sy-Yen Kuo. A new probabilistic induction method. Journal of Automated Reasoning, 18 (1):5–24, February 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005726727996.

Antoniou:1997:CLP

[380] Grigoris Antoniou and Elmar Langetepe. A correct logic programming computation of default logic extensions. Journal of Automated Reasoning, 18(1): 25–46, February 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005771523328.

Ramesh:1997:FSC

[381] Anavai Ramesh, Bernhard Beckert, Reiner Hähnle, and Neil V. Murray. Fast subsumption checks using antilinks. Journal of Automated Reasoning, 18(1):47–83, February 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1005711712356.

Finn:1997:PFT

[382] Simon Finn, Michael P. Fourman, and John Longley. Partial functions in a total setting. Journal of Automated Reasoning, 18(1):85– 104, February 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005702928286.

Dafa:1997:UAE

[383] Li Dafa. Unification algorithms for eliminating and introducing quantifiers in natural deduction automated theorem proving. Journal of Automated Reasoning, 18(1):105– 134, February 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005749401809.

Pelletier:1997:ESE

[384] Francis Jeffry Pelletier and Geoff Sutcliffe. An erratum for some errata to ATP problems. Journal of Automated Reasoning, 18(1):135, February 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1005764705033.

Sutcliffe:1997:CAS

[385] Geoff Sutcliffe and Christian Suttner. The CADE-13 ATP system competition. Journal of Automated Reasoning, 18(2):137–138, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1005839515219.

Suttner:1997:DCA

[386] Christian Suttner and Geoff Sutcliffe. The design of the CADE-13 ATP system competition. Journal of Automated Reasoning, 18(2): 139–162, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005802523220.

Sutcliffe:1997:PCA

[387] Geoff Sutcliffe and Christian Suttner. The procedures of the CADE-13 ATP system competition. Journal of Automated Reasoning, 18(2): 163–169, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005858625038.

Nieuwenhuis:1997:B

[388] Robert Nieuwenhuis, José Miguel Rivero, and Miguel Ángel Vallejo. Barcelona. Journal of Automated Reasoning, 18(2):171–176, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1023/A%3A1005862710017.

Alexander:1997:CSI

[389] Geoffrey D. Alexander. CLIN-E — smallest instance first hyper-linking. Journal of Automated Reasoning, 18 (2):177–182, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005818927764.

Chu:1997:CSG

[390] Heng Chu and David A. Plaisted. CLIN-S — a semantically guided first-order theorem prover. Journal of Automated Reasoning, 18(2): 183–188, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005875128672.

Denzinger:1997:DDL

[391] Jörg Denzinger, Martin Kronenburg, and Stephan Schulz. DIS-COUNT — a distributed and learning equational prover. Journal of Automated Reasoning, 18(2):189–198, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005879229581.

Tammet:1997:G

[392] Tanel Tammet. Gandalf. Journal of Automated Reasoning, 18(2): 199–204, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005887414560.

Letz:1997:LLI

[393] Reinhold Letz. LINUS — a link instantion prover with unit support. Journal of Automated Reasoning, 18(2): 205–210, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005839531398.

McCune:1997:OCC

[394] William McCune and Larry Wos. Otter- the CADE-13 competition incarnations. Journal of Automated Reasoning, 18(2):211–220, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1023/A%3A1005843632307.

Paramasivam:1997:RRR

[395] M. Paramasivam and David A. Plaisted. RRTP — a replacement rule theorem prover. Journal of Automated Reasoning, 18(2):221–226, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005847700447.

Geisler:1997:SCF

[396] Tim Geisler, Sven Panne, and Heribert Schütz. Satchmo- the compiling and functional variants. Journal of Automated Reasoning, 18(2): 227–236, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005851801356.

Moser:1997:SSC

[397] Max Moser, Ortrun Ibens, Reinhold Letz, Joachim Steinbach, Christoph Goller, Johann Schumann, and Klaus Mayr. SETHEO and E-SETHEO — the CADE-13 systems. Journal of Automated Reasoning, 18(2): 237–246, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005808119103.

Weidenbach:1997:SV

[398] Christoph Weidenbach. SPASS — version 0.49. Journal of Automated Reasoning, 18(2):247–252, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005812220011.

Suttner:1997:SPT

[399] Christian B. Suttner. SPTHEO — a parallel theorem prover. Journal of Automated Reasoning, 18(2): 253–258, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005868304990.

Greenbaum:1997:V

[400] Steve Greenbaum. Violet. Journal of Automated Reasoning, 18(2): 259–264, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005820421829.

Hillenbrand:1997:WHP

[401] Thomas Hillenbrand, Arnim Buch, and Roland Vogt. WALDMEISTER — high-performance equational deduction. Journal of Automated Reasoning, 18(2):265–270, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1005872405899.

Sutcliffe:1997:RCA

[402] Geoff Sutcliffe and Christian Suttner. The results- of the CADE-13 ATP system competition. Journal of Automated Reasoning, 18(2): 271–286, April 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005824522737.

Pelletier:1997:CAC

[403] Francis Jeffry Pelletier, Geoff Sutcliffe, and Christian Suttner. Conclusions about the CADE-13 ATP system competition. Journal of Automated Reasoning, 18(2):287–296, April 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005876506808.

Doherty:1997:CCR

[404] Patrick Doherty, Witold Łukaszewicz, and Andrzej Szałas. Computing circumscription revisited: A reduction algorithm. Journal of Automated Reasoning, 18(3):297–336, June 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005722130532.

Ramesh:1997:CDC

[405] Anavai Ramesh, George Becker, and Neil V. Murray. CNF and DNF considered harmful for computing prime Implicants/ implicates. Journal of Automated Reasoning, 18(3): 337–356, June 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005721905269.

Suchenek:1997:EQU

[406] Marek A. Suchenek. Evaluation of queries under closed-world assumption. Journal of Automated Reasoning, 18 (3):357–398, June 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005723423016.

Hutter:1997:CTC

[407] Dieter Hutter. Coloring terms to control equational reasoning. Journal of Automated Reasoning, 18(3): 399–442, June 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005772217686.

Giesl:1997:TNM

[408] Jürgen Giesl. Termination of nested and mutually recursive algorithms. Journal of Automated Reasoning, 19 (1):1–29, August 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005797629953.

Johnson:1997:DTV

[409] C. A. Johnson. Deduction trees and the view update problem in indefinite deductive databases. Journal of Automated Reasoning, 19(1): 31–85, August 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005779507994.

Narendran:1997:SVS

[410] Paliath Narendran and Friedrich Otto. Single versus simultaneous equational unification and equational unification for variable-permuting theories. Journal of Automated Reasoning, 19(1): 87–115, August 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005764526878.

Astrachan:1997:ULM

[411] O. L. Astrachan and D. W. Loveland. The use of lemmas in the model elimination procedure. Journal of Automated Reasoning, 19(1): 117–141, August 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005770705587.

Corsini:1997:TCF

[412] Marc-Michel Corsini and Antoine Rauzy. Toupie: The μ-calculus over finite domains as a constraint language. Journal of Automated Reasoning, 19 (2):143–171, October 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1005732517165.

Horwein:1997:SRP

[413] K. Hörwein. Structuring resolution proofs by introducing new lemmata. Journal of Automated Reasoning, 19 (2):173–203, October 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1005890730317.

Baumgartner:1997:DPR

[414] Peter Baumgartner and Stefan Brüning. A disjunctive positive refinement of model elimination and its application to subsumption deletion. Journal of Automated Reasoning, 19(2): 205–262, October 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005812703468.

Mccune:1997:SRP

[415] William Mccune. Solution of the Robbins problem. Journal of Automated Reasoning, 19(3):263–276, December 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005843212881.

Caferra:1997:NTV

[416] Ricardo Caferra and Nicolas Peltier. A new technique for verifying and correcting logic programs. Journal of Automated Reasoning, 19(3):277– 318, December 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005878609884.

Bundy:1997:APC

[417] Alan Bundy, Fausto Giunchiglia, Adolfo Villafiorita, and Toby Walsh. Abstract proof checking: An example motivated by an incompleteness theorem. Journal of Automated Reasoning, 19(3):319–346, December 1997. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005877613942.

Sturgill:1997:NDA

[418] David Sturgill and Alberto Maria Segre. Nagging: A distributed, adversarial search-pruning technique applied to first-order inference. Journal of Automated Reasoning, 19(3):347– 376, December 1997. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005885725562.

Alferes:1998:P

[419] José Júlio Alferes and Luís Moniz Pereira. Preface. Journal of Automated Reasoning, 20(1–2):1–3, April 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005925305967.

deNivelle:1998:ARU

[420] Hans de Nivelle. An algorithm for the retrieval of unifiers from discrimination trees. Journal of Automated Reasoning, 20(1–2):5–25, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005992522805.

Defourneaux:1998:SGP

[421] Gilles Défourneaux, Christophe Bourely, and Nicolas Peltier. Semantic generalizations for proving and disproving conjectures by analogy. Journal of Automated Reasoning, 20(1– 2):27–45, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005944606876.

Degtyarev:1998:WYA

[422] Anatoli Degtyarev and Andrei Voronkov. What you always wanted to know about rigid E-unification. Journal of Automated Reasoning, 20(1– 2):47–80, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005996623714.

Frohlich:1998:SMB

[423] Peter Fröhlich, Wolfgang Nejdl, and Michael Schroeder. Strategies in model-based diagnosis. Journal of Automated Reasoning, 20(1–2):81–105, April 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005948807784.

Alferes:1998:CNN

[424] José Júlio Alferes, Luís Moniz Pereira, and Teodor C. Przymusinski. ’classical’ negation in nonmonotonic reasoning and logic programming. Journal of Automated Reasoning, 20(1–2): 107–142, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005900924623.

Brass:1998:CDW

[425] Stefan Brass and Jürgen Dix. Characterizations of the disjunctive wellfounded semantics: Confluent calculi and iterated GCWA. Journal of Automated Reasoning, 20(1–2): 143–165, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005952908693.

Forcheri:1998:ILT

[426] Paola Forcheri, Paolo Gentilini, and Maria Teresa Molfino. Informational logic as a tool for automated reasoning. Journal of Automated Reasoning, 20(1– 2):167–190, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005905025531.

Besnard:1998:SSP

[427] Ph. Besnard and T. Schaub. Signed systems for paraconsistent reasoning. Journal of Automated Reasoning, 20(1– 2):191–213, April 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005957026440.

Huang:1998:SSS

[428] Guoxiang Huang and Dale Myers. Subgoal strategies for solving board puzzles. Journal of Automated Reasoning, 20(3):215–253, June 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005882724254.

Melis:1998:HBC

[429] Erica Melis. The Heine–Borel challenge problem. In honor of Woody Bledsoe. Journal of Automated Reasoning, 20 (3):255–282, June 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005843328643.

Hodgson:1998:SSA

[430] Kahlil Hodgson. Shortest single axioms for the equivalential calculus with CD and RCD. Journal of Automated Reasoning, 20(3): 283–316, June 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005731217123.

Ohlbach:1998:ESR

[431] Hans Jürgen Ohlbach. Elimination of self-resolving clauses. Journal of Automated Reasoning, 20(3): 317–336, June 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005894718064.

Paramasivam:1998:ADT

[432] M. Paramasivam and David A. Plaisted. Automated deduction techniques for classification in description logic systems. Journal of Automated Reasoning, 20(3):337–364, June 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005866922570.

Rody:1998:TDS

[433] Abdelilah Kandri Rody, Hamid Maârouf, and Mohamed Ssafini. Triviality and dimension of a system of algebraic differential equations. Journal of Automated Reasoning, 20(3): 365–385, June 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005821729260.

Anonymous:1998:EIC

[434] Anonymous. The Eleventh International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems (IEA/AIE–98). Journal of Automated Reasoning, 20(3):387, June 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1017193831844.

Hongbo:1998:CAR

[435] Li Hongbo and Cheng Minteh. Clifford algebraic reduction method for automated theorem proving in differential geometry. Journal of Automated Reasoning, 21(1):1–21, August 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005819428156.

Hong:1998:TPP

[436] Hoon Hong and Dalibor Jakuš. Testing positiveness of polynomials. Journal of Automated Reasoning, 21(1): 23–38, August 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005983105493.

Lu:1998:FAR

[437] James J. Lu, Neil V. Murray, and Erik Rosenthal. A framework for automated reasoning in multiple-valued logics. Journal of Automated Reasoning, 21(1):39–67, August 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1005784309139.

Kunen:1998:NCM

[438] Kenneth Kunen. Nonconstructive computational mathematics. Journal of Automated Reasoning, 21(1): 69–97, August 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005888712422.

Suttner:1998:CAS

[439] Christian Suttner and Geoff Sutcliffe. The CADE-14 ATP system competition. Journal of Automated Reasoning, 21(1):99–134, August 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1006006930186.

Wos:1998:ASE

[440] Larry Wos. Automating the search for elegant proofs. Journal of Automated Reasoning, 21(2):135–175, October 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005847113370.

Sutcliffe:1998:TPL

[441] Geoff Sutcliffe and Christian Suttner. The TPTP problem library. Journal of Automated Reasoning, 21(2): 177–203, October 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005806324129.

Voronkov:1998:PSI

[442] Andrei Voronkov. Proof search in intuitionistic logic with equality, or back to simultaneous rigid E-unification. Journal of Automated Reasoning, 21(2): 205–231, October 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1005934721802.

Chazarain:1998:ASR

[443] Jacques Chazarain and Serge Muller. Automated synthesis of recursive programs from a ∀∃ logical specification. Journal of Automated Reasoning, 21 (2):233–275, October 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1005903504159.

Anonymous:1998:A

[444] Anonymous. Announcements. Journal of Automated Reasoning, 21(2): 277, October 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1017187209418.

Kapur:1998:P

[445] Deepak Kapur and Dongming Wang. Preface. Journal of Automated Reasoning, 21(3):277–278, December 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1006075111638.

Harrison:1998:SAC

[446] John Harrison and L. Théry. A skeptic’s approach to combining HOL and Maple. Journal of Automated Reasoning, 21(3):279–294, December 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006023127567.

Bauer:1998:AEC

[447] Andrej Bauer, Edmund Clarke, and Xudong Zhao. Analytica– an experiment in combining theorem proving and symbolic computation. Journal of Automated Reasoning, 21(3):295– 325, December 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006079212546.

Kerber:1998:ICA

[448] Manfred Kerber, Michael Kohlhase, and Volker Sorge. Integrating computer algebra into proof planning. Journal of Automated Reasoning, 21(3):327– 355, December 1998. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006059810729.

Dolzmann:1998:NAA

[449] Andreas Dolzmann, Thomas Sturm, and Volker Weispfenning. A new approach for automatic theorem proving in real geometry. Journal of Automated Reasoning, 21(3):357–380, December 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006031329384.

Wu:1998:WBI

[450] Jinzhao Wu and Zhuojun Liu. Wellbehaved inference rules for first-order theorem proving. Journal of Automated Reasoning, 21(3):381–400, December 1998. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006027228476.

Wos:1999:HLS

[451] Larry Wos and Gail W. Pieper. The hot list strategy. Journal of Automated Reasoning, 22(1):1–44, January 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005909914693.

Stolzenburg:1999:AGS

[452] Frieder Stolzenburg. An algorithm for general set unification and its complexity. Journal of Automated Reasoning, 22(1):45–63, January 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006019227709.

Madden:1999:RPO

[453] Peter Madden, Alan Bundy, and Alan Smaill. Recursive program optimization through inductive synthesis proof transformation. Journal of Automated Reasoning, 22(1):65– 115, January 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005969312327.

Melis:1999:AIT

[454] Erica Melis and Jon Whittle. Analogy in inductive theorem proving. Journal of Automated Reasoning, 22(2): 117–147, February 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005936130801.

Shiny:1999:EAG

[455] A. K. Shiny and Arun K. Pujari. An efficient algorithm to generate prime implicants. Journal of Automated Reasoning, 22(2):149–170, February 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005940031099.

Aitken:1999:MFI

[456] William E. Aitken, Robert L. Constable, and Judith L. Underwood. Metalogical frameworks II: Developing a reflected decision procedure. Journal of Automated Reasoning, 22(2): 171–221, February 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1005929703675.

Lukaszewicz:1999:TDF

[457] Leon Łukaszewicz. Triple dots in a formal language. Journal of Automated Reasoning, 22(3):223–239, March 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006086330571.

Pesant:1999:RAS

[458] Gilles Pesant and Michel Boyer. Reasoning about solids using constraint logic programming. Journal of Automated Reasoning, 22(3):241–262, March 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006080931326.

Arieli:1999:MTA

[459] Ofer Arieli and Arnon Avron. A model-theoretic approach for recovering consistent data from inconsistent knowledge bases. Journal of Automated Reasoning, 22(3):263–309, March 1999. CODEN JAREEW.

ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006020529663.

Belinfante:1999:CPG

[460] Johan G. F. Belinfante. Computer proofs in Gödel’s class theory with equational definitions for composite and cross. Journal of Automated Reasoning, 22(3):311–339, March 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006050629424.

Belinfante:1999:CAP

[461] Johan G. F. Belinfante. On computerassisted proofs in ordinal number theory. Journal of Automated Reasoning, 22(3):341–378, March 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006010913494.

Schmidt:1999:DRP

[462] Renate A. Schmidt. Decidability by resolution for propositional modal logics. Journal of Automated Reasoning, 22(4):379–396, May 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006043519663.

Suttner:1999:SPS

[463] Christian B. Suttner. SPS-Parallelism + SETHEO = SPTHEO. Journal of Automated Reasoning, 22(4): 397–431, May 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006127022180.

Formisano:1999:RRM

[464] Andrea Formisano and Alberto Policriti. T -resolution: Refinements and model elimination. Journal of Automated Reasoning, 22(4):433– 483, May 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006170514174.

Anonymous:1999:AIV

[465] Anonymous. Author index to volume 22. Journal of Automated Reasoning, 22(4):485–486, May 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1017148420906.

Anonymous:1999:SIV

[466] Anonymous. Subject index to volume 22. Journal of Automated Reasoning, 22(4):487, May 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1017100504977.

Anonymous:1999:CV

[467] Anonymous. Contents to volume 22. Journal of Automated Reasoning, 22 (4):489–490, May 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1017144304068.

Sutcliffe:1999:CAS

[468] G. Sutcliffe and C. B. Suttner. The CADE-15 ATP system competition.

Journal of Automated Reasoning, 23 (1):1–23, July 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006285423991.

Kurihara:1999:CMR

[469] Masahito Kurihara and Hisashi Kondo. Completion for multiple reduction orderings. Journal of Automated Reasoning, 23(1):25–42, July 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006129631807.

Ramsdell:1999:TRS

[470] John D. Ramsdell. The tailrecursive SECD machine. Journal of Automated Reasoning, 23(1):43– 62, July 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006151910336.

Recio:1999:ADT

[471] T. Recio and M. P. Vélez. Automatic discovery of theorems in elementary geometry. Journal of Automated Reasoning, 23(1):63–82, July 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006135322108.

Brodsky:1999:SPO

[472] Alexander Brodsky, Catherine Lassez, Jean-Louis Lassez, and Michael J. Maher. Separability of polyhedra for optimal filtering of spatial and constraint data. Journal of Automated Reasoning, 23(1):83–104, July 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006171919920.

Anonymous:1999:A

[473] Anonymous. Announcements. Journal of Automated Reasoning, 23(1):105, July 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1017225615873.

Hermann:1999:CCS

[474] Miki Hermann and Phokion G. Kolaitis. Computational complexity of simultaneous elementary matching problems. Journal of Automated Reasoning, 23(2):107–136, August 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1006136609427.

VanGelder:1999:APP

[475] Allen Van Gelder. Autarky pruning in propositional model elimination reduces failure redundancy. Journal of Automated Reasoning, 23(2): 137–193, August 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006143621319.

Anonymous:1999:P

[476] Anonymous. Preface. Journal of Automated Reasoning, 23(3):195–196, November 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1017202010519.

Rudnicki:1999:EWF

[477] Piotr Rudnicki and Andrzej Trybulec. On equivalents of well-foundedness. Journal of Automated Reasoning, 23 (3):197–234, November 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006218513245.

Kammuller:1999:FPS

[478] Florian Kammüller and Lawrence C. Paulson. A formal proof of Sylow’s theorem. Journal of Automated Reasoning, 23(3):235–264, November 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006269330992.

Chou:1999:FVP

[479] Ching-Tsun Chou and Doron Peled. Formal verification of a partialorder reduction technique for model checking. Journal of Automated Reasoning, 23(3):265–298, November 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006225515062.

Naraschewski:1999:TIV

[480] Wolfgang Naraschewski and Tobias Nipkow. Type inference verified: Algorithm W in Isabelle/HOL. Journal of Automated Reasoning, 23(3):299– 318, November 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006277616879.

Dubois:1999:CTI

[481] Catherine Dubois and Valérie Ménissier-Morain. Certification of a type inference tool for ML: Damas–Milner within Coq. Journal of Automated Reasoning, 23(3):319–346, November 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006285817788.

Jaume:1999:FFS

[482] M. Jaume. A full formalization of SLD–Resolution in the calculus of inductive constructions. Journal of Automated Reasoning, 23(3):347– 371, November 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006242018697.

McKinna:1999:SLC

[483] James McKinna and Robert Pollack. Some Lambda calculus and type theory formalized. Journal of Automated Reasoning, 23(3):373– 409, November 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006294005493.

Reus:1999:FSD

[484] Bernhard Reus. Formalizing synthetic domain theory. Journal of Automated Reasoning, 23(3):411–444, November 1999. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006258506401.

Goldschlag:1999:MUP

[485] David M. Goldschlag. A mechanization of unity in PC–NQTHM-92. Journal of Automated Reasoning, 23(3): 445–498, November 1999. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006262609127.

Gent:2000:SY

[486] Ian P. Gent and Toby Walsh. Satisfiability in the year 2000. Journal of Automated Reasoning, 24(1–2): 1–3, February 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006338702529.

Abramovici:2000:SSU

[487] Miron Abramovici and Jose T. De Sousa. A SAT solver using reconfigurable hardware and virtual logic. Journal of Automated Reasoning, 24(1– 2):5–36, February 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006310219368.

deKlerk:2000:RSP

[488] Etienne de Klerk, Hans van Maaren, and Joost P. Warners. Relaxations of the satisfiability problem using semidefinite programming. Journal of Automated Reasoning, 24(1–2):37– 65, February 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006362203438.

Gomes:2000:HTP

[489] Carla P. Gomes, Bart Selman, Nuno Crato, and Henry Kautz. Heavy-tailed phenomena in satisfiability and constraint satisfaction problems. Journal of Automated Reasoning, 24(1–2): 67–100, February 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006314320276.

Groote:2000:PFC

[490] Jan Friso Groote and Joost P. Warners. The propositional formula checker HeerHugo. Journal of Automated Reasoning, 24(1–2):101–125, February 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006366304347.

Hirsch:2000:SLS

[491] Edward A. Hirsch. SAT local search algorithms: Worst-case study. Journal of Automated Reasoning, 24(1–2): 127–143, February 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006318521185.

Kuchlin:2000:PCA

[492] Wolfgang Küchlin and Carsten Sinz. Proving consistency assertions for automotive product data management. Journal of Automated Reasoning, 24(1– 2):145–163, February 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006370506164.

Massacci:2000:LCS

[493] Fabio Massacci and Laura Marraro. Logical cryptanalysis as a SAT problem. Journal of Automated Reasoning, 24(1–2):165–203, February 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006326723002.

Mills:2000:GLS

[494] Patrick Mills and Edward Tsang. Guided local search for solving SAT and weighted MAX–SAT problems. Journal of Automated Reasoning, 24(1– 2):205–223, February 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006343127545.

Rish:2000:RVS

[495] Irina Rish and Rina Dechter. Resolution versus search: Two strategies for SAT. Journal of Automated Reasoning, 24(1–2):225–275, February 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006303512524.

Zhang:2000:IDP

[496] Hantao Zhang and Mark Stickel. Implementing the Davis–Putnam method. Journal of Automated Reasoning, 24(1– 2):277–296, February 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1006351428454.

Balsiger:2000:BMP

[497] Peter Balsiger, Alain Heuerding, and Stefan Schwendimann. A benchmark method for the propositional modal logics K, KT, S4. Journal of Automated Reasoning, 24(3): 297–317, April 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006249507577.

Massacci:2000:SST

[498] Fabio Massacci. Single step tableaux for modal logics. Journal of Automated Reasoning, 24(3):319–364, April 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006155811656.

Padmanabhan:2000:HBL

[499] R. Padmanabhan and P. Penner. A hyperbase for binary lattice hyperidentities. Journal of Automated Reasoning, 24(3):365–370, April 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1006279127474.

Sutcliffe:2000:CAS

[500] Geoff Sutcliffe. The CADE-16 ATP system competition. Journal of Automated Reasoning, 24(3):371–396, April 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006393501098.

Hirsch:2000:NWC

[501] Edward A. Hirsch. New worstcase upper bounds for SAT. Journal of Automated Reasoning, 24(4): 397–420, May 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006340920104.

Hoos:2000:LSA

[502] Holger H. Hoos and Thomas Stützle. Local search algorithms for SAT: An empirical evaluation. Journal of Automated Reasoning, 24(4):421– 481, May 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006350622830.

Richards:2000:NSN

[503] E. Thomas Richards and Barry Richards. Nonsystematic search and no-good learning. Journal of Automated Reasoning, 24(4):483–533, May 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006362926464.

Spencer:2000:EAD

[504] Bruce Spencer and J. D. Horton. Efficient algorithms to detect and Restore minimality, an extension of the regular restriction of resolution. Journal of Automated Reasoning, 25 (1):1–34, July 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006241826565.

Bry:2000:PUH

[505] François Bry and Adnan Yahya. Positive unit hyperresolution tableaux and their application to minimal model generation. Journal of Automated Reasoning, 25(1):35–82, July 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1006291616338.

Li:2000:VES

[506] Hongbo Li. Vectorial equations solving for mechanical geometry theorem proving. Journal of Automated Reasoning, 25(2):83–121, August 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1006182023017.

Hutter:2000:MSI

[507] Dieter Hutter and Michael Kohlhase. Managing structural information by higher-order colored unification. Journal of Automated Reasoning, 25(2): 123–164, August 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006282725324.

Anonymous:2000:E

[508] Anonymous.  Erratum.  Jour-

nal of Automated Reasoning, 25(2): 165, August 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1023/A%3A1017253718896.pdf.

Plaisted:2000:OSH

[509] David A. Plaisted and Yunshan Zhu. Ordered semantic hyper-linking. Journal of Automated Reasoning, 25(3): 167–217, October 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006376231563.

Chou:2000:DDA

[510] Shang-Ching Chou, Xiao-Shan Gao, and Jing-Zhong Zhang. A deductive database approach to automated geometry theorem proving and discovering. Journal of Automated Reasoning, 25(3):219–246, October 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1006171315513.

Suchenek:2000:EQU

[511] Marek A. Suchenek. Evaluation of queries under closed-world assumption. Part II: The hierarchical case. Journal of Automated Reasoning, 25(4): 247–289, November 2000. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006319819647.

Sistla:2000:RAQ

[512] A. Prasad Sistla and Clement Yu. Reasoning about qualitative spatial relationships. Journal of Automated Reasoning, 25(4):291–328, November 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006322417869.

Anonymous:2000:AIV

[513] Anonymous. Author index to volume 25. Journal of Automated Reasoning, 25(4):329, November 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1017299103703.

Anonymous:2000:SIV

[514] Anonymous. Subject index to volume 25. Journal of Automated Reasoning, 25(4):331–332, November 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1017251220541.

Anonymous:2000:CV

[515] Anonymous. Contents to volume 25. Journal of Automated Reasoning, 25(4):333–334, November 2000. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1017296219633.

Giesl:2001:IPP

[516] Jürgen Giesl. Induction proofs with partial functions. Journal of Automated Reasoning, 26(1):1–49, January 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006408829523.

Nipkow:2001:MCR

[517] Tobias Nipkow. More Church– Rosser proofs. Journal of Automated Reasoning, 26(1):51–66, January 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006496715975.

Wu:2001:PHL

[518] Chih-Hung Wu and Shie-Jue Lee. Parallelization of a hyper-linking-based theorem prover. Journal of Automated Reasoning, 26(1):67–106, January 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006469202251.

Thery:2001:MCI

[519] Laurent Théry. A machine-checked implementation of Buchberger’s algorithm. Journal of Automated Reasoning, 26(2):107–137, February 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1026518331905.

Gehrke:2001:CAC

[520] Wolfgang Gehrke and Jochen Pfalzgraf. Computer-aided construction of finite geometric spaces: Automated verification of geometric constraints. Journal of Automated Reasoning, 26(2): 139–160, February 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1026477216294.

Kaufmann:2001:STD

[521] Matt Kaufmann and J. Strother Moore. Structured theory development for a mechanized logic. Journal of Automated Reasoning, 26(2): 161–203, February 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1026517200045.

Berger:2001:WAD

[522] Ulrich Berger, Helmut Schwichtenberg, and Monika Seisenberger. The Warshall algorithm and Dickson’s lemma: Two examples of realistic program extraction. Journal of Automated Reasoning, 26(2):205– 221, February 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1026748613865.

Avenhaus:2001:GAP

[523] Jürgen Avenhaus and David A. Plaisted. General algorithms for permutations in equational inference. Journal of Automated Reasoning, 26 (3):223–268, April 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006439522342.

Farmer:2001:SST

[524] William M. Farmer. STMM: A set theory for mechanized mathematics. Journal of Automated Reasoning, 26 (3):269–289, April 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1006437704595.

Giunchiglia:2001:RT

[525] F. Giunchiglia, P. Pecchiari, and C. Talcott. Reasoning theories. Journal of Automated Reasoning, 26(3): 291–331, April 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1006407208923.

Johnson:2001:CDW

[526] C. A. Johnson. On the computation of the disjunctive well-founded semantics. Journal of Automated Reasoning, 26(4):333–356, May 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010744127526.

Holmes:2001:WTP

[527] M. Randall Holmes and Jim Alves-Foss. The Watson theorem prover. Journal of Automated Reasoning, 26 (4):357–408, May 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010652212067.

McNaughton:2001:STS

[528] Robert McNaughton. Semi-Thue systems with an inhibitor. Journal of Automated Reasoning, 26(4): 409–431, May 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010759024900.

Anonymous:2001:AIVa

[529] Anonymous. Author index to volume 26. Journal of Automated Reasoning, 26(4):433, May 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1017344401104.

Anonymous:2001:SIVa

[530] Anonymous. Subject index to volume 26. Journal of Automated Reasoning, 26(4):435, May 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1017396417942.

Anonymous:2001:CVa

[531] Anonymous. Contents to volume 26. Journal of Automated Reasoning, 26 (4):437–438, May 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1017343017034.

Basin:2001:CTL

[532] David Basin and Amy Felty. Current trends in logical frameworks and metalanguages. Journal of Automated Reasoning, 27(1):1–2, July 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1010624702918.

Callaghan:2001:ILC

[533] Paul Callaghan and Zhaohui Luo. An implementation of LF with coercive subtyping & universes. Journal of Automated Reasoning, 27(1):3– 27, July 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010648911114.

Bognar:2001:CLC

[534] Mirna Bognar and Roel de Vrijer. A calculus of Lambda calculus contexts. Journal of Automated Reasoning, 27 (1):29–59, July 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010654904735.

Severi:2001:STS

[535] Paula Severi and Nora Szasz. Studies of a theory of specifications with built-in program extraction. Journal of Automated Reasoning, 27(1): 61–87, July 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010663224299.

Wos:2001:MRS

[536] Larry Wos. A milestone reached and a secret revealed. Journal of Automated Reasoning, 27(2):89–95, August 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010650624155.

Ulrich:2001:LRT

[537] Dolph Ulrich. A legacy recalled and a tradition continued. Journal of Automated Reasoning, 27(2): 97–122, August 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010683508225.

Veroff:2001:FSP

[538] Robert Veroff. Finding shortest proofs: An application of linked inference rules. Journal of Automated Reasoning, 27 (2):123–139, August 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1010635625063.

Harris:2001:DOS

[539] Kenneth Harris and Branden Fitelson. Distributivity in Łℵ_{0} and other sentential logics. Journal of Automated Reasoning, 27(2):141–156, August 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1010687609134.

Veroff:2001:SOQ

[540] Robert Veroff. Solving open questions and other challenge problems using proof sketches. Journal of Automated Reasoning, 27(2):157–174, August 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1010639725972.

Wos:2001:CMS

[541] Larry Wos. Conquering the Meredith single axiom. Journal of Automated Reasoning, 27(2):175–199, August 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1010691726881.

Fitelson:2001:MPF

[542] Branden Fitelson and Larry Wos. Missing proofs found. Journal of Automated Reasoning, 27(2):201–225, August 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1010695827789.

Sutcliffe:2001:CAS

[543] G. Sutcliffe. The CADE-17 ATP system competition. Journal of Automated Reasoning, 27(3):227–250, October 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017517027537.

Littman:2001:SBS

[544] Michael L. Littman, Stephen M. Majercik, and Toniann Pitassi. Stochastic Boolean satisfiability. Journal of Automated Reasoning, 27(3):251– 296, October 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017584715408.

Fiorino:2001:SDP

[545] Guido Fiorino. An O(n log n)-SPACE decision procedure for the propositional Dummett logic. Journal of Automated Reasoning, 27(3):297–311, October 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1017515831550.

He:2001:SIS

[546] Lifeng He. I-SATCHMO: An improvement of SATCHMO. Journal of Automated Reasoning, 27(3):313– 322, October 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1017594402123.

Gamboa:2001:NAA

[547] Ruben A. Gamboa and Matt Kaufmann. Nonstandard analysis in ACL2. Journal of Automated Reasoning, 27 (4):323–351, November 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1011908113514.

Bittencourt:2001:ADT

[548] Guilherme Bittencourt and Isabel Tonin. An algorithm for dual transformation in first-order logic. Journal of Automated Reasoning, 27(4): 353–389, November 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1011917032470.

Bonatti:2001:RSS

[549] P. A. Bonatti. Resolution for skeptical stable model semantics. Journal of Automated Reasoning, 27(4):391– 421, November 2001. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1011960831261.

Anonymous:2001:AIVb

[550] Anonymous. Author index to volume 27. Journal of Automated Reasoning, 27(4):423, November 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1017317122485.

Anonymous:2001:SIVb

[551] Anonymous. Subject index to volume 27. Journal of Automated Reasoning, 27(4):425, November 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1017369106556.

Anonymous:2001:CVb

[552] Anonymous. Contents to volume 27. Journal of Automated Reasoning, 27(4):427–428, November 2001. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1017348505647.

Yahya:2002:DGD

[553] Adnan H. Yahya. Duality for goaldriven query processing in disjunctive deductive databases. Journal of Automated Reasoning, 28(1):1–34, January 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020109502432.

Eker:2002:SEA

[554] Steven Eker. Single elementary associative–commutative matching. Journal of Automated Reasoning, 28 (1):35–51, January 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1020122610698.

Shen:2002:SRW

[555] Yi-Dong Shen, Li-Yan Yuan, and Jia-Huai You. SLT-Resolution for the well-founded semantics. Journal of Automated Reasoning, 28(1): 53–97, January 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1020116927466.

Gent:2002:SY

[556] Ian P. Gent and Toby Walsh. Satisfiability in the year 2000. Journal of Automated Reasoning, 28(2): 99, February 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1015051815026.

Cadoli:2002:AEQ

[557] Marco Cadoli, Marco Schaerf, Andrea Giovanardi, and Massimo Giovanardi. An algorithm to evaluate quantified Boolean formulae and its experimental evaluation. Journal of Automated Reasoning, 28(2):101– 142, February 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1015019416843.

Giunchiglia:2002:SBD

[558] Enrico Giunchiglia, Armando Tacchella, and Fausto Giunchiglia. SAT-Based decision procedures for classical modal logics. Journal of Automated Reasoning, 28(2):143–171, February 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015071400913.

Horrocks:2002:EOD

[559] Ian Horrocks and Peter F. Patel-Schneider. Evaluating optimized decision procedures for propositional modal K_{(m)} satisfiability. Journal of Automated Reasoning, 28(2):173– 204, February 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015063215934.

Hustadt:2002:URT

[560] Ullrich Hustadt and Renate A. Schmidt. Using resolution for testing modal satisfiability and building models. Journal of Automated Reasoning, 28(2):205–232, February 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015067300005.

Anonymous:2002:E

[561] Anonymous.  Erratum.  Jour-

nal of Automated Reasoning, 28(2): 233, February 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1023/A%3A1017482326325.pdf.

Cremanns:2002:CPF

[562] Robert Cremanns and Friedrich Otto. A completion procedure for finitely presented groups that is based on word cycles. Journal of Automated Reasoning, 28(3):235–256, April 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1023/A%3A1015741511536.

Janicic:2002:GSF

[563] Predrag Janičić and Alan Bundy. A general setting for flexibly combining and augmenting decision procedures. Journal of Automated Reasoning, 28 (3):257–305, April 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1015707001763.

Sutcliffe:2002:IAS

[564] G. Sutcliffe, C. B. Suttner, and F. J. Pelletier. The IJCAR ATP system competition. Journal of Automated Reasoning, 28(3):307–320, April 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1015736313131.

Barendregt:2002:ACF

[565] Henk Barendregt and Erik Barendsen. Autarkic computations in formal proofs. Journal of Automated Reasoning, 28(3):321–336, April 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1015761529444.

Knapik:2002:CSP

[566] Teodor Knapik. Checking simple properties of transition systems defined by Thue specifications. Journal of Automated Reasoning, 28(4):

337–369, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015850228563.

Hooker:2002:PIM

[567] J. N. Hooker, G. Rago, V. Chandru, and A. Shrivastava. Partial instantiation methods for inference in first-order logic. Journal of Automated Reasoning, 28(4):371–396, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015854101244.

Montanari:2002:ATT

[568] Angelo Montanari, Alberto Policriti, and Matteo Slanina. Alternative translation techniques for propositional and first-order modal logics. Journal of Automated Reasoning, 28(4): 397–415, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015849504706.

Goldberg:2002:PUC

[569] Eugene Goldberg. Proving unsatisfiability of CNFs locally. Journal of Automated Reasoning, 28(4): 417–434, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1015803927637.

Anonymous:2002:AIVa

[570] Anonymous. Author index to volume 28. Journal of Automated Reasoning, 28(4):435, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1015830623811.

Anonymous:2002:SIVa

[571] Anonymous. Subject index to volume 28. Journal of Automated Reasoning, 28(4):437, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1015850010607.

Anonymous:2002:CVa

[572] Anonymous. Contents to volume 28. Journal of Automated Reasoning, 28 (4):439–440, May 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/A%3A1015889609698.

McCune:2002:SSA

[573] William McCune, Robert Veroff, Branden Fitelson, Kenneth Harris, Andrew Feist, and Larry Wos. Short single axioms for Boolean algebra. Journal of Automated Reasoning, 29(1): 1–16, January 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1020542009983.

Yahya:2002:OSH

[574] Adnan Yahya and David A. Plaisted. Ordered semantic hyper tableaux. Journal of Automated Reasoning, 29 (1):17–57, January 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1020512924736.

Slaney:2002:MPA

[575] John Slaney. More proofs of an axiom of Łukasiewicz. Journal of Automated Reasoning, 29(1):59–66, January 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020541318160.

Thiele:2002:HTF

[576] Ruediger Thiele and Larry Wos. Hilbert’s twenty-fourth problem. Journal of Automated Reasoning, 29(1): 67–89, January 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020537107897.

Borkar:2002:MPE

[577] Vivek S. Borkar, Vijay Chandru, and Sanjoy K. Mitter. Mathematical programming embeddings of logic. Journal of Automated Reasoning, 29(1): 91–106, January 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1020521718123.

Wos:2002:VXQ

[578] Larry Wos, Dolph Ulrich, and Branden Fitelson. Vanquishing the XCB question: The methodological discovery of the last shortest single axiom for the equivalential calculus. Journal of Automated Reasoning, 29(2): 107–124, February 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1021693818601.

Amgoud:2002:IIP

[579] Leila Amgoud and Claudette Cayrol. Inferring from inconsistency in preference-based argumentation frameworks. Journal of Automated Reasoning, 29(2):125–169, February 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021603608656.

Mashevitzky:2002:UTC

[580] G. Mashevitzky. Unification types of completely regular semigroups. Journal of Automated Reasoning, 29(2): 171–182, February 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021695118634.

Kamareddine:2002:PMA

[581] Fairouz Kamareddine. Preface: Mechanizing and automating mathematics: In honour of N. G. de Bruijn. Journal of Automated Reasoning, 29(3–4): 183–188, September 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021977414812.

Bancerek:2002:CCL

[582] Grzegorz Bancerek and Piotr Rudnicki. A compendium of continuous lattices in MIZAR. Journal of Automated Reasoning, 29(3–4):189– 224, September 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021966832558.

Bertot:2002:PGS

[583] Yves Bertot, Nicolas Magaud, and Paul Zimmermann. A proof of GMP square root. Journal of Automated Reasoning, 29(3–4):225–252, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1023/A:1021987403425.

Bezem:2002:APC

[584] Marc Bezem, Dimitri Hendriks, and Hans de Nivelle. Automated proof construction in type theory using resolution. Journal of Automated Reasoning, 29(3–4):253–275, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1021939521172.

Hendriks:2002:PRC

[585] Dimitri Hendriks. Proof reflection in Coq. Journal of Automated Reasoning, 29(3–4):277–307, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1021923116629.

Nguyen:2002:ERS

[586] Quang Huy Nguyen, Claude Kirchner, and Hélène Kirchner. External rewriting for skeptical proof assistants. Journal of Automated Reasoning, 29(3–4): 309–336, September 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1021975117537.

Prevosto:2002:API

[587] Virgile Prevosto and Damien Doligez. Algorithms and proofs inheritance in the FOC language. Journal of Automated Reasoning, 29(3–4):337– 363, September 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021979218446.

Wiedijk:2002:NIA

[588] Freek Wiedijk. A new implementation of Automath. Journal of Automated Reasoning, 29(3–4):365– 387, September 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021983302516.

Wenzel:2002:CMI

[589] Markus Wenzel and Freek Wiedijk. A comparison of Mizar and Isar. Journal of Automated Reasoning, 29(3–4): 389–411, September 2002. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/A%
                                        3A1021935419355.

Anonymous:2002:AIVb

[590] Anonymous. Author index to volume 29. Journal of Automated Reasoning, 29(3–4):413, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1023/
                                        A%3A1021953115802.

Anonymous:2002:SIVb

[591] Anonymous. Subject index to volume 29. Journal of Automated Reasoning, 29(3–4):415, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1021910531731.

Anonymous:2002:CVb

[592] Anonymous. Contents to volume 29. Journal of Automated Reasoning, 29(3–4):417–418, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
A%3A1021905232640.

Tinelli:2003:CBR

[593] Cesare Tinelli. Cooperation of background reasoners in theory reasoning by residue sharing. Journal of Automated Reasoning, 30(1):1–31, January 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022587501759.

Santone:2003:MLS

[594] Antonella Santone and Gigliola Vaglini. Modifying LOTOS specifications by means of automatable formula-based integrations. Journal of Automated Reasoning, 30(1):33–58, January 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022561505762.

Kamareddine:2003:FSN

[595] Fairouz Kamareddine and Haiyan Qiao. Formalizing strong normalization proofs of explicit substitution calculi in ALF. Journal of Automated Reasoning, 30(1):59–98, January 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1022527914293.

Bofill:2003:PKB

[596] Miquel Bofill, Guillem Godoy, Robert Nieuwenhuis, and Albert Rubio. Paramodulation and Knuth–Bendix completion with nontotal and nonmonotonic orderings. Journal of Automated Reasoning, 30(1):99–120, January 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1022515030222.

Voronkov:2003:PSI

[597] Andrei Voronkov. Proof-search in intuitionistic logic with equality, or back to simultaneous rigid E-unification. Journal of Automated Reasoning, 30(2): 121–151, February 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023260415982.

Rusinowitch:2003:MVI

[598] Michaël Rusinowitch, Sorin Stratulat, and Francis Klay. Mechanical verification of an ideal incremental ABR conformance algorithm. Journal of Automated Reasoning, 30(2): 153–177, February 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023251327012.

Wos:2003:SC

[599] Larry Wos. The strategy of cramming. Journal of Automated Reasoning, 30 (2):179–204, February 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1023278717096.

Chatterjee:2003:BTT

[600] Krishnendu Chatterjee, Pallab Dasgupta, and P. P. Chakrabarti. A branching time temporal framework for quantitative reasoning. Journal of Automated Reasoning, 30(2):205– 232, February 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1023217515688.

Nipkow:2003:JBV

[601] Tobias Nipkow. Java bytecode verification. Journal of Automated Reasoning, 30(3–4):233, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1025086804452.

Leroy:2003:JBV

[602] Xavier Leroy. Java bytecode verification: Algorithms and formalizations. Journal of Automated Reasoning, 30 (3–4):235–269, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025055424017.

Freund:2003:TSJ

[603] Stephen N. Freund and John C. Mitchell. A type system for the Java bytecode language and verifier. Journal of Automated Reasoning, 30(3–4): 271–321, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025011624925.

Stark:2003:CBV

[604] Robert F. Stärk and Joachim Schmid. Completeness of a bytecode verifier and a certifying Java-to-JVM compiler. Journal of Automated Reasoning, 30(3– 4):323–361, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025003423108.

Klein:2003:VBS

[605] G. Klein and M. Wildmoser. Verified bytecode subroutines. Journal of Automated Reasoning, 30(3–4): 363–398, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025095122199.

Basin:2003:BVM

[606] David A. Basin, Stefan Friedrich, and Marek Gawkowski. Bytecode verification by model checking. Journal of Automated Reasoning, 30(3–4): 399–444, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1025059508087.

Anonymous:2003:AIVa

[607] Anonymous. Author index to volume 30. Journal of Automated Reasoning, 30(3–4):445, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1025000828525.

Anonymous:2003:SIVa

[608] Anonymous. Subject index to volume 30. Journal of Automated Reasoning, 30(3–4):447, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1025010812596.

Anonymous:2003:CVa

[609] Anonymous. Contents to volume 30. Journal of Automated Reasoning, 30 (3–4):449–450, May 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/A%3A1025062813504.

Veroff:2003:SBB

[610] Robert Veroff. A shortest 2-basis for Boolean algebra in terms of the Sheffer stroke. Journal of Automated Reasoning, 31(1):1–9, September 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1027322305654.

Hai:2003:TPB

[611] Lin Hai, Sun Jigui, and Zhang Yimin. Theorem proving based on the extension rule. Journal of Automated Reasoning, 31(1):11– 21, September 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1027339205632.

Sutcliffe:2003:CAS

[612] G. Sutcliffe and C. B. Suttner. The CADE-18 ATP system competition. Journal of Automated Reasoning, 31 (1):23–32, September 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/A%3A1027302123309.

Dowek:2003:TPM

[613] Gilles Dowek, Thérèse Hardin, and Claude Kirchner. Theorem proving modulo. Journal of Automated Reasoning, 31(1):33–72, September 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1027357912519.

Gentilini:2003:BSC

[614] R. Gentilini, C. Piazza, and A. Policriti. From bisimulation to simulation: Coarsest partition problems. Journal of Automated Reasoning, 31(1):73– 103, September 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/A%
3A1027328830731.

Kapur:2003:A

[615] Deepak Kapur. Announcement. Journal of Automated Reasoning, 31(2): 105, October 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/B%3AJARS.0000009622.67092.
8d.

Manolios:2003:PFA

[616] Panagiotis Manolios and J. Strother Moore. Partial functions in ACL2. Journal of Automated Reasoning, 31 (2):107–127, October 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3AJARS.0000009505.07087.
34.

Bachmair:2003:ACC

[617] Leo Bachmair, Ashish Tiwari, and Laurent Vigneron. Abstract congruence closure. Journal of Automated Reasoning, 31(2):129–168, October 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000009518.26415.49.

Andrews:2003:HAA

[618] Peter B. Andrews. Herbrand award acceptance speech. Journal of Automated Reasoning, 31(2):169–187, October 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000009552.54063.f3.

Felty:2003:P

[619] Amy P. Felty. Preface. Journal of Automated Reasoning, 31(3–4):189– 190, November 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1023/B%3AJARS.0000021173.25739.
c5.

Hamid:2003:SAF

[620] Nadeem A. Hamid, Zhong Shao, Valery Trifonov, Stefan Monnier, and Zhaozhong Ni. A syntactic approach to foundational proof-carrying code. Journal of Automated Reasoning, 31(3– 4):191–229, November 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1023/B%3AJARS.0000021012.97318.
e9.

Appel:2003:TPC

[621] Andrew W. Appel, Neophytos Michael, Aaron Stump, and Roberto Virga. A trustworthy proof checker. Journal of Automated Reasoning, 31(3–4):231– 260, November 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000021013.61329.58.

Aspinall:2003:HBA

[622] David Aspinall and Adriana Compagnoni. Heap-bounded assembly language. Journal of Automated Reasoning, 31(3–4):261–302, November 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000021014.79255.33.

Rose:2003:LBV

[623] Eva Rose. Lightweight bytecode verification. Journal of Automated Reasoning, 31(3–4):303–334, November 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000021015.15794.82.

Riely:2003:TPT

[624] James Riely and Matthew Hennessy. Trust and partial typing in open systems of mobile agents. Journal of Automated Reasoning, 31(3–4):335– 370, November 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000021016.61054.3b.

Anonymous:2003:AIVb

[625] Anonymous. Author index to volume 31. Journal of Automated Reasoning, 31(3–4):371, November 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
B%3AJARS.0000021187.23857.1b.

Anonymous:2003:SIVb

[626] Anonymous. Subject index to volume 31. Journal of Automated Reasoning, 31(3–4):373, November 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
B%3AJARS.0000021186.93911.99.

Anonymous:2003:CVb

[627] Anonymous. Contents to volume 31. Journal of Automated Reasoning, 31(3–4):375–376, November 2003. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
B%3AJARS.0000021188.52328.42.

Kapur:2004:Pa

[628] Deepak Kapur. Preface. Journal of Automated Reasoning, 32(1):1–2, January 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1023/
B%3AJARS.0000022020.08299.5b.

Matsuda:2004:GGT

[629] Noboru Matsuda and Kurt Van-Lehn. GRAMY: A geometry theorem prover capable of construction. Journal of Automated Reasoning, 32 (1):3–33, January 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000021960.39761.b7.

Walther:2004:VC

[630] Christoph Walther and Stephan Schweitzer. Verification in the classroom. Journal of Automated Reasoning, 32(1):35–73, January 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1023/B%3AJARS.0000021872.
64036.41.

Andrews:2004:ESH

[631] Peter B. Andrews, Chad E. Brown, Frank Pfenning, Matthew Bishop, Sunil Issar, and Hongwei Xi. ETPS: A system to help students write formal proofs. Journal of Automated Reasoning, 32(1):75–92, January 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000021871.18776.94.

Hai:2004:KCU

[632] Lin Hai and Sun Jigui. Knowledge compilation using the extension rule. Journal of Automated Reasoning, 32(2): 93–102, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000029959.45572.44.

Ganzinger:2004:FTI

[633] Harald Ganzinger, Robert Nieuwenhuis, and Pilar Nivela. Fast term indexing with coded context trees. Journal of Automated Reasoning, 32(2): 103–120, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000029963.64213.ac.

Kreitz:2004:MCM

[634] Christoph Kreitz and Heiko Mantel. A matrix characterization for multiplicative exponential linear logic. Journal of Automated Reasoning, 32(2): 121–166, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000029976.22387.ac.

Johnson:2004:TQP

[635] C. A. Johnson. Top-down query processing in first-order deductive databases under the DWFS. Journal of Automated Reasoning, 32(2): 167–184, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000030010.90643.41.

Kapur:2004:Pb

[636] Deepak Kapur. Preface. Journal of Automated Reasoning, 32(3):185– 186, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1023/B%3AJARS.0000044948.61098.
                                        6c.

Makatchev:2004:ATP

[637] Maxim Makatchev, Pamela W. Jordan, and Kurt VanLehn. Abductive theorem proving for analyzing student explanations to guide feedback in intelligent tutoring systems. Journal of Automated Reasoning, 32(3):187– 226, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/B%
                                        3AJARS.0000044823.50442.cd.

Sommer:2004:PET

[638] Richard Sommer and Gregory Nuckols. A proof environment for teaching mathematics. Journal of Automated Reasoning, 32(3):227–258, February 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/article/10.1023/B%
                                        3AJARS.0000044825.55318.95.

Baumgartner:2004:LBD

[639] P. Baumgartner, U. Furbach, M. Gross-Hardt, and A. Sinner. Living book– deduction, slicing, and interaction. Journal of Automated Reasoning, 32(3): 259–286, February 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1023/B%
3AJARS.0000044872.51237.c9.

Kapur:2004:O

[640] Deepak Kapur. Obituary. Journal of Automated Reasoning, 32(4): ??, December 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF03177741.

Martin-Mateos:2004:FVG

[641] Francisco-Jesús Martín-Mateos, José-Antonio Alonso, María-José Hidalgo, and José-Luis Ruiz-Reina. Formal verification of a generic framework to synthesize SAT-provers. Journal of Automated Reasoning, 32(4):??, December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03177742.

Urbain:2004:MIA

[642] Xavier Urbain. Modular and incremental automated termination proofs. Journal of Automated Reasoning, 32 (4):??, December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03177743.

Anantharaman:2004:UMA

[643] Siva Anantharaman, Paliath Narendran, and Michael Rusinowitch. Unification modulo ACUI plus distributivity axioms. Journal of Automated Reasoning, 33(1):1–28, July 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-004-2279-7.

Paulson:2004:ONT

[644] Lawrence C. Paulson. Organizing numerical theories using axiomatic type classes. Journal of Automated Reasoning, 33(1):29–49, July 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-004-3997-6.

Zimmer:2004:CSP

[645] Jürgen Zimmer and Erica Melis. Constraint solving for proof planning. Journal of Automated Reasoning, 33 (1):51–88, July 2004. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-004-2725-6.

Liang:2004:CRR

[646] Chuck Liang, Gopalan Nadathur, and Xiaochu Qi. Choices in representation and reduction strategies for Lambda terms in intensional contexts. Journal of Automated Reasoning, 33(2): 89–132, September 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-004-6885-1.

Peltier:2004:RBM

[647] N. Peltier. Representing and building models for decidable subclasses of equational clausal logic. Journal of Automated Reasoning, 33(2): 133–170, September 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-004-5555-7.

Newborn:2004:OCL

[648] Monty Newborn and Zongyan Wang. Octopus: Combining learning and parallel search. Journal of Automated Reasoning, 33(2):171–218, September 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-004-
3243-2.

Kapur:2004:Pc

[649] Deepak Kapur and Laurent Vigneron. Preface. Journal of Automated Reasoning, 33(3–4):219–220, October/ December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-004-6240-6.

Ghilardi:2004:MTM

[650] Silvio Ghilardi. Model-theoretic methods in combined constraint satisfiability. Journal of Automated Reasoning, 33(3–4):221–249, October/December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-004-
6241-5.

Zarba:2004:DPS

[651] Calogero G. Zarba, Domenico Cantone, and Jacob T. Schwartz. A decision procedure for a sublanguage of set theory involving monotone, additive, and multiplicative functions, i: The two-level case. Journal of Automated Reasoning, 33(3–4):251–269, October/December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-004-6243-3.

delaTour:2004:CDM

[652] Thierry Boy de la Tour and Mnacho Echenim. On the complexity of deduction modulo leaf permutative equations. Journal of Automated Reasoning, 33(3–4):271–317, October/December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-004-
                                        6244-2.

Urban:2004:MMI

[653] Josef Urban. MPTP — motivation, implementation, first experiments. Journal of Automated Reasoning, 33(3– 4):319–339, October/December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10817-004-6245-
                                        1.

Feuillade:2004:RAT

[654] Guillaume Feuillade, Thomas Genet, and Valérie Viet Triem Tong. Reachability analysis over term rewriting systems. Journal of Automated Reasoning, 33(3–4):341–383, October/December 2004. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-004-
                                        6246-0.

Zarba:2005:CSC

[655] Calogero G. Zarba. Combining sets with cardinals. Journal of Automated Reasoning, 34(1):1–29, January 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
3075-8.

Remshagen:2005:EAF

[656] Anja Remshagen and Klaus Truemper. An effective algorithm for the futile questioning problem. Journal of Automated Reasoning, 34(1):31–47, January 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
0981-8.

Cantone:2005:TBD

[657] Domenico Cantone, Calogero G. Zarba, and Rosa Ruggeri Cannata. A tableaubased decision procedure for a fragment of set theory with iterated membership. Journal of Automated Reasoning, 34(1):49–72, January 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-004-8271-4.

Jurkowiak:2005:PSB

[658] Bernard Jurkowiak, Chu Min Li, and Gil Utard. A parallelization scheme based on work stealing for a class of SAT solvers. Journal of Automated Reasoning, 34(1):73–101, January 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
1970-7.

Giesl:2005:Pa

[659] Jürgen Giesl and Deepak Kapur. Preface. Journal of Automated Reasoning, 34(2):103–104, December 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10817-006-7237-0.

Zantema:2005:TSR

[660] H. Zantema. Termination of string rewriting proved automatically. Journal of Automated Reasoning, 34(2): 105–139, December 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-005-6545-0.

Serebrenik:2005:TFP

[661] Alexander Serebrenik and Danny De Schreye. Termination of floating-point computations. Journal of Automated Reasoning, 34(2):141–177, December 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-005-
                                        6546-z.

Pientka:2005:VTR

[662] Brigitte Pientka. Verifying termination and reduction properties about higher-order logic programs. Journal of Automated Reasoning, 34(2): 179–207, December 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-005-6534-3.

Tinelli:2005:CNI

[663] Cesare Tinelli and Calogero G. Zarba. Combining nonstably infinite theories. Journal of Automated Reasoning, 34 (3):209–238, April 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-5204-9.

Arkoudas:2005:SPF

[664] Konstantine Arkoudas. Simplifying proofs in Fitch-style natural deduction systems. Journal of Automated Reasoning, 34(3):239–294, April 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9000-3.

Fernandez-Gago:2005:FOT

[665] M. C. Fernández-Gago, U. Hustadt, C. Dixon, M. Fisher, and B. Konev. First-order temporal verification in practice. Journal of Automated Reasoning, 34(3):295–321, April 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-005-7354-
1.

Giesl:2005:Pb

[666] Jügen Giesl and Deepak Kapur. Preface. Journal of Automated Reasoning, 34(4):323, May 2005. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-005-9021-y.

Geser:2005:TPS

[667] Alfons Geser, Dieter Hofbauer, and Johannes Waldmann. Termination proofs for string rewriting systems via inverse match-bounds. Journal of Automated Reasoning, 34(4):365–385, May 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9024-8.

Manolios:2005:OAA

[668] Panagiotis Manolios and Daron Vroon. Ordinal arithmetic: Algorithms and mechanization. Journal of Automated Reasoning, 34(4):387–423, May 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9023-9.

Contejean:2005:MPT

[669] Evelyne Contejean, Claude Marché, Ana Paula Tomás, and Xavier Urbain. Mechanically proving termination using polynomial interpretations. Journal of Automated Reasoning, 34(4):??, May 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9022-x.

Giunchiglia:2005:SY

[670] Enrico Giunchiglia and Toby Walsh. Satisfiability in the year 2005. Journal of Automated Reasoning, 35(1– 3):1–2, October 2005. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-006-9041-2.

Bhalla:2005:HBB

[671] Ateet Bhalla, Inês Lynce, José T. de Sousa, and João Marques-Silva. Heuristic-based backtracking relaxation for propositional satisfiability. Journal of Automated Reasoning, 35 (1–3):3–24, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-9005-y.

Pan:2005:STS

[672] Guoqiang Pan and Moshe Y. Vardi. Symbolic techniques in satisfiability solving. Journal of Automated Reasoning, 35(1–3):25–50, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9009-7.

Alekhnovich:2005:ELB

[673] Michael Alekhnovich, Edward A. Hirsch, and Dmitry Itsykson. Exponential lower bounds for the running time of DPLL algorithms on satisfiable formulas. Journal of Automated Reasoning, 35(1–3):51–72, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9006-x.

Szeider:2005:BSD

[674] Stefan Szeider. Backdoor sets for DLL subsolvers. Journal of Automated Reasoning, 35(1–3):73–88, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9007-9.

Johannsen:2005:CPL

[675] Jan Johannsen. The complexity of pure literal elimination. Journal of Automated Reasoning, 35(1–3):89–95, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-005-
                                        9008-8.

Thornton:2005:CWL

[676] John Thornton. Clause weighting local search for SAT. Journal of Automated Reasoning, 35(1–3):97–142, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-005-
                                        9010-1.

Frisch:2005:SNB

[677] Alan M. Frisch, Timothy J. Peugniez, Anthony J. Doggett, and Peter W. Nightingale. Solving non–Boolean satisfiability problems with stochastic local search: A comparison of encodings. Journal of Automated Reasoning, 35 (1–3):143–179, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-005-9011-0.

Boufkhad:2005:RRS

[678] Yacine Boufkhad, Olivier Dubois, Yannet Interian, and Bart Selman. Regular random k-SAT: Properties of balanced formulas. Journal of Automated Reasoning, 35(1–3):181–200, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-005-
                                        9012-z.

Meier:2005:ASS

[679] Andreas Meier and Volker Sorge. Applying SAT solving in classification of finite algebras. Journal of Automated Reasoning, 35(1–3):201–235, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9003-0.

Armando:2005:SBA

[680] Alessandro Armando, Claudio Castellini, Enrico Giunchiglia, and Marco Maratea. The SAT-based approach to separation logic. Journal of Automated Reasoning, 35(1–3):237–263, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9002-1.

Bozzano:2005:MAS

[681] Marco Bozzano, Roberto Bruttomesso, Alessandro Cimatti, Tommi Junttila, Peter van Rossum, Stephan Schulz, and Roberto Sebastiani. M ath SAT: Tight integration of SAT and mathematical decision procedures. Journal of Automated Reasoning, 35(1– 3):265–293, October 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-9004-z.

Barthe:2005:TAS

[682] Gilles Barthe, Pierre Courtieu, Guillaume Dufay, and Simão Melo de Sousa. Tool-assisted specification and verification of typed low– level languages. Journal of Automated Reasoning, 35(4):295–354, November 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10817-005-
                                        0084-6.

Boyer:2005:TAD

[683] Robert S. Boyer, Wilfred J. Legato, and Victor W. Marek. Toward automating the discovery of decreasing measures. Journal of Automated Reasoning, 35(4):355–371, November 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-005-
                                        9020-z.

Barrett:2005:DRF

[684] Clark Barrett, Leonardo de Moura, and Aaron Stump. Design and results of the first satisfiability modulo theories competition (SMT-COMP 2005). Journal of Automated Reasoning, 35 (4):373–390, November 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9026-1.

Peltier:2005:STP

[685] N. Peltier. Some techniques for proving termination of the hyperresolution calculus. Journal of Automated Reasoning, 35(4):391–427, November 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-006-
                                        9028-z.

Shah:2005:BRR

[686] Nimish Shah. Book review: Rippling: Meta-Level Guidance for Mathematical Reasoning, by Alan Bundy, David Basin, Dieter Hutter, and Andrew Ireland, Cambridge University Press, 2005. Journal of Automated Reasoning, 35(4):429–431, November 2005. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9027-0.

Armando:2006:ARS

[687] Alessandro Armando, David Basin, and Jorge Cuellar. Automated reasoning for security protocol analysis. Journal of Automated Reasoning, 36 (1–2):1–3, January 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-9014-x.

Bella:2006:VSP

[688] Giampaolo Bella, Fabio Massacci, and Lawrence C. Paulson. Verifying the SET purchase protocols. Journal of Automated Reasoning, 36(1–2):5– 37, January 2006. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-9018-6.

Chadha:2006:FAM

[689] Rohit Chadha, Steve Kremer, and Andre Scedrov. Formal analysis of multiparty contract signing. Journal of Automated Reasoning, 36(1– 2):39–83, January 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-9019-5.

Delaune:2006:DPS

[690] Stéphanie Delaune and Florent Jacquemard. Decision procedures for the security of protocols with probabilistic encryption against offline dictionary attacks. Journal of Automated Reasoning, 36(1–2):85–124, January 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9017-7.

Huttel:2006:DIE

[691] Hans Hüttel and Jiří Srba. Decidability issues for extended ping– Pong protocols. Journal of Automated Reasoning, 36(1–2):125–147, January 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-005-
9015-9.

Steel:2006:AGP

[692] Graham Steel and Alan Bundy. Attacking group protocols by refuting incorrect inductive conjectures. Journal of Automated Reasoning, 36(1– 2):149–176, January 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-005-9016-8.

Audemard:2006:PDS

[693] Gilles Audemard, Belaïd Benhamou, and Laurent Henocque. Predicting and detecting symmetries in FOL finite model search. Journal of Automated Reasoning, 36(3):177–212, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9040-3.

Kuncak:2006:DBA

[694] Viktor Kuncak, Huu Hai Nguyen, and Martin Rinard. Deciding Boolean algebra with Presburger arithmetic. Journal of Automated Reasoning, 36 (3):213–239, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9042-1.

Fiala:2006:SSS

[695] Nick C. Fiala and Keith M. Agre. Searching for shortest single axioms for groups of exponent 6. Journal of Automated Reasoning, 36(3):241–257, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9044-z.

Simpson:2006:EGZ

[696] Carlos Simpson. Explaining Gabriel– Zisman localization to the computer. Journal of Automated Reasoning, 36 (3):259–285, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9038-x.

Fischer:2006:ESAa

[697] Bernd Fischer, Geoff Sutcliffe, and Stephan Schulz. Empirically successful automated reasoning: Systems issue. Journal of Automated Reasoning, 36(4):287, April 2006. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-006-9059-5.

Lochner:2006:TKW

[698] Bernd Löchner. Things to know when implementing KBO. Journal of Automated Reasoning, 36(4):289–310, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-006-
                                        9031-4.

Beeson:2006:MIO

[699] Michael Beeson. Mathematical induction in otter–Lambda. Journal of Automated Reasoning, 36(4):311–344, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-006-
                                        9036-z.

Ireland:2006:IAH

[700] Andrew Ireland, Bill J. Ellis, Andrew Cook, Roderick Chapman, and Janet Barnes. An integrated approach to high integrity software verification. Journal of Automated Reasoning, 36 (4):379–410, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9034-1.

Giunchiglia:2006:ASP

[701] Enrico Giunchiglia, Yuliya Lierler, and Marco Maratea. Answer set programming based on propositional satisfiability. Journal of Automated Reasoning, 36(4):??, April 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9033-2.

Fischer:2006:ESAb

[702] Bernd Fischer, Geoff Sutcliffe, and Stephan Schulz. Empirically successful automated reasoning: Applications issue. Journal of Automated Reasoning, 37(1–2):1, August 2006. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-006-9056-8.

Veroff:2006:ASB

[703] R. Veroff and M. Spinks. Axiomatizing the skew Boolean propositional calculus. Journal of Automated Reasoning, 37(1–2):3–20, August 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9029-y.

Urban:2006:MDI

[704] Josef Urban. MPTP 0.2: Design, implementation, and initial experiments. Journal of Automated Reasoning, 37 (1–2):21–43, August 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9032-3.

Sinz:2006:CCC

[705] Carsten Sinz, Wolfgang Küchlin, and Dieter Feichtinger. Checking consistency and completeness of on-line product manuals. Journal of Automated Reasoning, 37(1–2):45–66, August 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9037-y.

Ruiz-Reina:2006:FCQ

[706] José-Luis Ruiz-Reina, Francisco-Jesús Martín-Mateos, José-Antonio Alonso, and María-José Hidalgo. Formal correctness of a quadratic unification algorithm. Journal of Automated Reasoning, 37(1–2):67–92, August 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9030-5.

Manolios:2006:FVB

[707] Panagiotis Manolios and Sudarshan K. Srinivasan. A framework for verifying bit-level pipelined machines based on automated deduction and decision procedures. Journal of Automated Reasoning, 37(1–2):93–116, August 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9035-0.

Pell:2006:VFL

[708] Oliver Pell. Verification of FPGA layout generators in higher-order logic. Journal of Automated Reasoning, 37 (1–2):117–152, August 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9039-9.

Giesl:2006:TSI

[709] Jürgen Giesl and Deepak Kapur. Third special issue on techniques for automated termination proofs. Journal of Automated Reasoning, 37(3): 153–154, October 2006. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-006-9052-z.

Giesl:2006:MID

[710] Jürgen Giesl, René Thiemann, Peter Schneider-Kamp, and Stephan Falke. Mechanizing and improving dependency pairs. Journal of Automated Reasoning, 37(3):155–203, October 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9057-7.

Keiichirou:2006:ETA

[711] Kusakari Keiichirou, Nakamura Masaki, and Toyama Yoshihito. Elimination transformations for associative– commutative rewriting systems. Journal of Automated Reasoning, 37(3): 205–229, October 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9053-y.

Bailleux:2006:SCA

[712] Olivier Bailleux and Pierre Marquis. Some computational aspects of distance-sat. Journal of Automated Reasoning, 37(4):231–260, November 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9063-9.

Dequen:2006:EAS

[713] Gilles Dequen and Olivier Dubois. An efficient approach to solving random k-SAT problems. Journal of Automated Reasoning, 37(4):261–276, November 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-006-
                                        9025-2.

Givant:2006:CRF

[714] Steven Givant. The calculus of relations as a foundation for mathematics. Journal of Automated Reasoning, 37 (4):277–322, November 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9062-x.

Limet:2006:TTL

[715] Sébastien Limet and Gernot Salzer. Tree tuple languages from the logic programming point of view. Journal of Automated Reasoning, 37(4): 323–349, November 2006. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9064-8.

Beckert:2007:P

[716] Bernhard Beckert and Lawrence C. Paulson. Preface. Journal of Automated Reasoning, 38(1–3):1–2, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10817-006-9058-6.

Antonsen:2007:LVS

[717] Roger Antonsen and Arild Waaler. Liberalized variable splitting. Journal of Automated Reasoning, 38(1–3):3–30, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s10817-006-
9055-9.

Cantone:2007:SFR

[718] Domenico Cantone and Marianna Nicolosi-Asmundo. A sound framework for δ-rule variants in free-variable semantic tableaux. Journal of Automated Reasoning, 38(1–3):31–56, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9045-y.

Jacobs:2007:CIG

[719] Swen Jacobs and Uwe Waldmann. Comparing instance generation methods for automated reasoning. Journal of Automated Reasoning, 38(1– 3):57–78, April 2007. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9046-x.

Letz:2007:DTC

[720] Reinhold Letz and Gernot Stenz. The disconnection tableau calculus. Journal of Automated Reasoning, 38(1– 3):79–126, April 2007. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9048-8.

Giese:2007:SBE

[721] Martin Giese. Superposition-based equality handling for analytic tableaux. Journal of Automated Reasoning, 38 (1–3):127–153, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9050-1.

Murray:2007:EQP

[722] Neil V. Murray and Erik Rosenthal. Efficient query processing with reduced implicate tries. Journal of Automated Reasoning, 38(1–3):155–172, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-006-
                                        9054-x.

Bresolin:2007:ODP

[723] Davide Bresolin, Angelo Montanari, and Guido Sciavicco. An optimal decision procedure for right propositional neighborhood logic. Journal of Automated Reasoning, 38(1–3):173–199, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-006-
                                        9051-0.

Larchey-Wendling:2007:GBD

[724] Dominique Larchey-Wendling. Graphbased decision for Gödel–Dummett logics. Journal of Automated Reasoning, 38(1–3):201–225, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9047-9.

Lutz:2007:TAD

[725] Carsten Lutz and Maja Miličić. A tableau algorithm for description logics with concrete domains and general TBoxes. Journal of Automated Reasoning, 38(1–3):227–259, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9049-7.

Raths:2007:IPL

[726] Thomas Raths, Jens Otten, and Christoph Kreitz. The ILTP problem library for intuitionistic logic. Journal of Automated Reasoning, 38 (1–3):261–271, April 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9060-z.

Gamboa:2007:TEA

[727] R. Gamboa and J. Cowles. Theory extension in ACL2(r). Journal of Automated Reasoning, 38(4):273–301, May 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-006-
9043-0.

Aiguier:2007:SAR

[728] Marc Aiguier and Diane Bahrami. Structures for abstract rewriting. Journal of Automated Reasoning, 38(4): 303–351, May 2007. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-006-9065-7.

Avigad:2007:DPL

[729] Jeremy Avigad and Kevin Donnelly. A decision procedure for linear “Big O” equations. Journal of Automated Reasoning, 38(4):353–373, May 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10817-007-
                                        9066-1.

Ciaffaglione:2007:RAO

[730] Alberto Ciaffaglione, Luigi Liquori, and Marino Miculan. Reasoning about object-based calculi in (co)inductive type theory and the theory of contexts. Journal of Automated Reasoning, 39(1):1–47, July 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-006-9061-y.

Buning:2007:BFM

[731] Hans Kleine Büning, K. Subramani, and Xishun Zhao. Boolean functions as models for quantified Boolean formulas. Journal of Automated Reasoning, 39(1):49–75, July 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-007-9067-0.

Morawska:2007:GUE

[732] Barbara Morawska. General Eunification with eager variable elimination and a nice cycle rule. Journal of Automated Reasoning, 39(1):77–106, July 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-007-
                                        9068-z.

Aspinall:2007:SIU

[733] David Aspinall and Christoph Lüth. Special issue on user interfaces in theorem proving: Preface. Journal of Automated Reasoning, 39(2): 107–108, August 2007. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-007-9075-0.

Asperti:2007:UIM

[734] Andrea Asperti, Claudio Sacerdoti Coen, Enrico Tassi, and Stefano Zacchiroli. User interaction with the Matita proof assistant. Journal of Automated Reasoning, 39(2):109–139, August 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9070-5.

Cairns:2007:ISA

[735] Paul Cairns and Jeremy Gow. Integrating searching and authoring in Mizar. Journal of Automated Reasoning, 39 (2):141–160, August 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-007-9073-2.

Narboux:2007:GUI

[736] Julien Narboux. A graphical user interface for formal proofs in geometry. Journal of Automated Reasoning, 39 (2):161–180, August 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-007-9071-4.

Billingsley:2007:SPE

[737] William Billingsley and Peter Robinson. Student proof exercises using MathsTiles and Isabelle/HOL in an intelligent book. Journal of Automated Reasoning, 39(2):181–218, August 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9072-3.

Sinz:2007:VSI

[738] Carsten Sinz. Visualizing SAT instances and runs of the DPLL algorithm. Journal of Automated Reasoning, 39(2):219–243, August 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-007-9074-
1.

Baader:2007:PSI

[739] Franz Baader. Preface to special issue on reasoning in description logics. Journal of Automated Reasoning, 39 (3):245–247, October 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-007-9081-2.

Horrocks:2007:TDP

[740] Ian Horrocks and Ulrike Sattler. A tableau decision procedure for SHOIQ. Journal of Automated Reasoning, 39(3):249–276, October 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9079-9.

Tsarkov:2007:OTR

[741] Dmitry Tsarkov, Ian Horrocks, and Peter F. Patel-Schneider. Optimizing terminological reasoning for expressive description logics. Journal of Automated Reasoning, 39(3):277–316, October 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9077-y.

Schlobach:2007:DIT

[742] Stefan Schlobach, Zhisheng Huang, Ronald Cornet, and Frank van Harmelen. Debugging incoherent terminologies. Journal of Automated Reasoning, 39(3):317–349, October 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-007-9076-z.pdf.

Hustadt:2007:RDL

[743] Ullrich Hustadt, Boris Motik, and Ulrike Sattler. Reasoning in description logics by a reduction to disjunctive Datalog. Journal of Automated Reasoning, 39(3):351–384, October 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9080-3.

Calvanese:2007:TRE

[744] Diego Calvanese, Giuseppe De Giacomo, Domenico Lembo, Maurizio Lenzerini, and Riccardo Rosati. Tractable reasoning and efficient query answering in description logics: The DL–Lite family. Journal of Automated Reasoning, 39(3):385–429, October 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9078-x.

Stapleton:2007:ATP

[745] Gem Stapleton, Judith Masthoff, Jean Flower, Andrew Fish, and Jane Southern. Automated theorem proving in Euler diagram systems. Journal of Automated Reasoning, 39(4):431–470, December 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-007-
                                        9069-y.

Forest:2007:ERS

[746] Julien Forest and Delia Kesner. Expression reduction systems with patterns. Journal of Automated Reasoning, 39(4):513–541, December 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10817-007-9083-
                                        0.

Brooks:2007:IPT

[747] Daniel R. Brooks, Esra Erdem, Selim T. Erdoğan, James W. Minett, and Don Ringe. Inferring phylogenetic trees using answer set programming. Journal of Automated Reasoning, 39(4):??, December 2007. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-007-9082-1.

Liffiton:2008:ACM

[748] Mark H. Liffiton and Karem A. Sakallah. Algorithms for computing minimal unsatisfiable subsets of constraints. Journal of Automated Reasoning, 40(1):1–33, January 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-007-9084-z.

Meng:2008:THO

[749] Jia Meng and Lawrence C. Paulson. Translating higher-order clauses to first-order clauses. Journal of Automated Reasoning, 40(1):35–60, January 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9085-y.

Bezem:2008:MPH

[750] Marc Bezem and Dimitri Hendriks. On the mechanization of the proof of Hessenberg’s theorem in coherent logic. Journal of Automated Reasoning, 40 (1):61–85, January 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-007-9086-x.

Furbach:2008:IP

[751] Ulrich Furbach. IJCAR preface. Journal of Automated Reasoning, 40(2– 3):87–88, March 2008. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-007-9088-8.

Kazakov:2008:RBDa

[752] Yevgeny Kazakov and Boris Motik. A resolution-based decision procedure for SHOIQ. Journal of Automated Reasoning, 40(2–3):89–116, March 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10817-007-
                                        9090-1.

Toman:2008:KFD

[753] David Toman and Grant Weddell. On keys and functional dependencies as first-class citizens in description logics. Journal of Automated Reasoning, 40 (2–3):117–132, March 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-007-9092-z.

Chaudhuri:2008:LCF

[754] Kaustuv Chaudhuri, Frank Pfenning, and Greg Price. A logical characterization of forward and backward chaining in the inverse method. Journal of Automated Reasoning, 40(2–3):133–177, March 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-007-
                                        9091-0.

Paskevich:2008:CTL

[755] Andrei Paskevich. Connection tableaux with lazy paramodulation. Journal of Automated Reasoning, 40(2– 3):179–194, March 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-007-9089-7.

Endrullis:2008:MIP

[756] Jörg Endrullis, Johannes Waldmann, and Hans Zantema. Matrix interpretations for proving termination of term rewriting. Journal of Automated Reasoning, 40(2–3):195–220, March 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9087-9.

Sorge:2008:ACV

[757] Volker Sorge, Andreas Meier, Roy Mc-Casland, and Simon Colton. Automatic construction and verification of isotopy invariants. Journal of Automated Reasoning, 40(2–3):221–243, March 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9093-y.

Ray:2008:MAP

[758] Sandip Ray, Warren A. Hunt, Jr.., John Matthews, and J. Strother Moore. A mechanical analysis of program verification strategies. Journal of Automated Reasoning, 40(4):245–269, May 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-008-
9098-1.

Aransay:2008:MPB

[759] Jesús Aransay, Clemens Ballarin, and Julio Rubio. A mechanized proof of the basic perturbation lemma. Journal of Automated Reasoning, 40(4):271–292, May 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9094-x.

Brock:2008:RER

[760] Bishop Brock, Matt Kaufmann, and J. Strother Moore. Rewriting with equivalence relations in ACL2. Journal of Automated Reasoning, 40(4): 293–306, May 2008. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-007-9095-9.

Rideau:2008:TWC

[761] Laurence Rideau, Bernard Paul Serpette, and Xavier Leroy. Tilting at windmills with Coq: Formal verification of a compilation algorithm for parallel moves. Journal of Automated Reasoning, 40(4):307–326, May 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-007-
9096-8.

Urban:2008:NTI

[762] Christian Urban. Nominal techniques in Isabelle/HOL. Journal of Automated Reasoning, 40(4):327–356, May 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-008-
9097-2.

Kazakov:2008:RBDb

[763] Yevgeny Kazakov and Boris Motik. A resolution-based decision procedure for SHOIQ. Journal of Automated Reasoning, 40(4):357, May 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-008-9100-y.pdf.

Leroy:2008:FVC

[764] Xavier Leroy and Sandrine Blazy. Formal verification of a C-like memory model and its uses for verifying program transformations. Journal of Automated Reasoning, 41(1):1–31, July 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-008-
9099-0.

Ortiz:2008:DCQ

[765] Magdalena Ortiz, Diego Calvanese, and Thomas Eiter. Data complexity of query answering in expressive description logics via tableaux. Journal of Automated Reasoning, 41(1):61–98, July 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-008-
9102-9.

Chaieb:2008:PSR

[766] Amine Chaieb and Tobias Nipkow. Proof synthesis and reflection for linear arithmetic. Journal of Automated Reasoning, 41(1):??, July 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-008-9101-x.

Haarslev:2008:SDL

[767] Volker Haarslev and Ralf Möller. On the scalability of description logic instance retrieval. Journal of Automated Reasoning, 41(2):99–142, August 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-008-
9104-7.

Platzer:2008:DDL

[768] André Platzer. Differential dynamic logic for hybrid systems. Journal of Automated Reasoning, 41(2):143–189, August 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-008-

9103-8. See correction [1266].

Autexier:2008:P

[769] Serge Autexier, Heiko Mantel, Stephan Merz, and Tobias Nipkow. Preface. Journal of Automated Reasoning, 41(3– 4):191–192, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s10817-008-9110-
                                        9.

Amjad:2008:DCP

[770] Hasan Amjad. Data compression for proof replay. Journal of Automated Reasoning, 41(3–4):193–218, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-008-
                                        9109-2.

Brucker:2008:EEO

[771] Achim D. Brucker and Burkhart Wolff. An extensible encoding of objectoriented data models in HOL. Journal of Automated Reasoning, 41(3–4): 219–249, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-008-9108-3.

Hasan:2008:UTP

[772] Osman Hasan and Sofiène Tahar. Using theorem proving to verify expectation and variance for discrete random variables. Journal of Automated Reasoning, 41(3–4):295–323, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-008-
9113-6.

Gorrieri:2008:FMA

[773] Roberto Gorrieri, Fabio Martinelli, and Marinella Petrocchi. Formal models and analysis of secure multicast in wired and wireless networks. Journal of Automated Reasoning, 41(3–4): 325–364, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-008-9112-7.

Moffat:2008:ACS

[774] Nick Moffat and Michael Goldsmith. Assumption–commitment support for CSP model checking. Journal of Automated Reasoning, 41(3–4):365– 398, November 2008. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-008-9111-8.

Donaldson:2008:ASD

[775] Alastair F. Donaldson and Alice Miller. Automatic symmetry detection for Promela. Journal of Automated Reasoning, 41(3–4):??, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10817-008-
                                        9107-4.

Hasan:2009:PAF

[776] Osman Hasan and Sofiène Tahar. Performance analysis and functional verification of the stop-and-wait protocol in HOL. Journal of Automated Reasoning, 42(1):1–33, January 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-008-9105-6.

Johnson:2009:COM

[777] C. A. Johnson. Computing only minimal answers in disjunctive deductive databases. Journal of Automated Reasoning, 42(1):35–76, January 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-008-
                                        9106-5.

Samer:2009:BSQ

[778] Marko Samer and Stefan Szeider. Backdoor sets of quantified Boolean formulas. Journal of Automated Reasoning, 42(1):77–97, January 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-008-9114-5.

Bjork:2009:FOS

[779] Magnus Björk. First order Stålmarck. Journal of Automated Reasoning, 42 (1):99–122, January 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-008-9115-4.

Klein:2009:OSV

[780] Gerwin Klein, Ralf Huuck, and Bastian Schlich. Operating system verification. Journal of Automated Reasoning, 42(2–4):123–124, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-009-9126-9.pdf.

Tuch:2009:FVC

[781] Harvey Tuch. Formal verification of C systems code. Journal of Automated Reasoning, 42(2–4):125–187, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9120-2.

Tews:2009:FMM

[782] Hendrik Tews, Marcus Völp, and Tjark Weber. Formal memory models for the verification of low-level operating-system code. Journal of Automated Reasoning, 42(2–4):189–227, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9122-0.

Gallardo:2009:MCD

[783] María del Mar Gallardo, Pedro Merino, and David Sanán. Model checking dynamic memory allocation in operating systems. Journal of Automated Reasoning, 42(2–4):229–264, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9124-y.

Tlili:2009:IFS

[784] Syrine Tlili and Mourad Debbabi. Interprocedural and flow-sensitive type analysis for memory and type safety of C code. Journal of Automated Reasoning, 42(2–4):265–300, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-009-
                                        9121-1.

Feng:2009:CLL

[785] Xinyu Feng, Zhong Shao, Yu Guo, and Yuan Dong. Certifying low-level programs with hardware interrupts and preemptive threads. Journal of Automated Reasoning, 42(2–4):301–347, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-009-
                                        9118-9.

Daum:2009:PFI

[786] Matthias Daum, Jan Dörrenbächer, and Burkhart Wolff. Proving fairness and implementation correctness of a microkernel scheduler. Journal of Automated Reasoning, 42(2–4):349–388, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-009-
                                        9119-8.

Alkassar:2009:BL

[787] Eyad Alkassar, Mark A. Hillebrand, Dirk C. Leinenbach, Norbert W. Schirmer, Artem Starostin, and Alexandra Tsyban. Balancing the load. Journal of Automated Reasoning, 42 (2–4):389–454, April 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9123-z.

Blanchette:2009:PPM

[788] Jasmin Christian Blanchette. Proof pearl: Mechanizing the textbook proof of Huffman’s algorithm. Journal of Automated Reasoning, 43(1):1–18, June 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9116-y.

Dufourd:2009:IPD

[789] Jean-François Dufourd. An intuitionistic proof of a discrete form of the Jordan curve theorem formalized in Coq with combinatorial hypermaps. Journal of Automated Reasoning, 43 (1):19–51, June 2009. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9117-x.

Monroy:2009:PEE

[790] Raúl Monroy, Alan Bundy, and Ian Green. On process equivalence = equation solving in CCS. Journal of Automated Reasoning, 43(1):53–80, June 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9125-x.

Maric:2009:FIM

[791] Filip Marić. Formalization and implementation of modern SAT solvers. Journal of Automated Reasoning, 43 (1):81–119, June 2009. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-009-9127-8.

Subramani:2009:OLR

[792] K. Subramani. Optimal length resolution refutations of difference constraint systems. Journal of Automated Reasoning, 43(2):121–137, August 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-009-
                                        9139-4.

Gamboa:2009:FPA

[793] Ruben A. Gamboa. A formalization of powerlist algebra in ACL2. Journal of Automated Reasoning, 43 (2):139–172, August 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-009-9140-y.

Zankl:2009:KO

[794] Harald Zankl, Nao Hirokawa, and Aart Middeldorp. KBO orientability. Journal of Automated Reasoning, 43 (2):173–201, August 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-009-9131-z.

Dalzotto:2009:PAD

[795] G. Dalzotto and T. Recio. On protocols for the automated discovery of theorems in elementary geometry. Journal of Automated Reasoning, 43 (2):203–236, August 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9133-x.

Boulton:2009:CAR

[796] Richard Boulton, Joe Hurd, and Konrad Slind. Computer assisted reasoning. Journal of Automated Reasoning, 43(3):237–242, October 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9144-7.

Harrison:2009:FAP

[797] John Harrison. Formalizing an analytic proof of the prime number theorem. Journal of Automated Reasoning, 43(3):243–261, October 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9145-6.

Blazy:2009:MSC

[798] Sandrine Blazy and Xavier Leroy. Mechanized semantics for the Clight subset of the C language. Journal of Automated Reasoning, 43(3): 263–288, October 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9148-3.

Nipkow:2009:SCT

[799] Tobias Nipkow. Social choice theory in HOL. Journal of Automated Reasoning, 43(3):289–304, October 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10817-009-
9147-4.

Norrish:2009:RCI

[800] Michael Norrish. Rewriting conversions implemented with continuations. Journal of Automated Reasoning, 43 (3):305–336, October 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9146-5.

Leroy:2009:FVC

[801] Xavier Leroy. A formally verified compiler back-end. Journal of Automated Reasoning, 43(4):??, December 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9155-4.

Sutcliffe:2009:TPL

[802] Geoff Sutcliffe. The TPTP problem library and associated infrastructure. Journal of Automated Reasoning, 43(4):??, December 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9143-8.

Bohme:2010:HBI

[803] Sascha Böhme, Michał Moskal, Wolfram Schulte, and Burkhart Wolff. HOL-Boogie — an interactive proverbackend for the verifying C compiler. Journal of Automated Reasoning, 44 (1–2):??, February 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9142-9.

Carette:2010:P

[804] Jacques Carette, Makarius Wenzel, and Freek Wiedijk. Preface. Journal of Automated Reasoning, 44(1–2):??, February 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
009-9141-x.pdf.

Coen:2010:DRP

[805] Claudio Sacerdoti Coen. Declarative representation of proof terms. Journal of Automated Reasoning, 44 (1–2):??, February 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9136-7.

Dietrich:2010:CIS

[806] Dominik Dietrich and Ewaryst Schulz. Crystal: Integrating structured queries into a tactic language. Journal of Automated Reasoning, 44(1–2):??, February 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9138-5.

Guidi:2010:PRC

[807] Ferruccio Guidi. Procedural representation of CIC proof terms. Journal of Automated Reasoning, 44(1– 2):??, February 2010. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9137-6.

James:2010:FMC

[808] Perry R. James and Patrice Chalin. Faster and more complete extended static checking for the Java modeling language. Journal of Automated Reasoning, 44(1–2):??, February 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9134-9.

Janicic:2010:GCL

[809] Predrag Janičić. Geometry constructions language. Journal of Automated Reasoning, 44(1–2):??, February 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9135-8.

Akbarpour:2010:MAT

[810] Behzad Akbarpour and Lawrence Charles Paulson. MetiTarski: An automatic theorem prover for real-valued special functions. Journal of Automated Reasoning, 44(3):175–205, March 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-009-9149-
2.

Mendler:2010:TCD

[811] Michael Mendler and Stephan Scheele. Towards constructive DL for abstraction and refinement. Journal of Automated Reasoning, 44(3):207–243, March 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10817-009-
9151-8.

Hermant:2010:RCF

[812] Olivier Hermant. Resolution is cutfree. Journal of Automated Reasoning, 44(3):245–276, March 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9153-6.

Pipatsrisawat:2010:MCL

[813] Knot Pipatsrisawat and Adnan Darwiche. On modern clause-learning satisfiability solvers. Journal of Automated Reasoning, 44(3):277–301, March 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9156-3.

Krauss:2010:PNR

[814] Alexander Krauss. Partial and nested recursive function definitions in higherorder logic. Journal of Automated Reasoning, 44(4):303–336, April 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-009-9157-
2.

Grau:2010:ICD

[815] Bernardo Cuenca Grau, Christian Halaschek-Wiener, Yevgeny Kazakov, and Boontawee Suntisrivaraporn. Incremental classification of description logics ontologies. Journal of Automated Reasoning, 44(4):337–369, April 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10817-009-
                                        9159-0.

Ferrari:2010:BCD

[816] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. BCDL: Basic constructive description logic. Journal of Automated Reasoning, 44(4):371–399, April 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-009-
                                        9160-7.

Piskac:2010:DEP

[817] Ruzica Piskac, Leonardo de Moura, and Nikolaj Bjørner. Deciding effectively propositional logic using DPLL and substitution sets. Journal of Automated Reasoning, 44(4):401–424, April 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-009-
                                        9161-6.

Xiang:2010:PSI

[818] Yang Xiang and Kevin Grant. Preface: Special issue on uncertain reasoning. Journal of Automated Reasoning, 45(1):1–2, June 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-009-9158-1.pdf.

Kern-Isberner:2010:PFD

[819] Gabriele Kern-Isberner and Manuela Ritterskamp. Preference fusion for default reasoning beyond System Z. Journal of Automated Reasoning, 45 (1):3–19, June 2010. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9129-6.

Palacios-Alonso:2010:ELD

[820] Miguel A. Palacios-Alonso, Carlos A. Brizuela, and L. Enrique Sucar. Evolutionary learning of dynamic naive Bayesian classifiers. Journal of Automated Reasoning, 45(1):21–37, June 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9130-0.

Benferhat:2010:NDT

[821] Salem Benferhat, Safa Yahi, and Habiba Drias. A new default theories compilation for MSP–Entailment. Journal of Automated Reasoning, 45 (1):39–59, June 2010. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9150-9.

Pichon:2010:UDR

[822] Frédéric Pichon and Thierry Denœux. The unnormalized Dempster’s rule of combination: A new justification from the least commitment principle and some extensions. Journal of Automated Reasoning, 45(1):61–87, June 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9152-7.

Armando:2010:P

[823] Alessandro Armando, Peter Baumgartner, and Gilles Dowek. Preface. Journal of Automated Reasoning, 45(2):89, August 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-010-9180-3.pdf.

Baader:2010:ABA

[824] Franz Baader and Rafael Peñaloza. Automata-based axiom pinpointing. Journal of Automated Reasoning, 45 (2):91–129, August 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9181-2.

Bozga:2010:QSL

[825] Marius Bozga, Radu Iosif, and Swann Perarnau. Quantitative separation logic and programs with lists. Journal of Automated Reasoning, 45(2): 131–156, August 2010. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9179-9.

Nigam:2010:FPS

[826] Vivek Nigam and Dale Miller. A framework for proof systems. Journal of Automated Reasoning, 45(2):157–188, August 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-010-
                                        9182-1.

Nipkow:2010:LQE

[827] Tobias Nipkow. Linear quantifier elimination. Journal of Automated Reasoning, 45(2):189–212, August 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-010-9183-
0.

Ye:2010:VDPa

[828] Zheng Ye, Shang-Ching Chou, and Xiao-Shan Gao. Visually dynamic presentation of proofs in plane geometry. Journal of Automated Reasoning, 45 (3):213–241, October 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9162-5.

Ye:2010:VDPb

[829] Zheng Ye, Shang-Ching Chou, and Xiao-Shan Gao. Visually dynamic presentation of proofs in plane geometry. Journal of Automated Reasoning, 45 (3):243–266, October 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9163-4.

Cheney:2010:EU

[830] James Cheney. Equivariant unification. Journal of Automated Reasoning, 45 (3):267–300, October 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9164-3.

Galdino:2010:FKB

[831] André L. Galdino and Mauricio Ayala-Rincón. A formalization of the Knuth– Bendix(–Huet) critical pair theorem. Journal of Automated Reasoning, 45 (3):301–325, October 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9165-2.

Beckert:2010:TP

[832] Bernhard Beckert and Reiner Hähnle. Tests and proofs. Journal of Automated Reasoning, 45(4):327–329, December 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9173-2.

Calvagna:2010:FLA

[833] Andrea Calvagna and Angelo Gargantini. A formal logic approach to constrained combinatorial testing. Journal of Automated Reasoning, 45(4): 331–358, December 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9171-4.

Dunets:2010:AFD

[834] Andriy Dunets, Gerhard Schellhorn, and Wolfgang Reif. Automated flaw detection in algebraic specifications. Journal of Automated Reasoning, 45 (4):359–395, December 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9166-1.

Angeletti:2010:UBM

[835] Damiano Angeletti, Enrico Giunchiglia, Massimo Narizzano, Alessandra Puddu, and Salvatore Sabina. Using bounded model checking for coverage analysis of safety-critical software in an industrial setting. Journal of Automated Reasoning, 45(4):397–414, December 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9172-3.

duBousquet:2010:RJS

[836] Lydie du Bousquet, Yves Ledru, Olivier Maury, Catherine Oriat, and Jean-Louis Lanet. Reusing a JML specification dedicated to verification for testing, and vice-versa: Case studies. Journal of Automated Reasoning, 45 (4):415–435, December 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-009-9132-y.

Longuet:2010:PGT

[837] Delphine Longuet, Marc Aiguier, and Pascale Le Gall. Proof-guided test selection from first-order specifications with equality. Journal of Automated Reasoning, 45(4):437–473, December 2010. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-009-
9128-7.

Areces:2011:ROS

[838] Carlos Areces and Daniel Gorín. Resolution with order and selection for hybrid logics. Journal of Automated Reasoning, 46(1):1–42, January 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9167-0.

Klarman:2011:AAD

[839] Szymon Klarman, Ulle Endriss, and Stefan Schlobach. ABox abduction in the description logic ALC. Journal of Automated Reasoning, 46(1): 43–80, January 2011. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9168-z.

Larrosa:2011:FCB

[840] Javier Larrosa, Robert Nieuwenhuis, Albert Oliveras, and Enric Rodríguez-Carbonell. A framework for certified Boolean branch-and-bound optimization. Journal of Automated Reasoning, 46(1):81–102, January 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9176-z.

Mengshoel:2011:PSL

[841] Ole J. Mengshoel, Dan Roth, and David C. Wilkins. Portfolios in stochastic local search: Efficiently computing most probable explanations in Bayesian networks. Journal of Automated Reasoning, 46(2): 103–160, February 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9170-5.

Albert:2011:CFU

[842] Elvira Albert, Puri Arenas, Samir Genaim, and Germán Puebla. Closedform upper bounds in static cost analysis. Journal of Automated Reasoning, 46(2):161–203, February 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-010-9174-
1.

Hernando:2011:PML

[843] Antonio Hernando, Eugenio Roanes-Lozano, and Luis M. Laita. A polynomial model for logics with a prime power number of truth values. Journal of Automated Reasoning, 46(2): 205–221, February 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9191-0.

Bauer:2011:PSI

[844] Lujo Bauer, Sandro Etalle, Jerry den Hartog, and Luca Viganò. Preface of special issue on “Computer Security: Foundations and Automated Reasoning”. Journal of Automated Reasoning, 46(3–4):223–224, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-010-9184-z.pdf.

Cortier:2011:SSM

[845] Véronique Cortier, Steve Kremer, and Bogdan Warinschi. A survey of symbolic methods in computational analysis of cryptographic systems. Journal of Automated Reasoning, 46(3– 4):225–259, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9187-9.

Courant:2011:APA

[846] J. Courant, M. Daubignard, C. Ene, P. Lafourcade, and Y. Lakhnech. Automated proofs for asymmetric encryption. Journal of Automated Reasoning, 46(3–4):261–291, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9186-x.

Martinho:2011:ECP

[847] Joana Martinho and António Ravara. Encoding cryptographic primitives in a calculus with polyadic synchronisation. Journal of Automated Reasoning, 46(3–4):293–323, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9189-7.

Kusters:2011:RPA

[848] Ralf Küsters and Tomasz Truderung. Reducing protocol analysis with XOR to the XOR–Free case in the Horn theory based approach. Journal of Automated Reasoning, 46(3–4):325–352, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-010-
                                        9188-8.

Arsac:2011:MAP

[849] Wihem Arsac, Giampaolo Bella, Xavier Chantry, and Luca Compagna. Multiattacker protocol validation. Journal of Automated Reasoning, 46(3– 4):353–388, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9185-y.

Kanovich:2011:CPC

[850] Max Kanovich, Paul Rowe, and Andre Scedrov. Collaborative planning with confidentiality. Journal of Automated Reasoning, 46(3–4):389–421, April 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9190-1.

Gordon:2011:RTJ

[851] Michael J. C. Gordon, Matt Kaufmann, and Sandip Ray. The right tools for the job: Correctness of cone of influence reduction proved using ACL2 and HOL4. Journal of Automated Reasoning, 47(1):1–16, June 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9169-y.

Thangarajah:2011:CER

[852] John Thangarajah and Lin Padgham. Computationally effective reasoning about goal interactions. Journal of Automated Reasoning, 47(1):17–56, June 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9175-0.

Dennis:2011:UEP

[853] Louise Abigail Dennis, Ian Green, and Alan Smaill. The use of embeddings to provide a clean separation of term and annotation for higher order rippling. Journal of Automated Reasoning, 47(1):57–105, June 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9177-y.

Schmidt:2011:PSI

[854] Renate A. Schmidt and Brigitte Pientka. Preface: Special issue of selected extended papers of CADE-22. Journal of Automated Reasoning, 47 (2):107–109, August 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-010-9212-z.pdf.

Claessen:2011:AIF

[855] Koen Claessen and Ann Lillieström. Automated inference of finite unsatisfiability. Journal of Automated Reasoning, 47(2):111–132, August 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-010-9216-
8.

Fuhs:2011:PTD

[856] Carsten Fuhs, Jürgen Giesl, Michael Parting, Peter Schneider-Kamp, and Stephan Swiderski. Proving termination by dependency pairs and inductive theorem proving. Journal of Automated Reasoning, 47(2):133–160, August 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9215-9.

Bonacina:2011:DST

[857] Maria Paola Bonacina, Christopher A. Lynch, and Leonardo de Moura. On deciding satisfiability by theorem proving with speculative inferences. Journal of Automated Reasoning, 47(2): 161–189, August 2011. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9213-y.

Baumgartner:2011:CSM

[858] Peter Baumgartner and Uwe Waldmann. A combined superposition and model evolution calculus. Journal of Automated Reasoning, 47(2): 191–227, August 2011. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9214-x.

Martin-Mateos:2011:PPF

[859] Francisco Jesús Martín-Mateos, José Luis Ruiz-Reina, José Antonio Alonso, and María José Hidalgo. Proof pearl: a formal proof of Higman’s lemma in ACL2. Journal of Automated Reasoning, 47 (3):229–250, October 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9178-x.

Johansson:2011:CSI

[860] Moa Johansson, Lucas Dixon, and Alan Bundy. Conjecture synthesis for inductive theories. Journal of Automated Reasoning, 47(3):251–289, October 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9193-y.

Chaieb:2011:FPS

[861] Amine Chaieb. Formal power series. Journal of Automated Reasoning, 47 (3):291–318, October 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9195-9.

Ciolli:2011:CPC

[862] Gianni Ciolli, Graziano Gentili, and Marco Maggesi. A certified proof of the Cartan fixed point theorems. Journal of Automated Reasoning, 47 (3):319–336, October 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9198-6.

Giesl:2011:PSI

[863] Jürgen Giesl and Reiner Hähnle. Preface: Special issue of selected extended papers of IJCAR 2010. Journal of Automated Reasoning, 47(4): 337–339, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-011-9240-3.pdf.

Brillout:2011:ISC

[864] Angelo Brillout, Daniel Kroening, Philipp Rümmer, and Thomas Wahl. An interpolating sequent calculus for quantifier-free Presburger arithmetic. Journal of Automated Reasoning, 47 (4):341–367, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-011-9237-y.

Blanchette:2011:MIH

[865] Jasmin Christian Blanchette and Alexander Krauss. Monotonicity inference for higher-order formulas. Journal of Automated Reasoning, 47(4): 369–398, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9234-1.

deNivelle:2011:CLP

[866] Hans de Nivelle. Classical logic with partial functions. Journal of Automated Reasoning, 47(4):399–425, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9236-z.

Magka:2011:TED

[867] Despoina Magka, Yevgeny Kazakov, and Ian Horrocks. Tractable extensions of the description logic EL with numerical datatypes. Journal of Automated Reasoning, 47(4):427–450, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9235-0.

Backes:2011:ATH

[868] Julian Backes and Chad Edward Brown. Analytic tableaux for higherorder logic with choice. Journal of Automated Reasoning, 47(4):451–479, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9233-2.

Hirokawa:2011:DDR

[869] Nao Hirokawa and Aart Middeldorp. Decreasing diagrams and relative termination. Journal of Automated Reasoning, 47(4):481–501, December 2011. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10817-011-
                                        9238-x.

Zucchelli:2012:DRM

[870] Daniele Zucchelli and Enrica Nicolini. A decidability result for the model checking of infinite-state systems. Journal of Automated Reasoning, 48(1):1–42, January 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9192-z.

Felty:2012:H

[871] Amy Felty and Alberto Momigliano. Hybrid. Journal of Automated Reasoning, 48(1):43–105, January 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10817-010-9194-
                                        x.

Borralleras:2012:SML

[872] Cristina Borralleras, Salvador Lucas, Albert Oliveras, Enric Rodríguez-Carbonell, and Albert Rubio. SAT modulo linear arithmetic for solving polynomial constraints. Journal of Automated Reasoning, 48(1): 107–131, January 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9196-8.

Comon-Lundh:2012:SIS

[873] Hubert Comon-Lundh and Catherine Meadows. Special issue on security and rewriting foreword. Journal of Automated Reasoning, 48(2): 133–134, February 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-011-9242-1.pdf.

Anantharaman:2012:UMH

[874] Siva Anantharaman, Hai Lin, Christopher Lynch, Paliath Narendran, and Michael Rusinowitch. Unification modulo homomorphic encryption. Journal of Automated Reasoning, 48(2): 135–158, February 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9205-y.

Guttman:2012:SPS

[875] Joshua D. Guttman. State and progress in Strand spaces: Proving fair exchange. Journal of Automated Reasoning, 48(2):159–195, February 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9202-1.

Kremer:2012:RET

[876] Steve Kremer, Antoine Mercier, and Ralf Treinen. Reducing equational theories for the decision of static equivalence. Journal of Automated Reasoning, 48(2):197–217, February 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9203-0.

Ciobaca:2012:CKS

[877] Ştefan Ciobâcă, Stéphanie Delaune, and Steve Kremer. Computing knowledge in security protocols under convergent equational theories. Journal of Automated Reasoning, 48(2): 219–262, February 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9197-7.

Chevalier:2012:DES

[878] Yannick Chevalier and Michaël Rusinowitch. Decidability of equivalence of symbolic derivations. Journal of Automated Reasoning, 48(2): 263–292, February 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9199-5.

Echenim:2012:ISS

[879] Mnacho Echenim and Nicolas Peltier. An instantiation scheme for satisfiability modulo theories. Journal of Automated Reasoning, 48(3):293–362, March 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9200-3.

Otop:2012:UCV

[880] Jan Otop. E-unification with constants vs. general E-unification. Journal of Automated Reasoning, 48(3):363–390, March 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9201-2.

Benhamou:2012:DSA

[881] Belaïd Benhamou, Lionel Paris, and Pierre Siegel. Dealing with satisfiability and n-ary CSPs in a logical framework. Journal of Automated Reasoning, 48(3):391–417, March 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9204-z.

Verbeek:2012:PPF

[882] Freek Verbeek and Julien Schmaltz. Proof pearl: A formal proof of Dally and Seitz’ necessary and sufficient condition for deadlock-free routing in interconnection networks. Journal of Automated Reasoning, 48(4):419–439, April 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9206-x.

Cortier:2012:DCR

[883] Véronique Cortier and Stéphanie Delaune. Decidability and combination results for two notions of knowledge in security protocols. Journal of Automated Reasoning, 48(4):441–487, April 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-010-
9208-8.

Janicic:2012:AM

[884] Predrag Janičić, Julien Narboux, and Pedro Quaresma. The area method. Journal of Automated Reasoning, 48 (4):489–532, April 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9209-7.

Chang:2012:FDD

[885] Liang Chang, Zhongzhi Shi, Tianlong Gu, and Lingzhong Zhao. A family of dynamic description logics for representing and reasoning about actions. Journal of Automated Reasoning, 49(1):1–52, June 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-010-9210-1.

Codish:2012:SST

[886] Michael Codish, Jürgen Giesl, Peter Schneider-Kamp, and René Thiemann. SAT solving for termination proofs with recursive path orders and dependency pairs. Journal of Automated Reasoning, 49(1):53–93, June 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-010-
                                        9211-0.

Krauss:2012:PPR

[887] Alexander Krauss and Tobias Nipkow. Proof pearl: Regular expression equivalence and relation algebra. Journal of Automated Reasoning, 49(1):95–106, June 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-011-
                                        9223-4.

Wiedijk:2012:BRH

[888] Freek Wiedijk. Book review: Handbook of Practical Logic and Automated Reasoning, by John R. Harrison, Cambridge University Press, 2009. Journal of Automated Reasoning, 49(1): 107–109, June 2012. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-012-9251-8.pdf.

Fernandez:2012:PTA

[889] Maribel Fernández and Christian Urban. Preface: Theory and applications of abstraction, substitution and naming. Journal of Automated Reasoning, 49(2):111–114, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-011-9217-2.pdf.

Lakin:2012:EAS

[890] Matthew R. Lakin and Andrew M. Pitts. Encoding abstract syntax without fresh names. Journal of Automated Reasoning, 49(2):115–140, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9220-7.

Benton:2012:STT

[891] Nick Benton, Chung-Kil Hur, Andrew J. Kennedy, and Conor McBride. Strongly typed term representations in Coq. Journal of Automated Reasoning, 49(2):141–159, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9219-0.

Bonchi:2012:PEE

[892] Filippo Bonchi, Maria Grazia Buscemi, Vincenzo Ciancia, and Fabio Gadducci. A presheaf environment for the explicit fusion calculus. Journal of Automated Reasoning, 49(2):161–183, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9224-3.

Pollack:2012:CLN

[893] Randy Pollack, Masahiko Sato, and Wilmer Ricciotti. A canonical locally named representation of binding. Journal of Automated Reasoning, 49 (2):185–207, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9229-y.

Cheney:2012:FAC

[894] James Cheney, Michael Norrish, and René Vestergaard. Formalizing adequacy: A case study for higher-order abstract syntax. Journal of Automated Reasoning, 49(2):209–239, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9221-6.

Gacek:2012:TLL

[895] Andrew Gacek, Dale Miller, and Gopalan Nadathur. A two-level logic approach to reasoning about computations. Journal of Automated Reasoning, 49(2):241–273, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9218-1.

Rose:2012:ESN

[896] Kristoffer H. Rose, Roel Bloo, and Frédéric Lang. On explicit substitution with names. Journal of Automated Reasoning, 49(2):275–300, August 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9222-5.

Pierce:2012:P

[897] Benjamin C. Pierce and Stephanie Weirich. Preface. Journal of Automated Reasoning, 49(3):301–302, October 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10817-012-9254-5.

Berghofer:2012:SPC

[898] Stefan Berghofer. A solution to the PoplMark challenge using de Bruijn indices in Isabelle/HOL. Journal of Automated Reasoning, 49(3): 303–326, October 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9231-4.

Vouillon:2012:SPC

[899] Jérôme Vouillon. A solution to the PoplMark challenge based on de Bruijn indices. Journal of Automated Reasoning, 49(3):327–362, October 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9230-5.

Chargueraud:2012:LNR

[900] Arthur Charguéraud. The locally nameless representation. Journal of Automated Reasoning, 49(3):363– 408, October 2012. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9225-2.

Hirschowitz:2012:NAS

[901] André Hirschowitz and Marco Maggesi. Nested abstract syntax in Coq. Journal of Automated Reasoning, 49 (3):409–426, October 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-010-9207-9.

Asperti:2012:FMP

[902] Andrea Asperti, Wilmer Ricciotti, Claudio Sacerdoti Coen, and Enrico Tassi. Formal metatheory of programming languages in the Matita interactive theorem prover. Journal of Automated Reasoning, 49(3):427–451, October 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9228-z.

Appel:2012:LMB

[903] Andrew W. Appel, Robert Dockins, and Xavier Leroy. A list-machine benchmark for mechanized metatheory. Journal of Automated Reasoning, 49 (3):453–491, October 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9226-1.

Schockaert:2012:SCL

[904] Steven Schockaert, Jeroen Janssen, and Dirk Vermeir. Satisfiability checking in Łukasiewicz logic as finite constraint satisfaction. Journal of Automated Reasoning, 49(4):493–550, December 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9227-0.

Tao:2012:PPT

[905] Jia Tao, Giora Slutzki, and Vasant Honavar. PS PACE tableau algorithms for acyclic modalized ALC. Journal of Automated Reasoning, 49(4): 551–582, December 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9232-3.

Jarvisalo:2012:SCL

[906] Matti Järvisalo, Armin Biere, and Marijn J. H. Heule. Simulating circuitlevel simplifications on CNF. Journal of Automated Reasoning, 49(4): 583–619, December 2012. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-011-9239-9.

Gaintzarain:2013:IFC

[907] Jose Gaintzarain, Montserrat Hermo, Paqui Lucio, Marisa Navarro, and Fernando Orejas. Invariant-free clausal temporal resolution. Journal of Automated Reasoning, 50(1):1–49, January 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9241-2.

Bofill:2013:PNM

[908] Miquel Bofill and Albert Rubio. Paramodulation with non-monotonic orderings and simplification. Journal of Automated Reasoning, 50(1): 51–98, January 2013. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-011-9244-z.

Bridge:2013:CSA

[909] James P. Bridge and Lawrence Charles Paulson. Case splitting in an automatic theorem prover for real-valued special functions. Journal of Automated Reasoning, 50(1):99–117, January 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-012-
                                        9245-6.

Trybulec:2013:FMM

[910] Andrzej Trybulec, Artur Kornilowicz, Adam Naumowicz, and Krystyna Kuperberg. Formal mathematics for mathematicians. Journal of Automated Reasoning, 50(2):119–121, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        012-9268-z.pdf.

Alama:2013:EIA

[911] Jesse Alama. Eliciting implicit assumptions of Mizar proofs by property omission. Journal of Automated Reasoning, 50(2):123–133, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s10817-012-
9264-3.

Belinfante:2013:GP

[912] Johan G. F. Belinfante. The GOEDEL program. Journal of Automated Reasoning, 50(2):135–146, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-012-
9259-0.

Caminati:2013:CAM

[913] Marco Bright Caminati and Giuseppe Rosolini. Custom automations in Mizar. Journal of Automated Reasoning, 50(2):147–160, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-012-9266-
1.

Futa:2013:FDT

[914] Yuichi Futa, Hiroyuki Okazaki, and Yasunari Shidama. Formalization of definitions and theorems related to an elliptic curve over a finite prime field by using Mizar. Journal of Automated Reasoning, 50(2):161–172, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
012-9265-2.pdf.

Harrison:2013:HLT

[915] John Harrison. The HOL Light theory of Euclidean space. Journal of Automated Reasoning, 50(2): 173–190, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-012-9250-9.

Iancu:2013:MML

[916] Mihnea Iancu, Michael Kohlhase, Florian Rabe, and Josef Urban. The Mizar mathematical library in OM-Doc: Translation and applications. Journal of Automated Reasoning, 50 (2):191–202, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-012-9271-4.

owicz:2013:RRM

[917] Artur Kornił owicz. On rewriting rules in Mizar. Journal of Automated Reasoning, 50(2):203–210, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        012-9261-6.pdf.

Mycielski:2013:FT

[918] Jan Mycielski. On the formalization of theories. Journal of Automated Reasoning, 50(2):211–216, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-012-
                                        9247-4.

Pak:2013:MLE

[919] Karol Pak. Methods of lemma extraction in natural deduction proofs. Journal of Automated Reasoning, 50 (2):217–228, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-012-9267-0.pdf.

Urban:2013:APS

[920] Josef Urban, Piotr Rudnicki, and Geoff Sutcliffe. ATP and presentation service for Mizar formalizations. Journal of Automated Reasoning, 50(2): 229–241, February 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-012-9269-y.

Barrett:2013:YSC

[921] Clark Barrett, Morgan Deters, Leonardo de Moura, Albert Oliveras, and Aaron Stump. 6 years of SMT–COMP. Journal of Automated Reasoning, 50 (3):243–277, March 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-012-9246-5.

Hirokawa:2013:UTC

[922] Nao Hirokawa, Aart Middeldorp, and Harald Zankl. Uncurrying for termination and complexity. Journal of Automated Reasoning, 50(3):279–315, March 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
012-9248-3.pdf.

Winkler:2013:MCT

[923] Sarah Winkler, Haruhiko Sato, Aart Middeldorp, and Masahito Kurihara. Multi-completion with termination tools. Journal of Automated Reasoning, 50(3):317–354, March 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-012-9249-2.pdf.

Gore:2013:ETU

[924] Rajeev Goré and Linh Anh Nguyen. ExpTime tableaux for using sound global caching. Journal of Automated Reasoning, 50(4):355–381, April 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-011-
9243-0.

Sabel:2013:TVL

[925] David Sabel and Manfred Schmidt-Schauß. A two-valued logic for properties of strict functional programs allowing partial functions. Journal of Automated Reasoning, 50(4):383–421, April 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-012-
9253-6.

Boldo:2013:WEN

[926] Sylvie Boldo, François Clément, Jean-Christophe Filliâtre, Micaela Mayero, Guillaume Melquiond, and Pierre Weis. Wave equation numerical resolution: a comprehensive mechanized proof of a C program. Journal of Automated Reasoning, 50(4):423–456, April 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-012-
9255-4.

Bjorner:2013:PSI

[927] Nikolaj Bjørner and Viorica Sofronie-Stokkermans. Preface: Special issue of selected extended papers of CADE-23. Journal of Automated Reasoning, 51(1):1–2, June 2013. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-013-9282-9.

Galmiche:2013:CBC

[928] Didier Galmiche and Daniel Méry. A connection-based characterization of bi-intuitionistic validity. Journal of Automated Reasoning, 51(1):3–26, June 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9279-4.

Noschinski:2013:AIR

[929] Lars Noschinski, Fabian Emmes, and Jürgen Giesl. Analyzing innermost runtime complexity of term rewriting by dependency pairs. Journal of Automated Reasoning, 51(1):27–56, June 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9277-6.

Brown:2013:RHO

[930] Chad E. Brown. Reducing higherorder theorem proving to a sequence of SAT problems. Journal of Automated Reasoning, 51(1):57–77, June 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9283-8.

Jovanovic:2013:CC

[931] Dejan Jovanović and Leonardo de Moura. Cutting to the chase. Journal of Automated Reasoning, 51(1):79–108, June 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-013-
                                        9281-x.

Blanchette:2013:ESS

[932] Jasmin Christian Blanchette, Sascha Böhme, and Lawrence C. Paulson. Extending Sledgehammer with SMT solvers. Journal of Automated Reasoning, 51(1):109–128, June 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10817-013-9278-
                                        5.

Ferrari:2013:CFL

[933] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. Contraction-free linear depth sequent calculi for intuitionistic propositional logic with the subformula property and minimal depth counter-models. Journal of Automated Reasoning, 51(2):129–149, August 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-012-
                                        9252-7.

Munoz:2013:FBP

[934] César Muñoz and Anthony Narkawicz. Formalization of Bernstein polynomials and applications to global optimization. Journal of Automated Reasoning, 51(2):151–196, August 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-012-9256-3.

Cerrito:2013:TBD

[935] Serenella Cerrito and Marta Cialdea Mayer. A tableau based decision procedure for an expressive fragment of hybrid logic with binders, converse and global modalities. Journal of Automated Reasoning, 51(2):197–239, August 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-012-
9257-2.

Ray:2013:SVC

[936] Sandip Ray and Rob Sumners. Specification and verification of concurrent programs through refinements. Journal of Automated Reasoning, 51 (3):241–280, October 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-012-9258-1.

Creus:2013:NLR

[937] Carles Creus, Guillem Godoy, Francesc Massanes, and Ashish Tiwari. Nonlinear rewrite closure and weak normalization. Journal of Automated Reasoning, 51(3):281–324, October 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-012-
9262-5.

Lynch:2013:SSM

[938] Christopher Lynch, Quang-Trung Ta, and Duc-Khanh Tran. SMELS: Satisfiability modulo equality with lazy superposition. Journal of Automated Reasoning, 51(3):325–356, October 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-012-
                                        9263-4.

Sternagel:2013:PPM

[939] Christian Sternagel. Proof pearl— a mechanized proof of GHC’s Mergesort. Journal of Automated Reasoning, 51 (4):357–370, December 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-012-9260-7.pdf.

Creus:2013:EFT

[940] Carles Creus, Adrià Gascón, and Guillem Godoy. Emptiness and finiteness for tree automata with global reflexive disequality constraints. Journal of Automated Reasoning, 51(4): 371–400, December 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-012-9270-5.

Baaz:2013:FVS

[941] Matthias Baaz, Ori Lahav, and Anna Zamansky. Finite-valued semantics for canonical labelled calculi. Journal of Automated Reasoning, 51(4): 401–430, December 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-013-9273-x.

Lisitsa:2013:FRS

[942] Alexei Lisitsa. Finite reasons for safety. Journal of Automated Reasoning, 51 (4):431–451, December 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-013-9274-9.

Barrett:2013:BRD

[943] Clark Barrett. Book review: Decision Procedures: An Algorithmic Point of View, by Daniel Kroening and Ofer Strichman, Springer–Verlag, 2008. Journal of Automated Reasoning, 51 (4):453–456, December 2013. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-013-9295-4.

Omodeo:2014:SGI

[944] Eugenio G. Omodeo and Alexandru I. Tomescu. Set graphs. III. Proof pearl: Claw-free graphs mirrored into transitive hereditarily finite sets. Journal of Automated Reasoning, 52(1):1–29, January 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-012-
9272-3.

Gwynne:2014:GUR

[945] Matthew Gwynne and Oliver Kullmann. Generalising unit-refutation completeness and SLUR via nested input resolution. Journal of Automated Reasoning, 52(1):31–65, January 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9275-8.

Fiorino:2014:TCP

[946] Guido Fiorino. Terminating calculi for propositional Dummett logic with subformula property. Journal of Automated Reasoning, 52(1):67–97, January 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-013-
                                        9276-7.

Hidalgo:2014:RPC

[947] Marta R. Hidalgo and Robert Joan-Arinyo. The reachability problem in constructive geometric constraint solving based dynamic geometry. Journal of Automated Reasoning, 52(1): 99–122, January 2014. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-013-9280-y.

Ballarin:2014:LMS

[948] Clemens Ballarin. Locales: A module system for mathematical theories. Journal of Automated Reasoning, 52 (2):123–153, February 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-013-9284-7.

Kutsia:2014:AUU

[949] Temur Kutsia, Jordi Levy, and Mateu Villaret. Anti-unification for unranked terms and hedges. Journal of Automated Reasoning, 52(2): 155–190, February 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-013-9285-6.pdf.

Alama:2014:PSM

[950] Jesse Alama, Tom Heskes, Daniel Kühlwein, Evgeni Tsivtsivadze, and Josef Urban. Premise selection for mathematics by corpus analysis and kernel methods. Journal of Automated Reasoning, 52(2):191–213, February 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9286-5.

Park:2014:MMT

[951] Jonghyun Park, Jeongbong Seo, Sungwoo Park, and Gyesik Lee. Mechanizing metatheory without typing contexts. Journal of Automated Reasoning, 52(2):215–239, February 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9287-4.

Alkassar:2014:FVC

[952] Eyad Alkassar, Sascha Böhme, Kurt Mehlhorn, and Christine Rizkallah. A framework for the verification of certifying computations. Journal of Automated Reasoning, 52(3):241–273, March 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9289-2.

Malkis:2014:AVS

[953] Alexander Malkis and Anindya Banerjee. On automation in the verification of software barriers: Experience report. Journal of Automated Reasoning, 52(3):275–329, March 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-013-9290-9.

Hidalgo-Doblado:2014:FVT

[954] M. J. Hidalgo-Doblado, J. A. Alonso-Jiménez, J. Borrego-Díaz, F. J. Martín-Mateos, and J. L. Ruiz-Reina. Formally verified tableau-based reasoners for a description logic. Journal of Automated Reasoning, 52(3):331–360, March 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-013-
                                        9291-8.

Stannett:2014:UIH

[955] Mike Stannett and István Németi. Using Isabelle/HOL to verify first-order relativity theory. Journal of Automated Reasoning, 52(4):361–378, April 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-013-
                                        9292-7.

Zhou:2014:ATB

[956] Min Zhou, Fei He, Bow-Yaw Wang, Ming Gu, and Jiaguang Sun. Array theory of bounded elements and its applications. Journal of Automated Reasoning, 52(4):379–405, April 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-013-
                                        9293-6.

Kaminski:2014:GDD

[957] Mark Kaminski and Gert Smolka. A goal-directed decision procedure for hybrid PDL. Journal of Automated Reasoning, 52(4):407–450, April 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9294-5.

Wu:2014:FMN

[958] Chunhan Wu, Xingyuan Zhang, and Christian Urban. A formalisation of the Myhill–Nerode Theorem based on regular expressions. Journal of Automated Reasoning, 52(4):451–480, April 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9297-2.

Kazakov:2014:IE

[959] Yevgeny Kazakov, Markus Krötzsch, and František Simančík. The incredible ELK. Journal of Automated Reasoning, 53(1):1–61, June 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-013-9296-3.

Affeldt:2014:FST

[960] Reynald Affeldt, Manabu Hagiwara, and Jonas Sénizergues. Formalization of Shannon’s theorems. Journal of Automated Reasoning, 53(1):63–103, June 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-013-
9298-1.

Zsido:2014:TTC

[961] Julianna Zsidó. Theorem of three circles in Coq. Journal of Automated Reasoning, 53(2):105–127, August 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-013-
                                        9299-0.

deGouw:2014:PPK

[962] Stijn de Gouw, Frank de Boer, and Jurriaan Rot. Proof pearl: The KeY to correct and stable sorting. Journal of Automated Reasoning, 53 (2):129–139, August 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-013-9300-y.

Bridge:2014:MLF

[963] James P. Bridge, Sean B. Holden, and Lawrence C. Paulson. Machine learning for first-order theorem proving. Journal of Automated Reasoning, 53 (2):141–172, August 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-014-9301-5.

Kaliszyk:2014:LAA

[964] Cezary Kaliszyk and Josef Urban. Learning-assisted automated reasoning with Flyspeck. Journal of Automated Reasoning, 53(2):173–213, August 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s10817-

014-9303-3.pdf. See erratum [975].

Wu:2014:AA

[965] Jiewen Wu, Alexander Hudek, David Toman, and Grant Weddell. Absorption for ABoxes. Journal of Automated Reasoning, 53(3):215–243, October 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-014-
9304-2.

Glimm:2014:HOR

[966] Birte Glimm, Ian Horrocks, Boris Motik, Giorgos Stoilos, and Zhe Wang. HermiT: An OWL 2 reasoner. Journal of Automated Reasoning, 53(3): 245–269, October 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-014-9305-1.

Braibant:2014:IRA

[967] Thomas Braibant, Jacques-Henri Jourdan, and David Monniaux. Implementing and reasoning about hash-consed data structures in Coq. Journal of Automated Reasoning, 53(3):271–304, October 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-014-
9306-0.

Mayer:2014:EDP

[968] Marta Cialdea Mayer. Extended decision procedure for a fragment of HL with binders. Journal of Automated Reasoning, 53(3):305–315, October 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-014-

9307-z. See erratum [988].

Bensaid:2014:CSC

[969] Hicham Bensaid and Nicolas Peltier. A complete superposition calculus for primal grammars. Journal of Automated Reasoning, 53(4):317–350, December 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-014-
                                        9309-x.

Steigmiller:2014:RNS

[970] Andreas Steigmiller, Birte Glimm, and Thorsten Liebig. Reasoning with nominal schemas through absorption. Journal of Automated Reasoning, 53(4): 351–405, December 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-014-9310-4.

Siddique:2014:FGF

[971] Umair Siddique and Osman Hasan. On the formalization of gamma function in HOL. Journal of Automated Reasoning, 53(4):407–429, December 2014. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-014-
                                        9311-3.

Martin-Dorel:2015:FVC

[972] Érik Martin-Dorel, Guillaume Hanrot, Micaela Mayero, and Laurent Théry. Formally verified certificate checkers for hardest-to-round computation. Journal of Automated Reasoning, 54(1):1–29, January 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-014-9312-2.

Ghosh:2015:APE

[973] Kamalesh Ghosh, Pallab Dasgupta, and S. Ramesh. Automated planning as an early verification tool for distributed control. Journal of Automated Reasoning, 54(1):31–68, January 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-014-
9313-1.

Bonacina:2015:IAT

[974] Maria Paola Bonacina and Moa Johansson. On interpolation in automated theorem proving. Journal of Automated Reasoning, 54(1):69–97, January 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-014-
9314-0.

Kaliszyk:2015:ELA

[975] Cezary Kaliszyk and Josef Urban. Erratum to: Learning-Assisted Automated Reasoning with Flyspeck. Journal of Automated Reasoning, 54(1):99, January 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s10817-014-
9315-z.pdf. See [964].

Zankl:2015:LDD

[976] Harald Zankl, Bertram Felgenhauer, and Aart Middeldorp. Labelings for decreasing diagrams. Journal of Automated Reasoning, 54(2):101– 133, February 2015. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-014-9316-y.pdf.

Boldo:2015:VCF

[977] Sylvie Boldo, Jacques-Henri Jourdan, Xavier Leroy, and Guillaume Melquiond. Verified compilation of floating-point computations. Journal of Automated Reasoning, 54(2): 135–163, February 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-014-9317-x.

Foster:2015:FSR

[978] Simon Foster and Georg Struth. On the fine-structure of regular algebra. Journal of Automated Reasoning, 54 (2):165–197, February 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-014-9318-9.

Reus:2015:SEP

[979] Bernhard Reus, Nathaniel Charlton, and Ben Horsfall. Symbolic execution proofs for higher order store programs. Journal of Automated Reasoning, 54 (3):199–284, March 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-014-9319-8.

Narkawicz:2015:FVD

[980] Anthony Narkawicz, César Muñoz, and Aaron Dutle. Formally-verified decision procedures for univariate polynomial computation based on Sturm’s and Tarski’s theorems. Journal of Automated Reasoning, 54(4):285–326, April 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9320-x.

Alberti:2015:DPF

[981] Francesco Alberti, Silvio Ghilardi, and Natasha Sharygina. Decision procedures for flat array properties. Journal of Automated Reasoning, 54(4): 327–352, April 2015. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-015-9323-7.

Bonacina:2015:ISG

[982] Maria Paola Bonacina and Moa Johansson. Interpolation systems for ground proofs in automated deduction: a survey. Journal of Automated Reasoning, 54(4):353–390, April 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9325-5.

Paulson:2015:MPG

[983] Lawrence C. Paulson. A mechanised proof of Gödel’s incompleteness theorems using nominal Isabelle. Journal of Automated Reasoning, 55(1):1–37, June 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9322-8.

Botana:2015:ATP

[984] Francisco Botana, Markus Hohenwarter, Predrag Janičić, Zoltán Kovács, Ivan Petrović, Tomás Recio, and Simon Weitzhofer. Automated theorem proving in GeoGebra: Current achievements. Journal of Automated Reasoning, 55(1):39–59, June 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-015-
                                        9326-4.

Cok:2015:ESC

[985] David R. Cok, Aaron Stump, and Tjark Weber. The 2013 evaluation of SMT–COMP and SMT–LIB. Journal of Automated Reasoning, 55(1):61–90, June 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-015-
                                        9328-2.

Kuhlwein:2015:MFA

[986] Daniel Kühlwein and Josef Urban. MaLeS: A framework for automatic tuning of automated theorem provers. Journal of Automated Reasoning, 55 (2):91–116, August 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-015-9329-1.pdf.

Davis:2015:RMT

[987] Jared Davis and Magnus O. Myreen. The reflective Milawa theorem prover is sound (down to the machine code that runs it). Journal of Automated Reasoning, 55(2):117–183, August 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9324-6.

Mayer:2015:EED

[988] Marta Cialdea Mayer. Erratum to: “Extension of a Decision Procedure for a Fragment of Hybrid Logic with Binders”. Journal of Automated Reasoning, 55(2):185, August 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
014-9308-y.pdf. See [968].

Kuperberg:2015:ATM

[989] Krystyna Trybulec Kuperberg. Andrzej Trybulec — in memoriam. Journal of Automated Reasoning, 55(3): 187–190, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-015-9343-3.

Grabowski:2015:FDM

[990] Adam Grabowski, Artur Kornił owicz, and Adam Naumowicz. Four decades of Mizar. Journal of Automated Reasoning, 55(3):191–198, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
015-9345-1.pdf.

Brown:2015:RPF

[991] Chad E. Brown. Reconsidering pairs and functions as sets. Journal of Automated Reasoning, 55(3):199–210, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-015-
                                        9340-6.

Grabowski:2015:MCL

[992] Adam Grabowski. Mechanizing complemented lattices within Mizar type system. Journal of Automated Reasoning, 55(3):211–221, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        content/pdf/10.1007/s10817-015-
                                        9333-5.pdf.

Harrison:2015:FPH

[993] John Harrison. Formal proofs of hypergeometric sums. Journal of Automated Reasoning, 55(3):223–243, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-015-
                                        9338-0.

Kaliszyk:2015:MM

[994] Cezary Kaliszyk and Josef Urban. MizAR 40 for Mizar 40. Journal of Automated Reasoning, 55(3):245–256, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        015-9330-8.pdf.

owicz:2015:DEM

[995] Artur Kornił owicz. Definitional expansions in Mizar. Journal of Automated Reasoning, 55(3):257–268, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.

com/content/pdf/10.1007/s10817-
015-9331-7.pdf.

Lyaletski:2015:EAI

[996] Alexander Lyaletski. Evidence algorithm and inference search in firstorder logics. Journal of Automated Reasoning, 55(3):269–284, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9346-0.

Naumowicz:2015:ABS

[997] Adam Naumowicz. Automating Boolean set operations in Mizar proof checking with the aid of an external SAT solver. Journal of Automated Reasoning, 55(3):285–294, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
015-9332-6.pdf.

Pak:2015:ILF

[998] Karol Pak. Improving legibility of formal proofs based on the close reference principle is NP–Hard. Journal of Automated Reasoning, 55(3): 295–306, October 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-015-9337-1.pdf.

Felty:2015:NCP

[999] Amy P. Felty, Alberto Momigliano, and Brigitte Pientka. The next 700 challenge problems for reasoning with higher-order abstract syntax representations. Journal of Automated Reasoning, 55(4):307–372, December 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-015-
                                        9327-3.

Asperti:2015:RC

[1000] Andrea Asperti. Reverse complexity. Journal of Automated Reasoning, 55 (4):373–388, December 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-015-9349-x.

Benzmuller:2015:HOP

[1001] Christoph Benzmüller, Nik Sultana, Lawrence C. Paulson, and Frank Theiß. The higher-order prover Leo-II. Journal of Automated Reasoning, 55(4): 389–404, December 2015. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-015-9348-y.pdf.

Bengtson:2016:PCI

[1002] Jesper Bengtson, Joachim Parrow, and Tjark Weber. Psi-calculi in Isabelle. Journal of Automated Reasoning, 56 (1):1–47, January 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-015-9336-2.

Gagne:2016:APB

[1003] Martin Gagné, Pascal Lafourcade, Yassine Lakhnech, and Reihaneh Safavi-Naini. Automated proofs of block cipher modes of operation. Journal of Automated Reasoning, 56(1):

49–94, January 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-015-9341-5.

Baumgartner:2016:MMS

[1004] Peter Baumgartner, Wolfgang Bibel, and Richard Waldinger. In memory of Mark Stickel. Journal of Automated Reasoning, 56(2):95–98, February 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9347-z.

Loveland:2016:MSH

[1005] Donald W. Loveland. Mark stickel:

His earliest work. Journal of Automated Reasoning, 56(2):99–112, February 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9342-4.

Bonacina:2016:SGG

[1006] Maria Paola Bonacina and David A.

Plaisted. Semantically-guided goalsensitive reasoning: Model representation. Journal of Automated Reasoning, 56(2):113–141, February 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-015-
9334-4.

Zhang:2016:ESM

[1007] Hantao Zhang. An experiment with satisfiability modulo SAT. Journal of Automated Reasoning, 56(2):143– 154, February 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-015-9354-0.

Blanchette:2016:SII

[1008] Jasmin Christian Blanchette, Sascha Böhme, Mathias Fleury, Steffen Juilf Smolka, and Albert Steckermeier. Semi-intelligible Isar proofs from machine-generated proofs. Journal of Automated Reasoning, 56(2):155– 200, February 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-015-9335-3.

Klein:2016:ITP

[1009] Gerwin Klein and Ruben Gamboa. Interactive theorem proving. Journal of Automated Reasoning, 56(3):201–203, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        016-9363-7.pdf.

Arthan:2016:DCT

[1010] Rob Arthan. On definitions of constants and types in HOL. Journal of Automated Reasoning, 56(3):205–219, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        016-9366-4.pdf.

Kumar:2016:SFH

[1011] Ramana Kumar, Rob Arthan, Magnus O. Myreen, and Scott Owens. Self-formalisation of higher-order logic. Journal of Automated Reasoning, 56 (3):221–259, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-015-9357-x.pdf.

Matichuk:2016:EPM

[1012] Daniel Matichuk, Toby Murray, and Makarius Wenzel. Eisbach: A proof method language for Isabelle. Journal of Automated Reasoning, 56(3): 261–282, March 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-015-9360-2.

Blazy:2016:VAI

[1013] Sandrine Blazy, Vincent Laporte, and David Pichardie. Verified abstract interpretation techniques for disassembling low-level self-modifying code. Journal of Automated Reasoning, 56 (3):283–308, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-015-9359-8.

Bourke:2016:MPA

[1014] Timothy Bourke, Robert J. van Glabbeek, and Peter Höfner. Mechanizing a process algebra for network protocols. Journal of Automated Reasoning, 56(3):309–341, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-015-9358-
9.

Doczkal:2016:CDR

[1015] Christian Doczkal and Gert Smolka.

Completeness and decidability results for CTL in constructive type theory. Journal of Automated Reasoning, 56 (3):343–365, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-016-9361-9.

Avigad:2016:HPR

[1016] Jeremy Avigad, Robert Y. Lewis, and Cody Roux. A heuristic prover for real inequalities. Journal of Automated Reasoning, 56(3):367–386, March 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-015-
                                        9356-y.

Dross:2016:ADP

[1017] Claire Dross, Sylvain Conchon, Johannes Kanig, and Andrei Paskevich. Adding decision procedures to SMT solvers using axioms with triggers. Journal of Automated Reasoning, 56 (4):387–457, April 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-015-9352-2.

Slivovsky:2016:QRQ

[1018] Friedrich Slivovsky and Stefan Szeider. Quantifier reordering for QBF. Journal of Automated Reasoning, 56 (4):459–477, April 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-015-9353-1.

Kroening:2016:PSI

[1019] Daniel Kroening and Andrey Rybalchenko. Preface: Special issue on interpolation. Journal of Automated Reasoning, 57(1):1–2, June 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10817-016-9375-
3.

Schlaipfer:2016:LIS

[1020] Matthias Schlaipfer and Georg Weissenbacher. Labelled interpolation systems for hyper-resolution, clausal, and local proofs. Journal of Automated Reasoning, 57(1):3–36, June 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
016-9364-6.pdf.

Totla:2016:CIB

[1021] Nishant Totla and Thomas Wies. Complete instantiation-based interpolation. Journal of Automated Reasoning, 57 (1):37–65, June 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-016-9371-7.

Christ:2016:PTP

[1022] Jürgen Christ and Jochen Hoenicke.

Proof tree preserving tree interpolation. Journal of Automated Reasoning, 57(1):67–95, June 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-016-9365-5.

Echenim:2016:SCA

[1023] M. Echenim and N. Peltier. A superposition calculus for abductive reasoning.

Journal of Automated Reasoning, 57 (2):97–134, August 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-015-9344-2.

Roux:2016:FPR

[1024] Pierre Roux. Formal proofs of rounding error bounds: With application to an automatic positive definiteness check. Journal of Automated Reasoning, 57 (2):135–156, August 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10817-015-9339-
z.

Franek:2016:QDF

[1025] Peter Franek, Stefan Ratschan, and Piotr Zgliczynski. Quasi-decidability of a fragment of the first-order theory of real numbers. Journal of Automated Reasoning, 57(2):157–185, August 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10817-015-9351-3.

Martin-Dorel:2016:PTB

[1026] Érik Martin-Dorel and Guillaume Melquiond. Proving tight bounds on univariate expressions with elementary functions in Coq. Journal of Automated Reasoning, 57(3):187– 217, October 2016. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-015-9350-4.

Blanchette:2016:LBF

[1027] Jasmin Christian Blanchette, David Greenaway, Cezary Kaliszyk, Daniel Kühlwein, and Josef Urban. A learning-based fact selector for Isabelle/HOL. Journal of Automated Reasoning, 57(3):219–244, October 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10817-016-9362-8.

BoydelaTour:2016:PGL

[1028] Thierry Boy de la Tour and Nicolas Peltier. Proof generalization in LK by second order unifier minimization. Journal of Automated Reasoning, 57 (3):245–280, October 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10817-016-9367-
3.

Pham:2016:RAA

[1029] Tuan-Hung Pham, Andrew Gacek, and Michael W. Whalen. Reasoning about algebraic data types with abstractions. Journal of Automated Reasoning, 57 (4):281–318, December 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s10817-016-9368-
2.

Krebbers:2016:FCM

[1030] Robbert Krebbers. A formal C memory model for separation logic. Journal of Automated Reasoning, 57(4): 319–387, December 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-016-9369-1.pdf.

Demri:2017:P

[1031] Stéphane Demri, Deepak Kapur, and Christoph Weidenbach. Preface. Journal of Automated Reasoning, 58(1): 1–2, January 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-016-9394-0.pdf.

Giesl:2017:APT

[1032] Jürgen Giesl, Cornelius Aschermann, Marc Brockschmidt, Fabian Emmes, Florian Frohn, Carsten Fuhs, Jera Hensel, Carsten Otto, Martin Plücker, Peter Schneider-Kamp, Thomas Ströder, Stephanie Swiderski, and René Thiemann. Analyzing program termination and complexity automatically with AProVE. Journal of Automated Reasoning, 58(1): 3–31, January 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-016-9388-y.

Stroder:2017:APT

[1033] Thomas Ströder, Jürgen Giesl, Marc Brockschmidt, Florian Frohn, Carsten Fuhs, Jera Hensel, Peter Schneider-Kamp, and Cornelius Aschermann. Automatically proving termination and memory safety for programs with pointer arithmetic. Journal of Automated Reasoning, 58(1):33–65, January 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

com/accesspage/article/10.1007/
s10817-016-9389-x.

Ceylan:2017:BOL

[1034] İsmail İlkan Ceylan and Rafael Peñaloza. The Bayesian ontology language BEL. Journal of Automated Reasoning, 58(1):67–95, January 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10817-016-9386-0.

Heule:2017:SVE

[1035] Marijn J. H. Heule, Martina Seidl, and Armin Biere. Solution validation and extraction for QBF preprocessing. Journal of Automated Reasoning, 58 (1):97–125, January 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-016-9390-4.

Zeljic:2017:AFS

[1036] Aleksandar Zeljić, Christoph M. Wintersteiger, and Philipp Rümmer. An approximation framework for solvers and decision procedures. Journal of Automated Reasoning, 58(1):127– 147, January 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-016-9393-1.pdf.

Blanchette:2017:SCP

[1037] Jasmin Christian Blanchette, Andrei Popescu, and Dmitriy Traytel. Soundness and completeness proofs by coinductive methods. Journal of Automated Reasoning, 58(1):149–179, January 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10817-016-9391-3.

Beeson:2017:FPT

[1038] Michael Beeson and Larry Wos. Finding proofs in Tarskian geometry. Journal of Automated Reasoning, 58(1): 181–207, January 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10817-016-9392-2.

Braun:2017:SPP

[1039] Gabriel Braun and Julien Narboux. A synthetic proof of Pappus’ Theorem in Tarski’s geometry. Journal of Automated Reasoning, 58(2):209– 230, February 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10817-016-9374-4.

Rocha-Oliveira:2017:COT

[1040] Ana Cristina Rocha-Oliveira, André Luiz Galdino, and Mauricio Ayala-Rincón. Confluence of orthogonal term rewriting systems in the prototype verification system. Journal of Automated Reasoning, 58(2):231–251, February 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10817-016-9376-2.

Ganesalingam:2017:FAT

[1041] M. Ganesalingam and W. T. Gowers. A fully automatic theorem prover with human-style output. Journal of Automated Reasoning, 58(2):253– 291, February 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s10817-016-9377-1.pdf.

Baumgartner:2017:HOP

[1042] Alexander Baumgartner, Temur Kutsia, Jordi Levy, and Mateu Villaret. Higher-order pattern anti-unification in linear time. Journal of Automated Reasoning, 58(2):293–310, February 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
016-9383-3.pdf.

Felty:2017:PSE

[1043] Amy Felty and Aart Middeldorp.

Preface: Selected extended papers of CADE 2015. Journal of Automated Reasoning, 58(3):311–312, March 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s10817-
016-9403-3.pdf.

Zulkoski:2017:CSS

[1044] Edward Zulkoski, Curtis Bright, Albert Heinle, Ilias Kotsireas, Krzysztof Czarnecki, and Vijay Ganesh. Combining SAT solvers with computer algebra systems to verify combinatorial conjectures. Journal of Automated Reasoning, 58(3):313–339, March 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10817-016-9396-y.

Reynolds:2017:DPC

[1045] Andrew Reynolds and Jasmin Christian Blanchette. A decision procedure for (co)datatypes in SMT solvers. Journal of Automated Reasoning, 58(3): 341–362, March 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s10817-016-9372-6.

DSilva:2017:AIA

[1046] Vijay D’Silva and Caterina Urban. Abstract interpretation as automated deduction. Journal of Automated Reasoning, 58(3):363–390, March 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10817-016-9382-4.

Iborra:2017:RTD

[1047] José Iborra, Naoki Nishida, Germán Vidal, and Akihisa Yamada. Relative termination via dependency pairs. Journal of Automated Reasoning, 58 (3):391–411, March 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-016-9373-5.pdf.

Lagniez:2017:PTT

[1048] Jean-Marie Lagniez and Pierre Marquis. On preprocessing techniques and their impact on propositional model counting. Journal of Automated Reasoning, 58(4):413–481, April 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.

com/accesspage/article/10.1007/
s10817-016-9370-8.

Eberl:2017:PDC

[1049] Manuel Eberl. Proving divide and conquer complexities in Isabelle/HOL. Journal of Automated Reasoning, 58 (4):483–508, April 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-016-9378-0.

Aransay:2017:FHF

[1050] Jesús Aransay and Jose Divasón. A formalisation in HOL of the Fundamental Theorem of Linear Algebra and its application to the solution of the least squares problem. Journal of Automated Reasoning, 58(4): 509–535, April 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s10817-016-9379-z.

Giesl:2017:PSI

[1051] Jürgen Giesl and Jan Hoffmann. Preface: Special issue on automatic resource bound analysis. Journal of Automated Reasoning, 59(1):1–2, June 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9399-8;   http://link.springer.
com/content/pdf/10.1007/s10817-
016-9399-8.pdf.

Sinn:2017:CRB

[1052] Moritz Sinn, Florian Zuleger, and Helmut Veith. Complexity and resource bound analysis of imperative programs using difference constraints. Journal of Automated Reasoning, 59(1):3–45, June 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9402-4;   http://link.springer.
com/content/pdf/10.1007/s10817-
016-9402-4.pdf.

Albert:2017:RGT

[1053] Elvira Albert, Antonio Flores-Montoya, Samir Genaim, and Enrique Martin-Martin. Rely-guarantee termination and cost analyses of loops with concurrent interleavings. Journal of Automated Reasoning, 59(1):47–85, June 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9400-6.

Jost:2017:TBC

[1054] Steffen Jost, Pedro Vasconcelos, Mário Florido, and Kevin Hammond. Typebased cost analysis for lazy functional languages. Journal of Automated Reasoning, 59(1):87–120, June 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9398-9.

Frohn:2017:LBR

[1055] Florian Frohn, Jürgen Giesl, Jera Hensel, Cornelius Aschermann, and Thomas Ströder. Lower bounds for runtime complexity of term rewriting. Journal of Automated Reasoning, 59(1):121–163, June 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-016-9397-x.

Bonacina:2017:SGG

[1056] Maria Paola Bonacina and David A.

Plaisted. Semantically-guided goalsensitive reasoning: Inference system and completeness. Journal of Automated Reasoning, 59(2):165–218, August 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9384-2.

Platzer:2017:CUS

[1057] André Platzer. A complete uniform substitution calculus for differential dynamic logic. Journal of Automated Reasoning, 59(2):219–265, August 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9385-1;   http://link.springer.
com/content/pdf/10.1007/s10817-
016-9385-1.pdf.

Walther:2017:FEW

[1058] Christoph Walther and Nathan Wasser.

Fermat, Euler, Wilson — three case studies in number theory. Journal of Automated Reasoning, 59(2):267– 286, August 2017. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-016-9387-z.

Chihani:2017:SFP

[1059] Zakaria Chihani, Dale Miller, and Fabien Renaud. A semantic framework for proof evidence. Journal of Automated Reasoning, 59(3):287–330, October 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9380-6.

Zhou:2017:ARG

[1060] Jie Zhou, Dingkang Wang, and Yao Sun. Automated reducible geometric theorem proving and discovery by Gröbner basis method. Journal of Automated Reasoning, 59(3):331–344, October 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-016-
9395-z.

Holzl:2017:MCM

[1061] Johannes Hölzl. Markov chains and Markov decision processes in Isabelle/ HOL. Journal of Automated Reasoning, 59(3):345–387, October 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-016-9401-
5.

Avigad:2017:FVP

[1062] Jeremy Avigad, Johannes Hölzl, and Luke Serafin. A formally verified proof of the Central Limit Theorem. Journal of Automated Reasoning, 59(4): 389–423, December 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9404-x.

Cruz-Filipe:2017:FPS

[1063] Luís Cruz-Filipe, Kim S. Larsen, and Peter Schneider-Kamp. Formally proving size optimality of sorting networks. Journal of Automated Reasoning, 59 (4):425–454, December 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9405-9.

Parsia:2017:ORE

[1064] Bijan Parsia, Nicolas Matentzoglu, Rafael S. Gonçalves, Birte Glimm, and Andreas Steigmiller. The OWL Reasoner Evaluation (ORE) 2015 Competition Report. Journal of Automated Reasoning, 59(4):455–482, December 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9406-8;   http://link.springer.
com/content/pdf/10.1007/s10817-
017-9406-8.pdf.

Sutcliffe:2017:TPL

[1065] Geoff Sutcliffe. The TPTP problem library and associated infrastructure. Journal of Automated Reasoning, 59 (4):483–502, December 2017. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9407-7.

Munoz:2018:SEP

[1066] César A. Muñoz, Sanjai Rayadurgam, and Oksana Tkachuk. Selected extended papers of NFM 2016: Preface. Journal of Automated Reasoning, 60(1):1–2, January 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9425-5.

Brunner:2018:FVE

[1067] Julian Brunner and Peter Lammich. Formal verification of an executable LTL model checker with partial order reduction. Journal of Automated Reasoning, 60(1):3–21, January 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-017-
                                        9418-4.

He:2018:VRS

[1068] Shaobo He, Shuvendu K. Lahiri, and Zvonimir Rakamarić. Verifying relative safety, accuracy, and termination for program approximations. Journal of Automated Reasoning, 60(1): 23–42, January 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9421-9.

Jha:2018:SAU

[1069] Susmit Jha, Vasumathi Raman, Dorsa Sadigh, and Sanjit A. Seshia. Safe autonomy under perception uncertainty using chance-constrained temporal logic. Journal of Automated Reasoning, 60(1):43–62, January 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9413-9.

Newell:2018:TIF

[1070] Josh Newell, Linna Pang, David Tremaine, Alan Wassyng, and Mark Lawford. Translation of IEC 61131-3 function block diagrams to PVS for formal verification with real-time nuclear application. Journal of Automated Reasoning, 60(1):63–84, January 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9415-7.

Sardar:2018:TPF

[1071] Muhammad Usama Sardar, Nida Afaq, Osman Hasan, and Khaza Anuarul Hoque. Towards probabilistic formal analysis of SATS–Simultaneously Moving Aircraft (SATS–SMA). Journal of Automated Reasoning, 60(1): 85–105, January 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9416-6.

Wu:2018:SOP

[1072] Yi-Chin Wu, Vasumathi Raman, Blake C. Rawlings, Stéphane Lafortune, and Sanjit A. Seshia. Synthesis of obfuscation policies to ensure privacy and utility. Journal of Automated Reasoning, 60(1):107–131, January 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9420-x.

Slaney:2018:CRF

[1073] John Slaney and Bruno Woltzenlogel Paleo. Conflict resolution: A first-order resolution calculus with decision literals and conflict-driven clause learning. Journal of Automated Reasoning, 60 (2):133–156, February 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9408-6. See erra-

tum [1088].

Kocsis:2018:GPV

[1074] Zoltan A. Kocsis and Jerry Swan. Genetic programming + proof search = automatic improvement. Journal of Automated Reasoning, 60(2): 157–176, February 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9409-5;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-017-9409-5.pdf.

Bozzelli:2018:VLT

[1075] Laura Bozzelli and César Sánchez. Visibly linear temporal logic. Journal of Automated Reasoning, 60(2): 177–220, February 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9410-z.

Casal:2018:MSE

[1076] Filipe Casal and João Rasga. Manysorted equivalence of shiny and strongly polite theories. Journal of Automated Reasoning, 60(2):221– 236, February 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9411-y.

Maggesi:2018:FMS

[1077] Marco Maggesi. A formalization of metric spaces in HOL Light. Journal of Automated Reasoning, 60(2): 237–254, February 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9412-x.

Blazy:2018:SEP

[1078] Sandrine Blazy and Marsha Chechik.

Selected extended papers of VSTTE 2016. Journal of Automated Reasoning, 60(3):255–256, March 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9430-8.

Tan:2018:BGM

[1079] Gang Tan and Greg Morrisett. Bidirectional grammars for machine-code decoding and encoding. Journal of Automated Reasoning, 60(3):257–277, March 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9429-1.

Kojima:2018:AVF

[1080] Kensuke Kojima, Akifumi Imanishi, and Atsushi Igarashi. Automated verification of functional correctness of race-free GPU programs. Journal of Automated Reasoning, 60(3):279–298, March 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9428-2.

Beyer:2018:UVS

[1081] Dirk Beyer, Matthias Dangl, and Philipp Wendler. A unifying view on SMT-based software verification.

Journal of Automated Reasoning, 60 (3):299–335, March 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9432-6. See correc-

tion [1237].

Kiefer:2018:RPR

[1082] Moritz Kiefer, Vladimir Klebanov, and Mattias Ulbrich. Relational program reasoning using compiler IR. Journal of Automated Reasoning, 60(3):337–363, March 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9433-5.

Clochard:2018:MRV

[1083] Martin Clochard, Léon Gondelman, and Mário Pereira. The matrix reproved (verification pearl). Journal of Automated Reasoning, 60(3):365–383, March 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9436-2.

Matentzoglu:2018:ORS

[1084] Nicolas Matentzoglu, Bijan Parsia, and Uli Sattler. OWL reasoning: Subsumption test hardness and modularity. Journal of Automated Reasoning, 60(4):385–419, April 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9414-8;   http:
//link.springer.com/content/pdf/
10.1007/s10817-017-9414-8.pdf.

Aguirre:2018:SNC

[1085] Luis Aguirre, Narciso Martí-Oliet, Miguel Palomino, and Isabel Pita. Sentence-normalized conditional narrowing modulo in rewriting logic and Maude. Journal of Automated Reasoning, 60(4):421–463, April 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-017-9417-
5.

Lucas:2018:ASL

[1086] Salvador Lucas and Raúl Gutiérrez.

Automatic synthesis of logical models for order-sorted first-order theories. Journal of Automated Reasoning, 60(4):465–501, April 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9419-3.

Braude:2018:GMO

[1087] Eric Braude and Satbek Abdyldayev.

Generalizing Morley’s and other theorems with automated realization. Journal of Automated Reasoning, 60(4): 503–526, April 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9435-3.

Slaney:2018:ECR

[1088] John Slaney and Bruno Woltzenlogel Paleo. Erratum to: Conflict resolution: A first-order resolution calculus with decision literals and conflictdriven clause learning. Journal of Automated Reasoning, 60(4):527, April 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9434-4;   http://link.springer.
com/content/pdf/10.1007/s10817-
017-9434-4.pdf. See [1073].

Avigad:2018:IMI

[1089] Jeremy Avigad, Jasmin Christian Blanchette, Gerwin Klein, Lawrence Paulson, Andrei Popescu, and Gregor Snelting. Introduction to milestones in interactive theorem proving. Journal of Automated Reasoning, 61(1–4):1–8, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9465-5;   http:
//link.springer.com/content/pdf/
10.1007/s10817-018-9465-5.pdf.

Bancerek:2018:RMM

[1090] Grzegorz Bancerek, Czesław Byliński, Adam Grabowski, Artur Korniłowicz, Roman Matuszewski, Adam Naumowicz, and Karol Pak. The role of the Mizar mathematical library for interactive proof development in Mizar. Journal of Automated Reasoning, 61 (1–4):9–32, June 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9440-6;   http:
//link.springer.com/content/pdf/
10.1007/s10817-017-9440-6.pdf.

Bertot:2018:DDP

[1091] Yves Bertot, Laurence Rideau, and Laurent Théry. Distant decimals of π: Formal proofs of some algorithms computing them and guarantees of exact computation. Journal of Automated Reasoning, 61(1–4):33–71, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9444-2.

Immler:2018:VOS

[1092] Fabian Immler. A verified ODE solver and the Lorenz attractor. Journal of Automated Reasoning, 61(1–4):73–111, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9448-y;   http://link.springer.
com/content/pdf/10.1007/s10817-
017-9448-y.pdf.

Bauereiss:2018:CCV

[1093] Thomas Bauereiß, Armando Pesenti Gritti, Andrei Popescu, and Franco Raimondi. CoSMed: A confidentialityverified social media platform. Journal of Automated Reasoning, 61(1–4): 113–139, June 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9443-3.

Chen:2018:TCV

[1094] Hao Chen, Xiongnan Wu, Zhong Shao, Joshua Lockerman, and Ronghui Gu. Toward compositional verification of interruptible OS kernels and device drivers. Journal of Automated Reasoning, 61(1–4):141–189, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9446-0.

Diekmann:2018:VIF

[1095] Cornelius Diekmann, Lars Hupel, Julius Michaelis, Maximilian Haslbeck, and Georg Carle. Verified iptables firewall analysis and verification. Journal of Automated Reasoning, 61(1–4): 191–242, June 2018. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9445-1;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-017-9445-1.pdf.

Lochbihler:2018:MTS

[1096] Andreas Lochbihler. Mechanising a type-safe model of multithreaded Java with a verified compiler. Journal of Automated Reasoning, 61(1–4):243–332, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9452-x.

Blanchette:2018:VSS

[1097] Jasmin Christian Blanchette, Mathias Fleury, Peter Lammich, and Christoph Weidenbach. A verified SAT solver framework with learn, forget, restart, and incrementality. Journal of Automated Reasoning, 61(1–4):333–365, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9455-7;   http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        018-9455-7.pdf.

Cao:2018:VFS

[1098] Qinxiang Cao, Lennart Beringer, Samuel Gruetter, Josiah Dodds, and Andrew W. Appel. VST-Floyd: A separation logic tool to verify correctness of C programs. Journal of Automated Reasoning, 61(1–4):367–422, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9457-5.

Czajka:2018:HCA

[1099] Łukasz Czajka and Cezary Kaliszyk.

Hammer for Coq: Automation for dependent type theory. Journal of Automated Reasoning, 61(1–4):423–453, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9458-4;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9458-4.pdf.

Schlichtkrull:2018:FRC

[1100] Anders Schlichtkrull. Formalization of the resolution calculus for first-order logic. Journal of Automated Reasoning, 61(1–4):455–484, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9447-z.

Abdulaziz:2018:FVA

[1101] Mohammad Abdulaziz, Michael Norrish, and Charles Gretton. Formally verified algorithms for upper-bounding state space diameters. Journal of Automated Reasoning, 61(1–4):485–520, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9450-z.

Doczkal:2018:RLR

[1102] Christian Doczkal and Gert Smolka. Regular language representations in the constructive type theory of Coq. Journal of Automated Reasoning, 61 (1–4):521–553, June 2018. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-018-9460-x.

Boutry:2019:PPC

[1103] Pierre Boutry, Charly Gries, Julien Narboux, and Pascal Schreck. Parallel postulates and continuity axioms: A mechanized study in intuitionistic logic using Coq. Journal of Automated Reasoning, 62(1):1–68, January 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-017-
                                        9422-8.

Li:2019:DUP

[1104] Wenda Li, Grant Olney Passmore, and Lawrence C. Paulson. Deciding univariate polynomial problems using untrusted certificates in Isabelle/ HOL. Journal of Automated Reasoning, 62 (1):69–91, January 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-017-9424-6;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-017-9424-6.pdf.

deGouw:2019:VOS

[1105] Stijn de Gouw, Frank S. de Boer, and Richard Bubel. Verifying OpenJDK’s sort method for generic collections. Journal of Automated Reasoning, 62 (1):93–126, January 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9426-4;   http:
//link.springer.com/content/pdf/
10.1007/s10817-017-9426-4.pdf.

Ferrari:2019:GOP

[1106] Mauro Ferrari and Camillo Fiorentini. Goal-oriented proof-search in natural deduction for intuitionistic propositional logic. Journal of Automated Reasoning, 62(1):127–167, January 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9427-3.

Blanchette:2019:SEP

[1107] Jasmin Christian Blanchette and Stephan Merz. Selected extended papers of ITP 2016: Preface. Journal of Automated Reasoning, 62(2): 169–170, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9470-8;   http:
//link.springer.com/content/pdf/
10.1007/s10817-018-9470-8.pdf.

Chan:2019:PPB

[1108] Hing-Lun Chan and Michael Norrish.

Proof pearl: Bounding least common multiples with triangles. Journal of Automated Reasoning, 62(2): 171–192, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9438-0.

Gregoire:2019:MAF

[1109] Thomas Grégoire and Adam Chlipala. Mostly automated formal verification of loop dependencies with applications to distributed stencil algorithms. Journal of Automated Reasoning, 62(2): 193–213, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-018-9451-y.

Immler:2019:FOF

[1110] Fabian Immler and Christoph Traut. The flow of ODEs: Formalization of variational equation and Poincaré map. Journal of Automated Reasoning, 62 (2):215–236, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-018-9449-5.

Kuncar:2019:TSL

[1111] Ondřej Kunčar and Andrei Popescu. From types to sets by local type definition in higher-order logic. Journal of Automated Reasoning, 62(2): 237–260, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-018-9464-6.

Lammich:2019:FNF

[1112] Peter Lammich and S. Reza Sefidgar. Formalizing network flow algorithms: a refinement approach in Isabelle/HOL. Journal of Automated Reasoning, 62 (2):261–280, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9442-4.

Mahboubi:2019:FVA

[1113] Assia Mahboubi, Guillaume Melquiond, and Thomas Sibut-Pinote. Formally verified approximations of definite integrals. Journal of Automated Reasoning, 62(2):281–300, February 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9463-7.

Maric:2019:FFP

[1114] Filip Marić. Fast formal proof of the Erdős–Szekeres conjecture for convex polygons with at most 6 points. Journal of Automated Reasoning, 62 (3):301–329, March 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9423-7.

Chargueraud:2019:VCA

[1115] Arthur Charguéraud and François Pottier. Verifying the correctness and amortized complexity of a union– find implementation in separation logic with time credits. Journal of Automated Reasoning, 62(3):331–365, March 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9431-7.

Nipkow:2019:ACV

[1116] Tobias Nipkow and Hauke Brinkop.

Amortized complexity verified. Journal of Automated Reasoning, 62(3):

367–391, March 2019. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9459-3.

Leitsch:2019:EET

[1117] Alexander Leitsch and Anela Lolic. Extraction of expansion trees. Journal of Automated Reasoning, 62(3):393–430, March 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9453-9;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9453-9.pdf.

Zhang:2019:SEP

[1118] Xingyuan Zhang and Christian Urban. Selected extended papers of ITP 2015: Preface. Journal of Automated Reasoning, 62(4):431–432, April 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09517-z;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09517-z.pdf.

Besson:2019:VCF

[1119] Frédéric Besson, Sandrine Blazy, and Pierre Wilke. A verified CompCert front-end for a memory model supporting pointer arithmetic and uninitialised data. Journal of Automated Reasoning, 62(4):433–480, April 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9439-z.

Lammich:2019:RIH

[1120] Peter Lammich. Refinement to Imperative HOL. Journal of Automated Reasoning, 62(4):481–503, April 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-017-
9437-1.

Boulme:2019:RCA

[1121] Sylvain Boulmé and Alexandre Maréchal.

Refinement to certify abstract interpretations: Illustrated on linearization for polyhedra. Journal of Automated Reasoning, 62(4):505–530, April 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9492-2.

Kuncar:2019:CFI

[1122] Ondřej Kunčar and Andrei Popescu.

A consistent foundation for Isabelle/ HOL. Journal of Automated Reasoning, 62(4):531–555, April 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL http://link.springer.com/
article/10.1007/s10817-018-9454-
8.

Barthe:2019:SLN

[1123] Gilles Barthe, Gustavo Betarte, Juan Diego Campo, and Carlos Luna. System-level non-interference of constant-time cryptography. Part I: Model. Journal of Automated Reasoning, 63(1):1–51, June 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-017-9441-5.

Lammich:2019:ARE

[1124] Peter Lammich and Andreas Lochbihler. Automatic refinement to efficient data structures: a comparison of two approaches. Journal of Automated Reasoning, 63(1):53–94, June 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9461-9.

Ebner:2019:GQL

[1125] Gabriel Ebner, Stefan Hetzl, Alexander Leitsch, Giselle Reis, and Daniel Weller. On the generation of quantified lemmas. Journal of Automated Reasoning, 63(1):95–126, June 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9462-8;   http:
//link.springer.com/content/pdf/
10.1007/s10817-018-9462-8.pdf.

Peitl:2019:LDR

[1126] Tomáš Peitl, Friedrich Slivovsky, and Stefan Szeider. Long-distance Qresolution with dependency schemes. Journal of Automated Reasoning, 63 (1):127–155, June 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9467-3;   http:
//link.springer.com/content/pdf/
10.1007/s10817-018-9467-3.pdf.

Lumsdaine:2019:PSI

[1127] Peter Lefanu Lumsdaine and Nicolas Tabareau. Preface: Special issue on homotopy type theory and univalent foundations. Journal of Automated Reasoning, 63(2):157–158, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9491-3;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9491-3.pdf.

Bezem:2019:UAC

[1128] Marc Bezem, Thierry Coquand, and Simon Huber. The univalence axiom in cubical sets. Journal of Automated Reasoning, 63(2):159–171, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9472-6;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9472-6.pdf.

Huber:2019:CCT

[1129] Simon Huber. Canonicity for cubical type theory. Journal of Automated Reasoning, 63(2):173–210, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9469-1;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9469-1.pdf.

Birkedal:2019:GCT

[1130] Lars Birkedal, Aleš Bizjak, Ranald Clouston, Hans Bugge Grathwohl, Bas Spitters, and Andrea Vezzosi. Guarded cubical type theory. Journal of Automated Reasoning, 63(2):211–253, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9471-7.

Brunerie:2019:JCP

[1131] Guillaume Brunerie. The James construction and π_{4}(S^{3}) in homotopy type theory. Journal of Automated Reasoning, 63(2):255–284, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9468-2.

Ahrens:2019:SMU

[1132] Benedikt Ahrens, Ralph Matthes, and Anders Mörtberg. From signatures to monads in UniMath. Journal of Automated Reasoning, 63(2):285–318, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9474-4;   http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        018-9474-4.pdf.

Ayala-Rincon:2019:SEP

[1133] Mauricio Ayala-Rincón and César Muñoz. Selected extended papers of ITP 2017. Journal of Automated Reasoning, 63(2):319–321, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        09500-0;   http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        018-09500-0.pdf.

Allamigeon:2019:FCP

[1134] Xavier Allamigeon and Ricardo D.

Katz. A formalization of convex polyhedra based on the simplex method. Journal of Automated Reasoning, 63 (2):323–345, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9477-1.

Bentkamp:2019:FPE

[1135] Alexander Bentkamp, Jasmin Christian Blanchette, and Dietrich Klakow. A formal proof of the expressiveness of deep learning. Journal of Automated Reasoning, 63(2):347–368, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9481-5;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9481-5.pdf.

Besson:2019:CMA

[1136] Frédéric Besson, Sandrine Blazy, and Pierre Wilke. CompCertS: a memoryaware verified C compiler using a pointer as integer semantics. Journal of Automated Reasoning, 63(2): 369–392, August 2019. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9496-y.

Forster:2019:CVL

[1137] Yannick Forster and Gert Smolka. Callby-value Lambda calculus as a model of computation in Coq. Journal of Automated Reasoning, 63(2):393–413, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9484-2.

Kirst:2019:CRL

[1138] Dominik Kirst and Gert Smolka. Categoricity results and large model constructions for second-order ZF in dependent type theory. Journal of Automated Reasoning, 63(2):415–438, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9480-6.

Lochbihler:2019:EPH

[1139] Andreas Lochbihler. Effect polymorphism in higher-order logic (proof pearl). Journal of Automated Reasoning, 63(2):439–462, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s10817-018-9476-
                                        2.

Ericsson:2019:VGG

[1140] Adam Sandberg Ericsson, Magnus O. Myreen, and Johannes Åman Pohjola. A verified generational garbage collector for CakeML. Journal of Automated Reasoning, 63(2):463–488, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9487-z;   http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        018-9487-z.pdf.

Zakowski:2019:VCG

[1141] Yannick Zakowski, David Cachera, Delphine Demange, Gustavo Petri, David Pichardie, Suresh Jagannathan, and Jan Vitek. Verifying a concurrent garbage collector with a relyguarantee methodology. Journal of Automated Reasoning, 63(2):489–515, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9489-x.

Zhan:2019:FFG

[1142] Bohua Zhan. Formalization of the fundamental group in untyped set theory using Auto2. Journal of Automated Reasoning, 63(2):517–538, August 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9478-0.

Tao:2019:PAA

[1143] Jia Tao. A PSpace algorithm for acyclic epistemic DL(ALCS)▽_{⇕}. Journal of Automated Reasoning, 63(3): 539–555, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9473-5.

Kaliszyk:2019:SMI

[1144] Cezary Kaliszyk and Karol Pak. Semantics of Mizar as an Isabelle object logic. Journal of Automated Reasoning, 63(3):557–595, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.

                                        1007/s10817-018-9479-z;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-018-9479-z.pdf.

Beyersdorff:2019:RDS

[1145] Olaf Beyersdorff, Joshua Blinkhorn, Leroy Chew, Renate Schmidt, and Martin Suda. Reinterpreting dependency schemes: Soundness meets incompleteness in DQBF. Journal of Automated Reasoning, 63(3):597–623, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9482-4;   http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        018-9482-4.pdf.

Miller:2019:MMR

[1146] Dale Miller. Mechanized metatheory revisited. Journal of Automated Reasoning, 63(3):625–665, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9483-3.

Chan:2019:CFF

[1147] Hing-Lun Chan and Michael Norrish. Classification of finite fields with applications. Journal of Automated Reasoning, 63(3):667–693, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-018-
                                        9485-1.

Cruz-Filipe:2019:FVS

[1148] Luís Cruz-Filipe, Joao Marques-Silva, and Peter Schneider-Kamp. Formally verifying the solution to the Boolean Pythagorean Triples problem. Journal of Automated Reasoning, 63(3): 695–722, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9490-4.

Wang:2019:WPB

[1149] Wenxi Wang, Harald Søndergaard, and Peter J. Stuckey. Wombit: a portfolio bit-vector solver using wordlevel propagation. Journal of Automated Reasoning, 63(3):723–762, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9493-1.

Abdulaziz:2019:IHF

[1150] Mohammad Abdulaziz and Lawrence C.

Paulson. An Isabelle/HOL formalisation of Green’s Theorem. Journal of Automated Reasoning, 63(3): 763–786, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9495-z.

Li:2019:FGA

[1151] Li-Ming Li, Zhi-Ping Shi, Yong Guan, Qian-Ying Zhang, and Yong-Dong Li. Formalization of geometric algebra in HOL Light. Journal of Automated Reasoning, 63(3):787–808, October 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9498-9.

Cervesato:2019:PSI

[1152] Iliano Cervesato and Maribel Fernández. Preface to the special issue on linearity. Journal of Automated Reasoning, 63 (4):809–811, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-019-09529-9;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-019-09529-9.pdf.

Baillot:2019:ICC

[1153] Patrick Baillot, Gilles Barthe, and Ugo Dal Lago. Implicit computational complexity of subrecursive definitions and applications to cryptographic proofs. Journal of Automated Reasoning, 63 (4):813–855, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-019-09530-2.

Heath:2019:PTM

[1154] Quentin Heath and Dale Miller. A proof theory for model checking. Journal of Automated Reasoning, 63(4): 857–885, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-018-9475-3.

Mackie:2019:LNS

[1155] Ian Mackie. Linear numeral systems. Journal of Automated Reasoning, 63 (4):887–909, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-018-9456-6;   http:

//link.springer.com/content/pdf/
10.1007/s10817-018-9456-6.pdf.

Acclavio:2019:PDM

[1156] Matteo Acclavio. Proof diagrams for multiplicative linear logic: Syntax and semantics. Journal of Automated Reasoning, 63(4):911–939, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9466-4.

Paolini:2019:QHO

[1157] Luca Paolini, Mauro Piccolo, and Margherita Zorzi. QPCF: Higher-order languages and quantum circuits. Journal of Automated Reasoning, 63(4): 941–966, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09518-y.

Mahmoud:2019:FMQ

[1158] Mohamed Yousri Mahmoud and Amy P. Felty. Formalization of metatheory of the Quipper quantum programming language in a linear logic. Journal of Automated Reasoning, 63 (4):967–1002, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09527-x.

Barrett:2019:SEP

[1159] Clark Barrett and Temesghen Kahsai. Selected extended papers of NFM 2017: Preface. Journal of Automated Reasoning, 63(4):1003–1004, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9488-y;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-9488-y.pdf.

Sogokon:2019:VSP

[1160] Andrew Sogokon, Paul B. Jackson, and Taylor T. Johnson. Verifying safety and persistence in hybrid systems using flowpipes and continuous invariants. Journal of Automated Reasoning, 63(4):1005–1029, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9497-x.

Dreossi:2019:CFC

[1161] Tommaso Dreossi, Alexandre Donzé, and Sanjit A. Seshia. Compositional falsification of cyber-physical systems with machine learning components. Journal of Automated Reasoning, 63(4):1031–1053, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09509-5.

Jha:2019:EAD

[1162] Susmit Jha, Tuhin Sahai, Vasumathi Raman, Alessandro Pinto, and Michael Francis. Explaining AI decisions using efficient methods for learning sparse Boolean formulae. Journal of Automated Reasoning, 63(4):1055–1075, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s10817-018-
9499-8.

Frenkel:2019:ATA

[1163] Hadar Frenkel, Orna Grumberg, and Sarai Sheinvald. An automatatheoretic approach to model-checking systems and specifications over infinite data domains. Journal of Automated Reasoning, 63(4):1077–1101, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
9494-0.

Aichernig:2019:EAA

[1164] Bernhard K. Aichernig and Martin Tappler. Efficient active automata learning via mutation testing. Journal of Automated Reasoning, 63(4): 1103–1134, December 2019. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-9486-0;   http:
//link.springer.com/content/pdf/
10.1007/s10817-018-9486-0.pdf.

Huang:2020:CDS

[1165] Cheng-Chao Huang, Ming Xu, and Zhi-Bin Li. A conflict-driven solving procedure for poly-power constraints. Journal of Automated Reasoning, 64 (1):1–20, January 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-09501-z.

Nguyen:2020:ETG

[1166] Linh Anh Nguyen. ExpTime tableaux with global caching for hybrid PDL. Journal of Automated Reasoning, 64 (1):21–52, January 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-09506-8.

Benzmuller:2020:AFL

[1167] Christoph Benzmüller and Dana S. Scott. Automating free logic in HOL, with an experimental application in category theory. Journal of Automated Reasoning, 64(1):53–72, January 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09507-7.

Zhang:2020:PIP

[1168] Xingyuan Zhang, Christian Urban, and Chunhan Wu. Priority inheritance protocol proved correct. Journal of Automated Reasoning, 64(1):73–95, January 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09511-5;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09511-5.pdf.

Chocron:2020:PCM

[1169] Paula Chocron, Pascal Fontaine, and Christophe Ringeissen. Politeness and combination methods for theories with bridging functions. Journal of Automated Reasoning, 64(1):97–134, January 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09512-4.

Mayer:2020:PDN

[1170] Marta Cialdea Mayer. A prover dealing with nominals, binders, transitivity and relation hierarchies. Journal of Automated Reasoning, 64(1): 135–165, January 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09513-3.

Lucas:2020:UWF

[1171] Salvador Lucas. Using well-founded relations for proving operational termination. Journal of Automated Reasoning, 64(2):167–195, February 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09514-2.

Baumgartner:2020:BOE

[1172] Peter Baumgartner and Renate A.

Schmidt. Blocking and other enhancements for bottom-up model generation methods. Journal of Automated Reasoning, 64(2):197–251, February 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09515-1;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09515-1.pdf.

Echenim:2020:CIS

[1173] M. Echenim and N. Peltier. Combining induction and saturation-based theorem proving. Journal of Automated Reasoning, 64(2):253–294, February 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09519-x.

Cristia:2020:SQF

[1174] Maximiliano Cristiá and Gianfranco Rossi. Solving quantifier-free firstorder constraints over finite sets and binary relations. Journal of Automated Reasoning, 64(2):295–330, February 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09520-4.

Li:2020:EWN

[1175] Wenda Li and Lawrence C. Paulson. Evaluating winding numbers and counting complex roots through Cauchy indices in Isabelle/HOL. Journal of Automated Reasoning, 64(2): 331–360, February 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09521-3;   http:
//link.springer.com/content/pdf/
10.1007/s10817-019-09521-3.pdf.

Biere:2020:PSI

[1176] Armin Biere, Cesare Tinelli, and Christoph Weidenbach. Preface to the special issue on automated reasoning systems. Journal of Automated Reasoning, 64(3):361–362, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09531-1;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09531-1.pdf.

Joosten:2020:VIA

[1177] Sebastiaan J. C. Joosten, René Thiemann, and Akihisa Yamada. A verified implementation of algebraic numbers in Isabelle/HOL. Journal of Automated Reasoning, 64(3):363–389, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09504-w;   http://link.springer.
com/content/pdf/10.1007/s10817-
018-09504-w.pdf.

Kaufmann:2020:LSO

[1178] Matt Kaufmann and J. Strother Moore. Limited second–order functionality in a first–order setting. Journal of Automated Reasoning, 64(3):391–422, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09505-9.

Sebastiani:2020:OTO

[1179] Roberto Sebastiani and Patrick Trentin.

OptiMathSAT: A tool for optimization modulo theories. Journal of Automated Reasoning, 64(3):423–460, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09508-6.

Nalon:2020:RBT

[1180] Cláudia Nalon, Ullrich Hustadt, and Clare Dixon. A resolution–based theorem prover for K_{n}: Architecture, refinements, strategies and experiments. Journal of Automated Reasoning, 64 (3):461–484, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-018-09503-x;   http:
//link.springer.com/content/pdf/
10.1007/s10817-018-09503-x.pdf.

Barbosa:2020:SFG

[1181] Haniel Barbosa, Jasmin Christian Blanchette, Mathias Fleury, and Pascal Fontaine. Scalable fine–grained proofs for formula processing. Journal of Automated Reasoning, 64(3):485–510, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09502-y.

deMoura:2020:PSE

[1182] Leonardo de Moura. Preface: Selected extended papers of CADE 2017. Journal of Automated Reasoning, 64(3):511, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09547-y;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09547-y.pdf.

Lammich:2020:EVS

[1183] Peter Lammich. Efficient verified (UN)SAT certificate checking. Journal of Automated Reasoning, 64(3): 513–532, March 2020. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09525-z;   http:
//link.springer.com/content/pdf/
10.1007/s10817-019-09525-z.pdf.

Heule:2020:SEF

[1184] Marijn J. H. Heule, Benjamin Kiesl, and Armin Biere. Strong extension– free proof systems. Journal of Automated Reasoning, 64(3):533–554, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09516-0;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09516-0.pdf.

Tellez:2020:AVT

[1185] Gadi Tellez and James Brotherston. Automatically verifying temporal properties of pointer programs with cyclic proof. Journal of Automated Reasoning, 64(3):555–578, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09532-0;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09532-0.pdf.

Bonacina:2020:CDS

[1186] Maria Paola Bonacina, Stéphane Graham-Lengrand, and Natarajan Shankar. Conflict-driven satisfiability for theory combination: Transition system and completeness. Journal of Automated Reasoning, 64(3):579–609, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-018-
09510-y.

Teucke:2020:SAF

[1187] Andreas Teucke and Christoph Weidenbach. SPASS-AR: A first–order theorem prover based on approximation– refinement into the monadic shallow linear fragment. Journal of Automated Reasoning, 64(3):611–640, March 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09546-z.

Gheri:2020:FGT

[1188] Lorenzo Gheri and Andrei Popescu. A formalized general theory of syntax with bindings: Extended version. Journal of Automated Reasoning, 64 (4):641–675, April 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09522-2.

Gadgil:2020:HLF

[1189] Siddhartha Gadgil. Homogeneous length functions on groups: Intertwined computer and human proofs. Journal of Automated Reasoning, 64 (4):677–688, April 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09523-1.

Moghaddam:2020:ARP

[1190] G. I. Moghaddam, R. Padmanabhan, and Yang Zhang. Automated reasoning with power maps. Journal of Automated Reasoning, 64(4):689–697, April 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09524-0.

Divason:2020:VIB

[1191] Jose Divasón, Sebastiaan J. C. Joosten, René Thiemann, and Akihisa Yamada. A verified implementation of the Berlekamp–Zassenhaus factorization algorithm. Journal of Automated Reasoning, 64(4):699–735, April 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09526-y;   http://link.springer.
com/content/pdf/10.1007/s10817-
019-09526-y.pdf.

Echenim:2020:FCR

[1192] Mnacho Echenim, Hervé Guiol, and Nicolas Peltier. Formalizing the Cox– Ross–Rubinstein pricing of European derivatives in Isabelle/HOL. Journal of Automated Reasoning, 64(4):737–765, April 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09528-w.

Pena:2020:APR

[1193] Ricardo Peña. An assertional proof of red–black trees using Dafny. Journal of Automated Reasoning, 64(4):767–791, April 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09534-y.

Avigad:2020:PSE

[1194] Jeremy Avigad and Assia Mahboubi.

Preface: Selected extended papers from Interactive Theorem Proving 2018. Journal of Automated Reasoning, 64 (5):793–794, June 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09557-w;   http:
//link.springer.com/content/pdf/
10.1007/s10817-020-09557-w.pdf.

Doczkal:2020:GTC

[1195] Christian Doczkal and Damien Pous. Graph theory in Coq: Minors, treewidth, and isomorphisms. Journal of Automated Reasoning, 64(5): 795–825, June 2020. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09543-2.

Thiemann:2020:FLB

[1196] René Thiemann, Ralph Bottesch, Jose Divasón, Max W. Haslbeck, Sebastiaan J. C. Joosten, and Akihisa Yamada. Formalizing the LLL basis reduction algorithm and the LLL factorization algorithm in Isabelle/HOL. Journal of Automated Reasoning, 64 (5):827–856, June 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09552-1;   http:
//link.springer.com/content/pdf/
10.1007/s10817-020-09552-1.pdf.

Daggitt:2020:RUD

[1197] Matthew L. Daggitt, Ran Zmigrod, and Timothy G. Griffin. A relaxation of Üresin and Dubois’ asynchronous fixed–point theory in Agda. Journal of Automated Reasoning, 64 (5):857–877, June 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09536-w;   http:
//link.springer.com/content/pdf/
10.1007/s10817-019-09536-w.pdf.

Eberl:2020:VAR

[1198] Manuel Eberl, Max W. Haslbeck, and Tobias Nipkow. Verified analysis of random binary tree structures. Journal of Automated Reasoning, 64(5):879–910, June 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09545-0;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09545-0.pdf.

Syeda:2020:FRU

[1199] Hira Taqdees Syeda and Gerwin Klein.

Formal reasoning under cached address translation. Journal of Automated Reasoning, 64(5):911–945, June 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09539-7.

Sozeau:2020:MP

[1200] Matthieu Sozeau, Abhishek Anand, Simon Boulier, Cyril Cohen, Yannick Forster, Fabian Kunze, Gregory Malecha, Nicolas Tabareau, and Théo Winterhalter. The MetaCoq project. Journal of Automated Reasoning, 64 (5):947–999, June 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-019-09540-0.

Burel:2020:FOA

[1201] Guillaume Burel, Guillaume Bury, Raphaël Cauderlier, David Delahaye, Pierre Halmagrand, and Olivier Hermant. First-order automated reasoning with theories: When deduction modulo theory meets practice. Journal of Automated Reasoning, 64(6): 1001–1050, August 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-019-09533-z.

Hajdu:2020:ESC

[1202] Ákos Hajdu and Zoltán Micskei. Efficient strategies for CEGAR-based model checking. Journal of Automated Reasoning, 64(6):1051–1091, August 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-019-
                                        09535-x;   http://link.springer.
                                        com/content/pdf/10.1007/s10817-
                                        019-09535-x.pdf.

Ballarin:2020:ESA

[1203] Clemens Ballarin. Exploring the structure of an algebra text with locales. Journal of Automated Reasoning, 64 (6):1093–1121, August 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-019-09537-9.

Affeldt:2020:LFL

[1204] Reynald Affeldt, Jacques Garrigue, and Takafumi Saikawa. A library for formalization of linear errorcorrecting codes. Journal of Automated Reasoning, 64(6):1123–1164, August 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-019-
09538-8.

Galmiche:2020:PSI

[1205] Didier Galmiche, Stephan Schulz, and Roberto Sebastiani. Preface: Special issue of selected extended papers from IJCAR 2018. Journal of Automated Reasoning, 64(7):1165–1167, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09556-x;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09556-x.pdf.

Schlichtkrull:2020:FBG

[1206] Anders Schlichtkrull, Jasmin Blanchette, Dmitriy Traytel, and Uwe Waldmann. Formalizing Bachmair and Ganzinger’s ordered resolution prover. Journal of Automated Reasoning, 64(7):1169– 1195, October 2020. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09561-0.

Larchey-Wendling:2020:CDR

[1207] Dominique Larchey-Wendling. Constructive decision via redundancy-free proof-search. Journal of Automated Reasoning, 64(7):1197–1219, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09555-y.

Das:2020:QMB

[1208] Anupam Das. From QBFs to MALL and back via focussing. Journal of Automated Reasoning, 64(7):1221– 1245, October 2020. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09564-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-020-09564-x.pdf.

Kiesl:2020:SSP

[1209] Benjamin Kiesl, Adrián Rebola-Pardo, Marijn J. H. Heule, and Armin Biere. Simulating strong practical proof systems with extended resolution. Journal of Automated Reasoning, 64(7): 1247–1267, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09554-z;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-020-09554-z.pdf.

Finger:2020:PPT

[1210] Marcelo Finger and Sandro Preto. Probably partially true: Satisfiability for Łukasiewicz infinitely valued probabilistic logic and related topics. Journal of Automated Reasoning, 64(7): 1269–1286, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09558-9.

Abrahamsson:2020:PPS

[1211] Oskar Abrahamsson, Son Ho, Hrutvik Kanabar, Ramana Kumar, Magnus O. Myreen, Michael Norrish, and Yong Kiam Tan. Proof-producing synthesis of CakeML from monadic HOL functions. Journal of Automated Reasoning, 64(7):1287–1306, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09559-8;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09559-8.pdf.

Conchon:2020:PMC

[1212] Sylvain Conchon, David Declerck, and Fatiha Zaïdi. Parameterized model checking on the TSO weak memory model. Journal of Automated Reasoning, 64(7):1307–1330, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09565-w.

Beyer:2020:SEP

[1213] Dirk Beyer and Marieke Huisman.

Selected and extended papers from TACAS 2018: Preface. Journal of Automated Reasoning, 64(7):1331– 1332, October 2020. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09575-8;   http:
//link.springer.com/content/pdf/
10.1007/s10817-020-09575-8.pdf.

Bansal:2020:SPU

[1214] Kshitij Bansal, Eric Koskinen, and Omer Tripp. Synthesizing precise and useful commutativity conditions. Journal of Automated Reasoning, 64(7): 1333–1359, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09573-w.

Bryant:2020:CRB

[1215] Randal E. Bryant. Chain reduction for binary and zero-suppressed decision diagrams. Journal of Automated Reasoning, 64(7):1361–1391, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10817-020-
                                        09569-6.

Champion:2020:IBR

[1216] Adrien Champion, Tomoya Chiba, Naoki Kobayashi, and Ryosuke Sato. ICE-based refinement type discovery for higher-order functional programs. Journal of Automated Reasoning, 64 (7):1393–1418, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09571-y.

Chini:2020:FGC

[1217] Peter Chini, Roland Meyer, and Prakash Saivasan. Fine-grained complexity of safety verification. Journal of Automated Reasoning, 64(7): 1419–1444, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09572-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10817-020-09572-x.pdf.

Costa:2020:NPP

[1218] Gabriele Costa, Letterio Galletta, Pierpaolo Degano, David Basin, and Chiara Bodei. Natural projection as partial model checking. Journal of Automated Reasoning, 64(7):1445–1481, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09568-7;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09568-7.pdf.

Hartmanns:2020:MCB

[1219] Arnd Hartmanns, Sebastian Junges, Joost-Pieter Katoen, and Tim Quatmann. Multi-cost bounded tradeoff analysis in MDP. Journal of Automated Reasoning, 64(7):1483–1522, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09574-9;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09574-9.pdf.

Neider:2020:LBA

[1220] Daniel Neider, P. Madhusudan, Shambwaditya Saha, Pranav Garg, and Daejun Park. A learning-based approach to synthesizing invariants for incomplete verification engines. Journal of Automated Reasoning, 64(7): 1523–1552, October 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09570-z;   http:
//link.springer.com/content/pdf/
10.1007/s10817-020-09570-z.pdf.

Hustadt:2020:TPP

[1221] Ullrich Hustadt, Ana Ozaki, and Clare Dixon. Theorem proving for pointwise metric temporal logic over the naturals via translations. Journal of Automated Reasoning, 64(8):1553–1610, December 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09541-4;   http://link.springer.
com/content/pdf/10.1007/s10817-
020-09541-4.pdf.

Lucas:2020:DPF

[1222] Salvador Lucas, José Meseguer, and Raúl Gutiérrez. The 2D dependency pair framework for conditional rewrite systems— Part II: Advanced processors and implementation techniques. Journal of Automated Reasoning, 64(8):1611–1662, December 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09542-3.

Alonderis:2020:LTS

[1223] R. Alonderis, R. Pliuškevičius, A. Pliuškevičienė, and H. Giedra. Loop-type sequent calculi for temporal logic. Journal of Automated Reasoning, 64(8): 1663–1684, December 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http:

//link.springer.com/article/10.
1007/s10817-020-09544-1.

Barthe:2020:SLN

[1224] Gilles Barthe, Gustavo Betarte, Juan Diego Campo, Carlos Luna, and David Pichardie. System-level noninterference of constant-time cryptography. Part II: Verified static analysis and stealth memory. Journal of Automated Reasoning, 64(8):1685–1729, December 2020. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10817-020-
09548-x.

Guan:2021:FEL

[1225] Yong Guan, Jingzhi Zhang, and Yongdong Li. Formalization of Euler– Lagrange equation set based on variational calculus in HOL light. Journal of Automated Reasoning, 65(1): 1–29, January 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09549-w.

Simic:2021:FPD

[1226] Danijela Simić, Filip Marić, and Pierre Boutry. Formalization of the Poincaré disc model of hyperbolic geometry. Journal of Automated Reasoning, 65 (1):31–73, January 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09551-2.

Ambal:2021:PC

[1227] Guillaume Ambal, Sergueï Lenglet, and Alan Schmitt. HOπ in Coq. Journal of Automated Reasoning, 65 (1):75–124, January 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09553-0.

Beyersdorff:2021:BSQ

[1228] Olaf Beyersdorff, Joshua Blinkhorn, and Meena Mahajan. Building strategies into QBF proofs. Journal of Automated Reasoning, 65(1):125– 154, January 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09560-1.

Blanchette:2021:MNE

[1229] Jasmin Blanchette. Message from the new Editor-in-Chief. Journal of Automated Reasoning, 65(2): 155, February 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-021-09587-y.

Cattaruzza:2021:UTS

[1230] Dario Cattaruzza, Alessandro Abate, and Daniel Kroening. Unboundedtime safety verification of guarded LTI models with inputs by abstract acceleration. Journal of Automated Reasoning, 65(2):157– 203, February 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09562-z.

Chan:2021:MAA

[1231] Hing Lun Chan and Michael Norrish. Mechanisation of the AKS algorithm. Journal of Automated Reasoning, 65 (2):205–256, February 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09563-y.

Gauthier:2021:TLP

[1232] Thibault Gauthier, Cezary Kaliszyk, and Michael Norrish. TacticToe:

Learning to prove with tactics. Journal of Automated Reasoning, 65(2): 257–286, February 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09580-x.

Farber:2021:MLG

[1233] Michael Färber, Cezary Kaliszyk, and Josef Urban. Machine learning guidance for connection tableaux. Journal of Automated Reasoning, 65(2): 287–320, February 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09576-7.

Popescu:2021:CCM

[1234] Andrei Popescu, Peter Lammich, and Ping Hou. CoCon: A conference management system with formally verified document confidentiality. Journal of Automated Reasoning, 65(2): 321–356, February 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09566-9.

Voigt:2021:DEF

[1235] Marco Voigt. Decidable ∃^{∗}∀^{∗} firstorder fragments of linear rational arithmetic with uninterpreted predicates. Journal of Automated Reasoning, 65 (3):357–423, March 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09567-8.

Ghilardi:2021:HOQ

[1236] Silvio Ghilardi and Elena Pagani. Higher-order quantifier elimination, counter simulations and fault-tolerant systems. Journal of Automated Reasoning, 65(3):425–460, March 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10817-020-09578-
                                        5.

Beyer:2021:CUV

[1237] Dirk Beyer, Matthias Dangl, and Philipp Wendler. Correction to: A Unifying View on SMT-Based Software Verification. Journal of Automated Reasoning, 65(3):461, March 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10817-020-09585-

6. See [1081].

Cristia:2021:APB

[1238] Maximiliano Cristiá and Gianfranco Rossi. Automated proof of Bell– LaPadula security properties. Journal of Automated Reasoning, 65(4): 463–478, April 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09577-6.

Cortier:2021:DCS

[1239] Véronique Cortier, Stéphanie Delaune, and Vaishnavi Sundararajan. A decidable class of security protocols for both reachability and equivalence properties. Journal of Automated Reasoning, 65(4):479–520, April 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09582-9.

Butler:2021:FVP

[1240] D. Butler, A. Lochbihler, and A. Gascón. Formalising Σ-protocols and commitment schemes using CryptHOL. Journal of Automated Reasoning, 65(4):521–567, April 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-020-09581-
w.

Hou:2021:IHF

[1241] Zhé Hóu, David Sanan, and Jin Song Dong. An Isabelle/HOL formalisation of the SPARC instruction set architecture and the TSO memory model. Journal of Automated Reasoning, 65 (4):569–598, April 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09579-4.

Cerna:2021:SRF

[1242] David M. Cerna, Alexander Leitsch, and Anela Lolic. Schematic refutations of formula schemata. Journal of Automated Reasoning, 65(5): 599–645, June 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-020-09583-8.

Wernhard:2021:CIC

[1243] Christoph Wernhard. Craig interpolation with clausal first-order tableaux. Journal of Automated Reasoning, 65 (5):647–690, June 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-021-09590-3.

Bordg:2021:CQC

[1244] Anthony Bordg, Hanna Lachnitt, and Yijun He. Certified quantum computation in Isabelle/HOL. Journal of Automated Reasoning, 65(5): 691–709, June 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-020-09584-7.

Peng:2021:ADG

[1245] Xicheng Peng, Qihang Chen, and Mao Chen. Automated discovery of geometric theorems based on vector equations. Journal of Automated Reasoning, 65 (6):711–726, August 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-021-09591-2.

Vallati:2021:IDM

[1246] Mauro Vallati, Lukáš Chrpa, and Frank Hutter. On the importance of domain model configuration for automated planning engines. Journal of Automated Reasoning, 65(6): 727–773, August 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-021-09592-1.

Steen:2021:EHO

[1247] Alexander Steen and Christoph Benzmüller. Extensional higher-order paramodulation in Leo-III. Journal of Automated Reasoning, 65(6):775– 807, August 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09588-x.

Cristia:2021:ARR

[1248] Maximiliano Cristiá and Gianfranco Rossi. Automated reasoning with restricted intensional sets. Journal of Automated Reasoning, 65(6):809– 890, August 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09589-w.

Fontaine:2021:PSI

[1249] Pascal Fontaine. Preface: Special issue of selected extended papers of CADE 2019. Journal of Automated Reasoning, 65(7):891–892, October 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09601-3.

Bentkamp:2021:SL

[1250] Alexander Bentkamp, Jasmin Blanchette, and Uwe Waldmann. Superposition with lambdas. Journal of Automated Reasoning, 65(7):893– 940, October 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09595-y.

Calvanese:2021:MCU

[1251] Diego Calvanese, Silvio Ghilardi, and Andrey Rivkin. Model completeness, uniform interpolants and superposition calculus. Journal of Automated Reasoning, 65(7):941– 969, October 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09596-x.

Havlena:2021:ATL

[1252] Vojtěch Havlena, Lukáš Holík, and Tomáš Vojnar. Automata terms in a lazy WSkS decision procedure. Journal of Automated Reasoning, 65(7): 971–999, October 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09597-w.

Niemetz:2021:TSM

[1253] Aina Niemetz, Mathias Preiner, and Cesare Tinelli. Towards satisfiability modulo parametric bit-vectors. Journal of Automated Reasoning, 65(7): 1001–1025, October 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09598-9.

Popescu:2021:DRG

[1254] Andrei Popescu and Dmitriy Traytel. Distilling the Requirements of Gödel’s Incompleteness Theorems with a Proof Assistant. Journal of Automated Reasoning, 65(7):1027– 1070, October 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09599-8.

Trentin:2021:OMT

[1255] Patrick Trentin and Roberto Sebastiani. Optimization modulo the theories of signed bit-vectors and floating-point numbers. Journal of Automated Reasoning, 65(7):1071– 1096, October 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09600-4.

Claessen:2021:HTR

[1256] Koen Claessen and Ann Lillieström.

Handling transitive relations in firstorder automated reasoning. Journal of Automated Reasoning, 65(8): 1097–1124, December 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09605-z.

Cristia:2021:AVP

[1257] Maximiliano Cristiá and Gianfranco Rossi. An automatically verified prototype of the tokeneer ID station specification. Journal of Automated Reasoning, 65(8):1125–1151, December 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09602-2.

Sieg:2021:HCA

[1258] Wilfried Sieg and Farzaneh Derakhshan. Human-centered automated proof search. Journal of Automated Reasoning, 65(8):1153– 1190, December 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09594-z.

Lucas:2021:DCC

[1259] Salvador Lucas. Derivational complexity and context-sensitive rewriting. Journal of Automated Reasoning, 65(8):1191–1229, December 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-021-09603-
1.

deLima:2021:FRT

[1260] Thaynara Arielly de Lima, André Luiz Galdino, and Mauricio Ayala-Rincón. Formalization of ring theory in PVS. Journal of Automated Reasoning, 65(8):1231–1263, December 2021. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-021-09593-
0.

Kohlhase:2021:EEM

[1261] Michael Kohlhase and Florian Rabe. Experiences from exporting major proof assistant libraries. Journal of Automated Reasoning, 65(8):1265– 1298, December 2021. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09604-0.

Sigley:2022:PCM

[1262] Sarah Sigley and Olaf Beyersdorff. Proof complexity of modal resolution. Journal of Automated Reasoning, 66 (1):1–41, February 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09609-9.

Bonacina:2022:CDS

[1263] Maria Paola Bonacina, Stéphane Graham-Lengrand, and Natarajan Shankar. Conflict-driven satisfiability for theory combination: Lemmas, modules, and proofs. Journal of Automated Reasoning, 66(1): 43–91, February 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09606-y.

HuertayMunive:2022:PTS

[1264] Jonathan Julián Huerta y Munive and Georg Struth. Predicate transformer semantics for hybrid systems. Journal of Automated Reasoning, 66 (1):93–139, February 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09607-x.

Dalvandi:2022:IOG

[1265] Sadegh Dalvandi, Brijesh Dongol, and Heike Wehrheim. Integrating Owicki–Gries for C11-style memory models into Isabelle/HOL. Journal of Automated Reasoning, 66(1): 141–171, February 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09610-2.

Platzer:2022:CDD

[1266] André Platzer. Correction to: Differential Dynamic Logic for Hybrid Systems. Journal of Automated Reasoning, 66(1):173, February 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09608-w. See [768].

Boldo:2022:CFL

[1267] Sylvie Boldo, François Clément, and Micaela Mayero. A Coq formalization of Lebesgue integration of nonnegative functions. Journal of Automated Reasoning, 66(2): 175–213, May 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09612-0.

Lewis:2022:BDE

[1268] Robert Y. Lewis and Minchao Wu. A bi-directional extensible interface between Lean and Mathematica. Journal of Automated Reasoning, 66(2): 215–238, May 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09611-1.

Wojciechowski:2022:ARO

[1269] Piotr Wojciechowski, K. Subramani, and R. Chandrasekaran. Analyzing read-once cutting plane proofs in Horn systems. Journal of Automated Reasoning, 66(2):239–274, May 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09618-2.

Peltier:2022:SIS

[1270] Nicolas Peltier and Viorica Sofronie-Stokkermans. Special issue of selected extended papers of IJCAR 2020. Journal of Automated Reasoning, 66 (3):275–276, August 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09622-6.

Brakensiek:2022:RKC

[1271] Joshua Brakensiek, Marijn Heule, and David Narváez. The resolution of Keller’s Conjecture. Journal of Automated Reasoning, 66(3): 277–300, August 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09623-5.

Baader:2022:DWP

[1272] Franz Baader and Deepak Kapur.

Deciding the word problem for ground and strongly shallow identities w.r.t. extensional symbols. Journal of Automated Reasoning, 66(3): 301–329, August 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09624-4.

Sheng:2022:PCA

[1273] Ying Sheng, Yoni Zohar, and Clark Barrett. Polite combination of algebraic datatypes. Journal of Automated Reasoning, 66(3):331– 355, August 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09625-3.

Baader:2022:UMT

[1274] Franz Baader and Jakub Rydval. Using model theory to find decidable and tractable description logics with concrete domains. Journal of Automated Reasoning, 66(3): 357–407, August 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09626-2.

Calvanese:2022:CUI

[1275] Diego Calvanese, Silvio Ghilardi, and Andrey Rivkin. Combination of uniform interpolants via Beth definability. Journal of Automated Reasoning, 66(3):409–435, August 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09627-1.

Bonacina:2022:SDA

[1276] Maria Paola Bonacina. Six decades of automated reasoning: Papers in memory of Larry Wos. Journal of Automated Reasoning, 66(4):437– 438, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09637-z.

Beeson:2022:LWV

[1277] Michael Beeson, Maria Paola Bonacina, Michael Kinyon, and Geoff Sutcliffe. Larry Wos: Visions of automated reasoning. Journal of Automated Reasoning, 66(4):439–461, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09620-8.

Bonacina:2022:SSD

[1278] Maria Paola Bonacina. Set of support, demodulation, paramodulation: a historical perspective. Journal of Automated Reasoning, 66(4):463– 497, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09628-0.

Waldmann:2022:CFS

[1279] Uwe Waldmann, Sophie Tourret, Simon Robillard, and Jasmin Blanchette. A comprehensive framework for saturation theorem proving. Journal of Automated Reasoning, 66(4):499– 539, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09621-7.

Vukmirovic:2022:MHO

[1280] Petar Vukmirović, Alexander Bentkamp, Jasmin Blanchette, Simon Cruanes, Visa Nummelin, and Sophie Tourret. Making higher-order superposition work. Journal of Automated Reasoning, 66(4):541–564, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09613-z.

Veroff:2022:WCM

[1281] Robert Veroff. A Wos challenge met.

Journal of Automated Reasoning, 66 (4):565–574, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09614-y.

Tourret:2022:PCL

[1282] Sophie Tourret and Christoph Weidenbach. A posthumous contribution by Larry Wos: Excerpts from an unpublished column. Journal of Automated Reasoning, 66(4):575– 584, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09617-3.

Lochbihler:2022:MPM

[1283] Andreas Lochbihler. A mechanized proof of the max–flow min–cut theorem for countable networks with applications to probability theory. Journal of Automated Reasoning, 66(4): 585–610, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09616-4.

Baanen:2022:FDD

[1284] Anne Baanen, Sander R. Dahmen, Ashvni Narayanan, and Filippo A. E. Nuccio Mortarino Majno di Capriglio. A formalization of Dedekind domains and class groups of global fields. Journal of Automated Reasoning, 66(4): 611–637, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09644-0.

Papacchini:2022:LBE

[1285] Fabio Papacchini, Cláudia Nalon, Ullrich Hustadt, and Clare Dixon. Local is best: Efficient reductions to modal logic K. Journal of Automated Reasoning, 66(4):639–666, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09630-6. See cor-

rection [1298].

Yamada:2022:TIT

[1286] Akihisa Yamada. Tuple interpretations for termination of term rewriting. Journal of Automated Reasoning, 66 (4):667–688, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09640-4.

Janicic:2022:TPC

[1287] Predrag Janičić and Julien Narboux.

Theorem proving as constraint solving with coherent logic. Journal of Automated Reasoning, 66(4):689– 746, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09629-z.

Pearce:2022:VWP

[1288] David J. Pearce, Mark Utting, and Lindsay Groves. Verifying Whiley programs with Boogie. Journal of Automated Reasoning, 66(4):747– 803, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09619-1.

Meyers:2022:FLK

[1289] Joshua Meyers, David I. Spivak, and Ryan Wisnesky. Fast left Kan extensions using the chase. Journal of Automated Reasoning, 66(4):805– 844, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09634-2.

deChampeaux:2022:FLU

[1290] Dennis de Champeaux. Faster linear unification algorithm. Journal of Automated Reasoning, 66(4):845– 860, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09635-1.

Macedo:2022:PTR

[1291] Nuno Macedo, Julien Brunel, David Chemouil, and Alcino Cunha. Pardinus: a temporal relational model finder. Journal of Automated Reasoning, 66(4):861–904, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-022-09642-
2.

Soldevila:2022:STS

[1292] Mallku Soldevila, Beta Ziliani, and Bruno Silvestre. From specification to testing: Semantics engineering for Lua 5.2. Journal of Automated Reasoning, 66(4):905–952, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09638-y.

Schmoetten:2022:TFS

[1293] Richard Schmoetten, Jake E. Palmer, and Jacques D. Fleuriot. Towards formalising Schutz’ axioms for Minkowski spacetime in Isabelle/HOL. Journal of Automated Reasoning, 66(4): 953–988, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09643-1. See cor-

rection [1307].

Ricciotti:2022:FSN

[1294] Wilmer Ricciotti and James Cheney. A formalization of SQL with nulls. Journal of Automated Reasoning, 66(4): 989–1030, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09632-4.

Ramos:2022:FCT

[1295] Thiago Mendonça Ferreira Ramos, Ariane Alves Almeida, and Mauricio Ayala-Rincón. Formalization of the computational theory of a Turing complete functional language model. Journal of Automated Reasoning, 66(4): 1031–1063, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-021-09615-x.

Divason:2022:FSN

[1296] Jose Divasón and René Thiemann.

A formalization of the Smith normal form in higher–order logic. Journal of Automated Reasoning, 66(4): 1065–1095, November 2022. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09631-5. See cor-

rection [1297].

Divason:2022:CFS

[1297] Jose Divasón and René Thiemann. Correction to: A Formalization of the Smith Normal Form in Higher–Order Logic. Journal of Automated Reasoning, 66(4):1097, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09636-0.  See

[1296].

Papacchini:2022:CLB

[1298] Fabio Papacchini, Cláudia Nalon, Ullrich Hustadt, and Clare Dixon. Correction to: Local is Best: Efficient Reductions to Modal Logic K. Journal of Automated Reasoning, 66(4): 1099, November 2022. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09633-3.  See

[1285].

Rosskopf:2023:FPC

[1299] Simon Roßkopf and Tobias Nipkow. A formalization and proof checker for Isabelle’s metalogic. Journal of Automated Reasoning, 67(1): ??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09648-w.

Edmonds:2023:FSR

[1300] Chelsea Edmonds, Angeliki Koutsoukou-Argyraki, and Lawrence C. Paulson. Formalising Szemerédi’s regularity lemma and Roth’s theorem on arithmetic progressions in Isabelle/ HOL. Journal of Automated Reasoning, 67(1):??, March 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09650-2.

Fiorino:2023:LDD

[1301] Guido Fiorino. Linear depth deduction with subformula property for intuitionistic epistemic logic. Journal of Automated Reasoning, 67(1): ??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09653-z.

Wang:2023:SAC

[1302] Qinshi Wang and Andrew W. Appel.

A solver for arrays with concatenation. Journal of Automated Reasoning, 67(1):??, March 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09654-y.

Quaresma:2023:MRG

[1303] Pedro Quaresma and Pierluigi Graziani.

Measuring the readability of geometric proofs: The area method case. Journal of Automated Reasoning, 67 (1):??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09652-0.

Bonacina:2023:SGG

[1304] Maria Paola Bonacina and Sarah Winkler. Semantically-guided goalsensitive reasoning: Decision procedures and the Koala Prover. Journal of Automated Reasoning, 67(1): ??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09656-w.

Velasquez:2023:ODC

[1305] Alvaro Velasquez, Ismail Alkhouri, K. Subramani, Piotr Wojciechowski, and George Atia. Optimal deterministic controller synthesis from steady-state distributions. Journal of Automated Reasoning, 67(1): ??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09657-9.

Appel:2023:EEB

[1306] Andrew W. Appel and Xavier Leroy. Efficient extensional binary tries. Journal of Automated Reasoning, 67(1): ??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09655-x.

Schmoetten:2023:CTF

[1307] Richard Schmoetten, Jake E. Palmer, and Jacques D. Fleuriot. Correction: Towards formalising Schutz’ axioms for Minkowski spacetime in Isabelle/ HOL. Journal of Automated Reasoning, 67(1):??, March 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09651-1.  See

[1293].

Bentkamp:2023:SHO

[1308] Alexander Bentkamp, Jasmin Blanchette, Sophie Tourret, and Petar Vukmirović. Superposition for higher-order logic. Journal of Automated Reasoning, 67 (1):??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-022-09649-9.

Andrlon:2023:FNB

[1309] Mak Andrlon. Finding normal binary floating-point factors efficiently. Journal of Automated Reasoning, 67 (1):??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09659-1.

Windsor:2023:CAC

[1310] Aaron A. Windsor. Computer-aided constructions of commafree codes. Journal of Automated Reasoning, 67 (1):??, March 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09662-6.

Kirst:2023:SUI

[1311] Dominik Kirst and Marc Hermes. Synthetic undecidability and incompleteness of first-order axiom systems in Coq. Journal of Automated Reasoning, 67(1):??, March 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09647-x.

Middeldorp:2023:FOT

[1312] Aart Middeldorp, Alexander Lochmann, and Fabian Mitterwallner. Firstorder theory of rewriting for linear variable-separated rewrite systems: Automation, formalization, certification. Journal of Automated Reasoning, 67(2):??, June 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09661-7.

Yolcu:2023:AAC

[1313] Emre Yolcu, Scott Aaronson, and Marijn J. H. Heule. An automated approach to the Collatz conjecture. Journal of Automated Reasoning, 67 (2):??, June 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-022-09658-8.

Ebner:2023:US

[1314] Gabriel Ebner, Jasmin Blanchette, and Sophie Tourret. Unifying splitting. Journal of Automated Reasoning, 67(2):??, June 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09660-8.

Cristia:2023:AVP

[1315] Maximiliano Cristiá, Guido De Luca, and Carlos Luna. An automatically verified prototype of the Android permissions system. Journal of Automated Reasoning, 67 (2):??, June 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09666-2.

Nash:2023:ETM

[1316] Oliver Nash. Engel’s theorem in Mathlib. Journal of Automated Reasoning, 67(2):??, June 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09668-0.

Abate:2023:SPN

[1317] Alessandro Abate, Haniel Barbosa, Clark Barrett, Cristina David, Pascal Kesseli, Daniel Kroening, Elizabeth Polgreen, Andrew Reynolds, and Cesare Tinelli. Synthesising programs with non-trivial constants. Journal of Automated Reasoning, 67 (2):??, June 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09664-4.

Kaliszyk:2023:CHO

[1318] Cezary Kaliszyk and Karol Pak.

Combining higher-order logic with set theory formalizations. Journal of Automated Reasoning, 67 (2):??, June 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09663-5.

Cruz-Filipe:2023:FTC

[1319] Luís Cruz-Filipe, Fabrizio Montesi, and Marco Peressotti. A formal theory of choreographic programming. Journal of Automated Reasoning, 67 (2):??, June 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09665-3.

Leidinger:2023:SES

[1320] Hendrik Leidinger and Christoph Weidenbach. SCL(EQ): SCL for first– order logic with equality. Journal of Automated Reasoning, 67(3): ??, September 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09673-3.

Popescu:2023:RRB

[1321] Andrei Popescu. Rensets and renaming-based recursion for syntax with bindings extended version. Journal of Automated Reasoning, 67(3): ??, September 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09672-4.

Urban:2023:PLD

[1322] Christian Urban. POSIX lexing with derivatives of regular expressions. Journal of Automated Reasoning, 67 (3):??, September 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09667-1.

Holub:2023:BCD

[1323] Štěpán Holub, Martin Raška, and Štěpán Starosta. Binary codes that do not preserve primitivity. Journal of Automated Reasoning, 67(3): ??, September 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09674-2.

Luo:2023:RPS

[1324] Yun-Rong Luo, Che Cheng, and Jie-Hong R. Jiang. A resolution proof system for dependency stochastic Boolean satisfiability. Journal of Automated Reasoning, 67(3): ??, September 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09670-6.

Das:2023:CHS

[1325] Anupam Das and Marianna Girlando.

Cyclic hypersequent system for transitive closure logic. Journal of Automated Reasoning, 67(3):??, September 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-023-09675-
1.

Affeldt:2023:MCE

[1326] Reynald Affeldt and Cyril Cohen.

Measure construction by extension in dependent type theory with application to integration. Journal of Automated Reasoning, 67(3): ??, September 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09671-5.

Maggesi:2023:MGO

[1327] Marco Maggesi and Cosimo Perini Brogi. Mechanising Gödel–Löb provability logic in HOL light. Journal of Automated Reasoning, 67(3): ??, September 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09677-z.

Echenim:2023:PPS

[1328] Mnacho Echenim and Nicolas Peltier. A proof procedure for separation logic with inductive definitions and data. Journal of Automated Reasoning, 67 (3):??, September 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09680-4.

Reeves:2023:PPR

[1329] Joseph E. Reeves, Marijn J. H. Heule, and Randal E. Bryant. Preprocessing of propagation redundant clauses. Journal of Automated Reasoning, 67 (3):??, September 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-023-09681-3.

Sheng:2023:RAV

[1330] Ying Sheng, Andres Nötzli, Andrew Reynolds, Yoni Zohar, David Dill, Wolfgang Grieskamp, Junkil Park, Shaz Qadeer, Clark Barrett, and Cesare Tinelli. Reasoning about vectors:

Satisfiability modulo a theory of sequences. Journal of Automated Reasoning, 67(3):??, September 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09682-2.

Martin-Dorel:2023:EFP

[1331] Érik Martin-Dorel, Guillaume Melquiond, and Pierre Roux. Enabling floatingpoint arithmetic in the Coq proof assistant. Journal of Automated Reasoning, 67(4):??, December 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09679-x.

Sheng:2023:CSI

[1332] Ying Sheng, Yoni Zohar, Christophe Ringeissen, Andrew Reynolds, Clark Barrett, and Cesare Tinelli. Combining stable infiniteness and (strong) politeness. Journal of Automated Reasoning, 67(4):??, December 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09684-0.

Bohm:2023:LBQ

[1333] Benjamin Böhm and Olaf Beyersdorff.

Lower bounds for QCDCL via formula gauge. Journal of Automated Reasoning, 67(4):??, December 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09683-1.

Haselwarter:2023:FTT

[1334] Philipp G. Haselwarter and Andrej Bauer. Finitary type theories with and without contexts. Journal of Automated Reasoning, 67(4): ??, December 2023. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09678-y.

Indrzejczak:2023:BCF

[1335] Andrzej Indrzejczak. Bisequent calculus for four-valued quasi-relevant logics: Cut elimination and interpolation. Journal of Automated Reasoning, 67(4):??, December 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09685-z.

Peng:2023:SEA

[1336] Xicheng Peng, Jingzhong Zhang, Mao Chen, and Sannyuya Liu. Self-evident automated geometric theorem proving based on complex number identity. Journal of Automated Reasoning, 67(4):??, December 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09688-w.

Zheng:2023:SBB

[1337] Sen Zheng and Renate A. Schmidt. Saturation-based Boolean conjunctive query answering and rewriting for the guarded quantification fragments. Journal of Automated Reasoning, 67 (4):??, December 2023. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09687-x.

Faissole:2024:FVR

[1338] Florian Faissole. Formally-verified round-off error analysis of Runge– Kutta methods. Journal of Automated Reasoning, 68(1):??, March 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-023-09686-
y.

Echenim:2024:FCI

[1339] Mnacho Echenim and Mehdi Mhalla.

A formalization of the CHSH inequality and Tsirelson’s upper-bound in Isabelle/HOL. Journal of Automated Reasoning, 68(1):??, March 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-023-09689-
9.

Braun:2024:MBA

[1340] David Braun, Nicolas Magaud, and Pascal Schreck. A matroid-based automatic prover and Coq proof generator for projective incidence geometry. Journal of Automated Reasoning, 68(1):??, March 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09690-2.

Payet:2024:NTT

[1341] Étienne Payet. Non-termination in term rewriting and logic programming. Journal of Automated Reasoning, 68(1):??, March 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09693-z.

Bohm:2024:SDQ

[1342] Benjamin Böhm, Tomáš Peitl, and Olaf Beyersdorff. Should decisions in QCDCL follow prefix order? Journal of Automated Reasoning, 68(1): ??, March 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09694-6.

Geatti:2024:SMT

[1343] Luca Geatti, Nicola Gigante, Angelo Montanari, and Gabriele Venturato. SAT meets tableaux for linear temporal logic satisfiability. Journal of Automated Reasoning, 68 (2):??, June 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-023-09691-1.

Steinhofel:2024:SPP

[1344] Dominic Steinhöfel and Reiner Hähnle. Schematic program proofs with abstract execution. Journal of Automated Reasoning, 68(2):??, June 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-023-09692-
0.

Bernreiter:2024:SCC

[1345] Michael Bernreiter, Anela Lolic, Jan Maly, and Stefan Woltran. Sequent calculi for choice logics. Journal of Automated Reasoning, 68 (2):??, June 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09695-5.

Smola:2024:LRI

[1346] Filip Smola and Jacques D. Fleuriot. Linear resources in Isabelle/ HOL. Journal of Automated Reasoning, 68(2):??, June 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09698-2.

Li:2024:AGG

[1347] Lei Li, Zongkai Yang, Mao Chen, Xicheng Peng, Jianwen Sun, Zhonghua Yan, and Sannyuya Liu. Automated generation of geometry proof problems based on point geometry identity. Journal of Automated Reasoning, 68(2):??, June 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09699-1.

Nipkow:2024:GSV

[1348] Tobias Nipkow. Gale–Shapley verified. Journal of Automated Reasoning, 68(2):??, June 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09700-x.

Fiorentini:2024:GCS

[1349] Camillo Fiorentini and Mauro Ferrari.

General clauses for SAT–based proof search in intuitionistic propositional logic. Journal of Automated Reasoning, 68(3):??, September 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-024-09703-8.

Lammich:2024:RPA

[1350] Peter Lammich. Refinement of parallel algorithms down to LLVM: Applied to practically efficient parallel sorting. Journal of Automated Reasoning, 68(3):??, September 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-024-09701-w.

From:2024:VSC

[1351] Asta Halkjær From and Frederik Krogsdal Jacobsen. Verifying a sequent calculus prover for first-order logic with functions in Isabelle/HOL. Journal of Automated Reasoning, 68(3): ??, September 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-024-09697-3.

Choudhury:2024:DSC

[1352] Abhimanyu Choudhury and Meena Mahajan. Dependency schemes in CDCL–based QBF solving: a prooftheoretic study. Journal of Automated Reasoning, 68(3):??, September 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10817-024-09707-
                                        4.

Amir:2024:VGD

[1353] Guy Amir, Osher Maayan, Tom Zelazny, Guy Katz, and Michael Schapira. Verifying the generalization of deep learning to out-of-distribution domains. Journal of Automated Reasoning, 68(3):??, September 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09704-7.

Wang:2024:VPL

[1354] Zhongye Wang, Qinxiang Cao, and Yichen Tao. Verifying programs with logic and extended proof rules: Deep embedding vs. shallow embedding. Journal of Automated Reasoning, 68 (3):??, September 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09706-5.

Gross:2024:TSP

[1355] Jason Gross, Andres Erbsen, Jade Philipoom, Rajashree Agrawal, and Adam Chlipala. Towards a scalable proof engine: a performant prototype rewriting primitive for Coq. Journal of Automated Reasoning, 68(3): ??, September 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09705-6.

Malbos:2024:SSC

[1356] Philippe Malbos, Tanguy Massacrier, and Georg Struth. Single-set cubical categories and their formalisation with a proof assistant. Journal of Automated Reasoning, 68(4): ??, December 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09710-9.

HuertayMunive:2024:IIV

[1357] Jonathan Julián Huerta y Munive, Simon Foster, Mario Gleirscher, Georg Struth, Christian Pardillo Laursen, and Thomas Hickman. IsaVODEs: Interactive verification of cyberphysical systems at scale. Journal of Automated Reasoning, 68(4): ??, December 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-024-09709-2.

Chan:2024:WMH

[1358] Hing Lun Chan. Windmills of the minds: a hopping algorithm for Fermat’s two squares theorem. Journal of Automated Reasoning, 68(4): ??, December 2024. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-024-09708-3.

Cristia:2024:PDP

[1359] Maximiliano Cristiá and Gianfranco Rossi. A practical decision procedure for quantifier-free, decidable languages extended with restricted quantifiers. Journal of Automated Reasoning, 68(4):??, December 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-024-09713-6.

Wernhard:2024:IPS

[1360] Christoph Wernhard and Wolfgang Bibel. Investigations into proof structures. Journal of Automated Reasoning, 68(4):??, December 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09711-8.

Ayala-Rincon:2024:CFO

[1361] Mauricio Ayala-Rincón, Maribel Fernández, Gabriel Ferreira Silva, Temur Kutsia, and Daniele Nantes-Sobrinho. Certified first-order AC–unification and applications. Journal of Automated Reasoning, 68(4):??, December 2024. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09714-5. See cor-

rection [1367].

Baanen:2025:UAI

[1362] Anne Baanen. Use and abuse of instance parameters in the Lean mathematical library. Journal of Automated Reasoning, 69(1):??, March 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic).

URL https://link.springer.com/
article/10.1007/s10817-024-09712-
7.

Devadze:2025:CAP

[1363] Grigory Devadze, Victor Magron, and Stefan Streif. Computer-assisted proofs for Lyapunov stability via sums of squares certificates and constructive analysis. Journal of Automated Reasoning, 69(1):??, March 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09717-2.

Lipparini:2025:SNL

[1364] Enrico Lipparini and Stefan Ratschan.

Satisfiability of non-linear transcendental arithmetic as a certificate search problem. Journal of Automated Reasoning, 69(1):??, March 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09716-3.

Song:2025:FPN

[1365] Shuhao Song and Bowen Yao. Formalization of the prime number theorem with a remainder term. Journal of Automated Reasoning, 69(1): ??, March 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09718-9.

Beyer:2025:ISB

[1366] Dirk Beyer, Nian-Ze Lee, and Philipp Wendler. Interpolation and SATbased model checking revisited: Adoption to software verification. Journal of Automated Reasoning, 69(1): ??, March 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09702-9.

Ayala-Rincon:2025:CCF

[1367] Mauricio Ayala-Rincón, Maribel Fernández, Gabriel Ferreira Silva, Temur Kutsia, and Daniele Nantes-Sobrinho. Correction to: Certified first-order AC– unification and applications. Journal of Automated Reasoning, 69(1): ??, March 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-024-09715-4.  See

[1361].

Abrahamsson:2025:FVC

[1368] Oskar Abrahamsson, Magnus O.

Myreen, Michael Norrish, Hrutvik Kanabar, and Johannes Åman Pohjola. Fast, verified computation for HOL ITPs. Journal of Automated Reasoning, 69(1):??, March 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09719-8.

Borges:2025:LIT

[1369] Ana de Almeida Borges, Annalí Casanueva Artís, Jean-Rémy Falleri, Emilio Jesús Gallego Arias, Érik Martin-Dorel, Karl Palmskog, Alexander Serebrenik, and Théo Zimmermann. Lessons for interactive theorem proving researchers from a survey of Coq users. Journal of Automated Reasoning, 69(1):??, March 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09720-1.

Avigad:2025:PPC

[1370] Jeremy Avigad, Lior Goldberg, and Alon Titelman. A proof-producing compiler for blockchain applications. Journal of Automated Reasoning, 69 (2):??, June 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09723-y.

Echenim:2025:DPT

[1371] Mnacho Echenim and Nicolas Peltier.

A direct procedure to test entailment in a separation logic of relations. Journal of Automated Reasoning, 69(2):??, June 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09728-7.

Shamsu:2025:MVG

[1372] Sheera Shamsu, Dipesh Kafle, and KC Sivaramakrishnan. A mechanically verified garbage collector for OCaml. Journal of Automated Reasoning, 69(2):??, June 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09721-0.

Michaelson:2025:PPU

[1373] Dawn Michaelson, Dominik Schreiber, and Michael W. Whalen. Producing proofs of unsatisfiability with distributed clause-sharing SAT solvers. Journal of Automated Reasoning, 69 (2):??, June 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09725-w.

Bonacina:2025:QAQ

[1374] Maria Paola Bonacina, Stéphane Graham-Lengrand, and Christophe Vauthier. The QSMA algorithm for quantifiers in SMT. Journal of Automated Reasoning, 69 (2):??, June 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09727-8.

Tirore:2025:SCP

[1375] Dawit Tirore, Jesper Bengtson, and Marco Carbone. A sound and complete projection for global types. Journal of Automated Reasoning, 69 (2):??, June 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09726-9.

Sankur:2025:TAV

[1376] Ocan Sankur. Timed automata verification and synthesis via finite automata learning. Journal of Automated Reasoning, 69 (2):??, June 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09730-z.

Fuchs:2025:ALS

[1377] Tobias Fuchs, Jakob Bach, and Markus Iser. Active learning for SAT solver benchmarking. Journal of Automated Reasoning, 69(3): ??, September 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09729-6.

Garg:2025:DAF

[1378] Mohit Garg, N. Raja, and Abhishek Kr Singh. Double auctions: Formalization and automated checkers. Journal of Automated Reasoning, 69(3): ??, September 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09732-x.

Schmoetten:2025:CLA

[1379] Richard Schmoetten and Jacques D.

Fleuriot. Constructing the Lie algebra of smooth vector fields on a Lie group in Isabelle/HOL. Journal of Automated Reasoning, 69(3): ??, September 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09724-x.

Lisitsa:2025:ARP

[1380] Alexei Lisitsa, Zipei Nie, and Alexei Vernitski. Automated reasoning for proving non-orderability of groups. Journal of Automated Reasoning, 69 (3):??, September 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09734-9.

Blazek:2025:SPG

[1381] Jiří Blažek and Pavel Pech. Solving a problem with GeoGebra current possibilities and limits of CAS tools. Journal of Automated Reasoning, 69(3):??, September 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09733-w.

Cheung:2025:FVS

[1382] Louis Cheung, Alistair Moffat, and Christine Rizkallah. Formally verified suffix array construction. Journal of Automated Reasoning, 69(3): ??, September 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09735-8.

Dunn:2025:SMG

[1383] Lawrence Dunn, Val Tannen, and Steve Zdancewic. Structured monads for generic first-order syntax metatheory. Journal of Automated Reasoning, 69(3):??, September 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09731-y.

Mertens:2025:CEV

[1384] Hannah Mertens, Joost-Pieter Katoen, and Tobias Winkler. Computing expected visiting times and stationary distributions in Markov chains: Fast and accurate. Journal of Automated Reasoning, 69(3): ??, September 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09736-7.

Balachander:2025:LRS

[1385] Mrudula Balachander, Emmanuel Filiot, and Jean-François Raskin. LTL reactive synthesis with a few hints. Journal of Automated Reasoning, 69 (3):??, September 2025. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

//link.springer.com/article/10.
1007/s10817-025-09737-6.

Walther:2025:RAI

[1386] Christoph Walther. Reasoning about incompletely defined programs. Journal of Automated Reasoning, 69(3): ??, September 2025. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10817-025-09722-z.